### v1

# Principal stratification with $U$-statistics under principal ignorability

Abstract
Principal stratification is a popular framework for causal inference in the presence of an intermediate outcome. While the principal average treatment effects have traditionally been the default target of inference, it may not be sufficient when the interest lies in the relative favorability of one potential outcome over the other within the principal stratum. We thus introduce the principal generalized causal effect estimands, which extend the principal average causal effects to accommodate nonlinear contrast functions. Under principal ignorability, we expand the theoretical results in Jiang et al., (2022) to a much wider class of causal estimands in the presence of a binary intermediate variable. We develop identification formulas and derive the efficient influence functions of the generalized estimands for principal stratification analyses. These efficient influence functions motivate a set of multiply robust estimators and lay the ground for obtaining efficient debiased machine learning estimators via cross-fitting based on $U$-statistics. The proposed methods are illustrated through simulations and the analysis of a data example.
Keywords: causal inference, efficient influence function, principal stratification, multiply robust estimation, $U$-statistics, win-ratio.
1 Introduction
For causal inference under the Neyman-Rubin model (Rubin,, 1974), the average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest. Specifically, let $\{Y(1),Y(0)\}$ denote a pair of potential outcomes for a randomly selected individual from a super-population, and then the average causal effect estimand is defined as $\mathbb{E}\{Y(1)-Y(0)\}$. Conceptually, the average causal effect estimand can be expanded to a wider class, allowing arbitrary comparisons. That is, given a contrast function $h(u,v)$ defined on the product space, $\mathcal{Y}\times\mathcal{Y}$ ($\mathcal{Y}$: the potential outcome space), a generalized estimand can be defined as $\mathbb{E}[h\{Y(1),Y(0)\}]$. Examples include the average causal effect when $h(u,v)=u-v$, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$ (Lu et al.,, 2020). Because of the fundamental problem of causal inference, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$. However, a closely related quantity, the generalized causal effect (O’brien,, 1988; Mao,, 2017), can be formulated as:
$\tau_{h}=\int\int h(u,v)\mathrm{d}F_{1}(u)\mathrm{d}F_{0}(v),$ (1)
where $F_{z}(\cdot)$ is the marginal distribution of the potential outcome $Y(z)\in\mathcal{Y}$ for $z=0,1$. Consider $\{Y_{1}(1),Y_{1}(0)\}$ and $\{Y_{2}(1),Y_{2}(0)\}$ as two independent replicates of $\{Y(1),Y(0)\}$, and then $\tau_{h}$ can be re-expressed in a familiar form, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}]$. Two notable examples are the probability index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al.,, 2015), and the win-ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al.,, 2011). These estimable quantities are alternative measurements for the population causal effect and can be straightforwardly interpreted. For example, the probability index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than $F_{0}$, and is robust against extreme values in the tails of the potential outcome distributions.
When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al.,, 1996), mortality event (Zhang et al.,, 2009) or dropout (Frangakis and Rubin,, 1999), often have important implications even on the definition of a causal effect and require careful adjustment. The intermediate variable is not guaranteed to be balanced via randomization, and direct conditioning on the intermediate variable leads to sample selection bias and non-causal comparisons even in randomized experiments. As a solution, the principal stratification framework (Frangakis and Rubin,, 2002) segments the population into subpopulations (principal strata) according to joint potential values of the intermediate variable under alternative conditions. The stratum-specific causal effect—the principal causal effect—can be defined and is causally interpretable. However, a predominant focus in prior works has been on the average causal effect within each principal stratum, and there is a lack of related development when the generalized causal effect is of interest within each subpopulation. Therefore, in this article, we propose a new class of estimands, the principal generalized causal effect (PGCE), which accommodates an arbitrary and potentially nonlinear contrast function. Under a linear contrast function, the PGCE reduces to the usual principal average causal effect. As we demonstrate in due course, the PGCE can represent the stratum-specific probability index and strata-specific win-ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within important subpopulations.
Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al.,, 1996) and principal ignorability (Jo and Stuart,, 2009; Ding and Lu,, 2016), which are typically invoked for principal stratification analyses. The monotonicity assumption states that the treatment has a nonnegative causal effect on the intermediate outcome of each unit, effectively restricting the number of potential strata. This condition is considered plausible when, for example, the intermediate variable is death, and the treatment is known to benefit survival or when the intermediate variable is the actual treatment received and the treatment does not discourage treatment uptake. The principal ignorability assumption, on the other hand, requires that the observed pre-treatment covariates are sufficient to control for confounding between the intermediate variable and the final outcome. For the principal average causal effect, under monotonicity and principal ignorability, Ding and Lu, (2016) formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang et al., (2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and further developed multiply robust and locally efficient estimators under the additional assumption of treatment unconfoundedness. In contrast, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
The contributions of this article are several-folded. First, we propose a class of PGCE estimands in the presence of an intermediate variable, and develop complementary nonparametric identification formulas under treatment unconfoundedness, monotonicity, and principal ignorability. Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands, and seek efficiency improvement by deriving their corresponding efficient influence functions (Bickel et al.,, 1993, EIFs). Specifically, we show that, when the contrast function $h(u,v)=u-v$, our EIFs reduce to those derived by Jiang et al., (2022) as a special case, but are otherwise more general to accommodate arbitrary contrast functions. A major complexity in the context of a nonlinear contrast function, however, is that the estimating equations based on the EIFs requires nuisance functions defined over pairs of observations. Thus, to obtain a feasible estimator, we define ratio $U$-statistics whose influence functions are the same as the EIFs. We show that the $U$-estimators are triply robust in the sense that they are consistent and asymptotically normal when two of the three nuisance functions (including the treatment propensity score, principal score, and the potential outcome models) are correctly specified. When all nuisance functions are correctly specified, the $U$-estimators are locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds. The asymptotic properties of the $U$-estimators are developed upon the theory of empirical process (van der Vaart,, 1998, §19), $U$-process (Arcones and Gine,, 1993), as well as the theory of $U$-statistics (van der Vaart,, 1998, §11-12). Finally, the construction of the multiply robust $U$-estimators facilitates the integration of machine learning estimators for the nuisance functions. To achieve full asymptotic efficiency, we learn each nuisance function using machine learning methods, and operationalize the debiased machine learning $U$-estimators via cross-fitting (Chernozhukov et al.,, 2018). We discuss the modification needed for cross-fitting when a nonlinear contrast function is considered when defining the PGCE. We prove the multiple robustness and efficiency property of the debiased machine learning estimators under mild regulation conditions (Chernozhukov et al.,, 2022; Escanciano and Terschuur,, 2023).
The rest of this article is organized as follows. Section 2 introduces our causal estimands and identification assumptions for principal stratification. Section 3 discusses the nonparametric identification formulas for the PGCE and associated moment estimators. Section 4 derives the EIF for each PGCE estimand, and develops the EIF-motivated feasible $U$-estimators and the debiased machine learning estimators. Section 5 demonstrates the finite-sample performance of the proposed estimators via simulations. Section 6 illustrates the proposed methods with a data example from the United States Jobs Corps study (Schochet et al.,, 2001), and Section 7 concludes. Detailed proofs of all Theorems are available in the Online Supplement.
2 Notation, estimands and assumptions
2.1 Setting and notation
We consider a study (experimental or observational) with $n$ participants and two treatment conditions, and let $Z_{i}$ denote the treatment assignment of participant $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 00 the control. We use $D_{i}(z)$ and $Y_{i}(z)$ to denote the potential outcomes of the intermediate variable and the final outcome variable of interest for participant $i$ under treatment $Z_{i}=z$ ($z\in\{0,1\}$), respectively. The intermediate variable $D_{i}\in\{0,1\}$ is assumed to be binary, but there is no restriction on the type of variable for $Y_{i}$. Under the stable unit treatment value assumption (Rubin,, 1980, SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$. Let $X_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{X_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d).
2.2 Principal generalized causal effects
We operate under the principal stratification framework (Frangakis and Rubin,, 2002), and define distinct principal strata according to the joint potential values of the intermediate variable under alternative conditions. Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{11,10,01,00\}$. A key insight under the principal stratification framework is that, although $D$ is a post-randomization variable, the strata variable $S$ is unaffected by the assignment and can be viewed as a pre-treatment variable. Therefore, based on any contrast function $h$, we define the principal generalized causal effect estimand as
$\tau_{h}^{s}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=S_{2}=s],~{}\text{for}~{}s\in\{11,10,01,00\},$ (2)
where $Y_{1}(1)$ and $Y_{2}(0)$ denote the potential outcomes for two randomly selected units from the subpopulation with strata membership $s$ ($S_{1}=s$ and $S_{2}=s$ denote the strata membership for these two selected units). Of note, to ensure a proper causal interpretation of $\tau_{h}^{s}$, it is required that both elements of the pair for comparison are from the same principal stratum such that the comparison is made within a common subpopulation. If the two units are selected from different principal strata, the resulting quantity, $\tau_{h}^{s_{1},s_{2}}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2}]$ with $s_{1}\neq s_{2}$, remains estimable under this framework but generally lacks a meaningful causal interpretation; we thus do not further pursue this direction. Evidently, when a linear contrast function is specified as $h(u,v)=u-v$, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature; see, for example, Frangakis and Rubin, (2002), Jiang et al., (2022) and Lipkovich et al., (2022). When a nonlinear contrast function is considered in defining the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum. The following two examples illustrate the PGCE with a specific contrast function.
Example 1 (Probability index under truncation by death).
When $D$ represents the survival status (alive if $D=1$) that is ascertained before the measurement of the final outcome $Y$ (such as quality of life, which is only well-defined when $D=1$), there exist four principal strata based on the potential survival status. That is, $s=11$ represents the always-survivors who would survive until the measurement of the final outcome regardless of the assignment; $s=10$ and $s=01$ represent the protected and harmed, who would survive until the measurement of the final outcome only under the treatment condition and control condition, respectively; $s=00$ represents the never-survivors who would not survive until the measurement of the final outcome regardless of the assignment. Because the final outcome is oftentimes only well-defined without ambiguity when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al.,, 2009; Chen et al.,, 2024). In clinical research, Acion et al., (2006) argued that a more intuitive effect size can be measured by the probability index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$. Based on this contrast function, we can define the survivor probability index (SPI) as
$\mathrm{SPI}=\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=11\}.$ (3)
Compared to the SACE, the SPI is less sensitive to skewness of the underlying distributions of the potential outcomes and heteroskedasticity between conditions and outliers. In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg,, 1971).
Example 2 (Win estimands under noncompliance).
In experimental studies with ordinal outcomes, Wang and Pocock, (2016) introduced the win-ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\allowbreak\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$. This estimand corresponds to a two-dimensional contrast function specified as $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$. If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{11,10,01,00\}$ refers to always-takers, compliers, defiers, and never-takers, respectively (Angrist et al.,, 1996). Therefore, the win-ratio estimands can be extended to define the compliers win-ratio (CWR) as
$\mathrm{CWR}=\frac{\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=S_{2}=10\}}{\mathbb{P}\{Y_{1}(1)<Y_{2}(0)|S_{1}=S_{2}=10\}}.$ (4)
The CWR estimand represents the odds of a randomly selected unit from all-treated compliers having a more favorable outcome than one from the all-control compliers subpopulation. Alternatively, the complier win-difference (CWD) or complier net benefit estimand on the difference scale is given as
$\mathrm{CWD}=\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=S_{2}=10\}-\mathbb{P}\{Y_{1}(1)<Y_{2}(0)|S_{1}=S_{2}=10\}.$ (5)
The win estimands within other principal strata can be defined similarly. Mao, (2024) considered the point and interval identification of a similar estimand using the instrumental variable approach.
2.3 Identification assumptions
Even in experimental studies, the PGCE estimands cannot be identified without additional assumptions because the pair of potential intermediate variables and that of potential final outcomes are never observed simultaneously. To identify the PGCE from the observed data, we require the following assumptions.
Assumption 1 (Treatment ignorability).
[…]
Assumption 1 rules out confounding between the treatment assignment and intermediate as well as final outcomes conditional on observed pre-treatment covariates. This assumption holds by design in randomized experiments. For observational studies, Assumption 1 can be satisfied if all confounders affecting the treatment assignment, the intermediate variable, and the final outcome are measured and included in the measured set of pre-treatment covariates, $X$.
Assumption 2 (Monotonicity).
$D_{i}(1)\geq D_{i}(0)$ for all $i$.
Assumption 2 states that the treatment assignment only affects the intermediate outcome in a non-negative (or non-positive if assume $D_{i}(1)\leq D_{i}(0)$) direction, and restricts the number of possible strata by excluding the harmed or defiers stratum $S=01$ (or $S=10$). For example, when $D$ represents the survival status, monotonicity holds when the treatment is expected not to harm the survival of each unit. When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to active treatment.
Assumption 3 (Mean principal ignorability).
Given the same values of pre-treatment covariates for a pair of units ($X_{1}=x_{1}$ and $X_{2}=x_{2}$), the expected contrast function is exchangeable across certain pairs of principal strata. That is,
$\begin{split}&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2},X_{1}=x_{1},X_{2}=x_{2}]\\ =~{}&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1}^{\prime},S_{2}=s_{2}^{\prime},X_{1}=x_{1},X_{2}=x_{2}]\end{split}$ (6)
for (a) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{10}=\{(10,10),(11,10),(10,00),(11,00)\}$; (b) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{11}=\{(11,11),(10,11)\}$; (c) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{00}=\{(00,00),(00,10)\}$.
Assumption 3 rules out unmeasured confounding between the principal strata membership and the final outcome, ensuring that the PGCE components can be directly estimated from the observed data. For example, under Assumptions 1-3, the conditional expectations in group $\mathcal{S}_{10}$ are all equal to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$. In other words, within the observed data stratum defined by $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the combinations of possible latent principal strata, $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$, are exchangeable with respect to the expected contrast. Similarly, the conditional expected contrasts within group $\mathcal{S}_{11}$ and $\mathcal{S}_{00}$ are all equal to estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=2,X_{1},X_{2}]$ and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, respectively. That is, with Assumption 3, each one of the subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only.
Remark 1.
[…]
Assumptions 2 and 3 are instrumental in achieving nonparametric identification of the PGCE, however, their applicability, in general, cannot be directly tested using observed data alone without prior knowledge. Although our primary focus is on identifying and efficiently estimating the PGCE under these two structural assumptions, we provide a confounding function approach for sensitivity analyses under assumed violations of these two structural assumptions in Section S7 of the Online Supplement.
3 Nonparametric identification
3.1 Identification of the PGCE estimands
To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution, i.e., the treatment probability, the principal score, and the mean of the outcome contrast conditional on covariates (referred to as pairwise outcome mean hereafter for brevity). We consider both the observational and experimental studies, and define the treatment probability or the propensity score, by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin,, 1983). In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu,, 2016), for $s\in\{10,00,11\}$ under Assumption 2. Furthermore, we use averages over the distribution of covariates, $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$, to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively. Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\},$
for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. We provide some examples of the pairwise outcome model models for specific contrast functions below.
Example 3 (Pairwise mean function for average treatment effect).
For studying the average treatment effect, the contrast function is specified as $h(u,v)=u-v$. If the potential outcomes are continuous, a Gaussian regression model is often assumed for the potential outcomes, i.e., $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$.Here, $f_{z}(D,X)$ is the mean function of $D$ and $X$ given treatment status $Z=z$. We then have the pairwise outcome mean function given by
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2}).$ (7)
Example 4 (Pairwise mean function for probability index).
For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3. In this case, following Wolfe and Hogg, (1971), one can derive the pairwise outcome mean as
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\Phi\left\{\frac{f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2})}{\sqrt{2}\sigma}\right\}.$ (8)
Example 5 (Pairwise outcome mean for win estimands).
Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand). For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao, (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ (9)
=$\quad=\left\{\sum_{q=1}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2}),\sum_{q=1}^{Q}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2})\right\},$
where $p_{zq}(D,X)$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $X$, $Q$ is the largest category of the ordinal outcome.
For estimating the principal scores, we first define conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$. Then under Assumption 2, the principal scores can be expressed as (Jiang et al.,, 2022)
$e_{10}(X)=p_{1}(X)-p_{0}(X),~{}~{}~{}e_{00}(X)=1-p_{1}(X),~{}~{}~{}e_{11}(X)=p_{0}(X).$
This is because, with monotonicity, the observed stratum $(Z=0,D=1)$ only contains stratum $S=11$, and observed stratum $(Z=1,D=0)$ only contains stratum $S=00$. The remaining principal score for stratum $S=10$ can thus be obtained by subtracting $e_{11}(X)$ and $e_{00}(X)$ from one. The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$. These elements then prepare us to state the set of nonparametric identification results in the following Theorem 6.
Theorem 6 (Nonparametric identification).
Suppose Assumptions 1-3 hold, $e_{s}>0$ for $s\in\{10,00,11\}$, and $10<\pi(X)<$ 1 for all $X\in\mathcal{X}$, the support of covariates. Then, based on different components of the observed data distribution, the following identification formulas for each PGCE can be obtained:
  • (a)
    Based on the treatment probability and the principal score,
    $\tau_{h}^{10}$ $=\mathbb{E}\left[\left\{\frac{e_{10}(X_{1})}{p_{1}-p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{10}(X_{2})}{p_{1}-p_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right],$
    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-D_{1}}{1-p_{1}}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{00}(X_{2})}{1-p_{1}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right],$
    $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{e_{11}(X_{1})}{p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{D_{2}}{p_{0}}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right];$
  • (b)
    Based on the treatment probability and the pairwise outcome mean,
    $\tau_{h}^{10}$ $=\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
    ×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2})-D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{1}-p_{0}}\right]\mu_{1100}(X_{1},X_{2})\right),$
    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-D_{1}Z_{1}/\pi(X_{1})}{1-p_{1}}\right\}\left\{\frac{1-D_{2}Z_{2}/\pi(X_{2})}{1-p_{1}}\right\}\mu_{1000}(X_{1},X_{2})\right],$
    $\tau_{h}^{11}$ $=\mathbb{E}\left(\left[\frac{D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{0}}\right]\left[\frac{D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{0}}\right]\mu_{1101}(X_{1},X_{2})\right);$
  • (c)
    Based on the principal score and the pairwise outcome mean,
    $\tau_{h}^{10}$ $=\mathbb{E}\left[\left\{\frac{p_{1}(X_{1})-p_{0}(X_{1})}{p_{1}-p_{0}}\right\}\left\{\frac{p_{1}(X_{2})-p_{0}(X_{2})}{p_{1}-p_{0}}\right\}\mu_{1100}(X_{1},X_{2})\right],$
    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-p_{1}(X_{1})}{1-p_{1}}\right\}\left\{\frac{1-p_{1}(X_{2})}{1-p_{1}}\right\}\mu_{1000}(X_{1},X_{2})\right],$
    $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{p_{0}(X_{1})}{p_{0}}\right\}\left\{\frac{p_{0}(X_{2})}{p_{0}}\right\}\mu_{1101}(X_{1},X_{2})\right].$
We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 6, and the identification of the other two PGCEs follows the same strategy. In Theorem 6(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$. Specifically, under Assumptions 1 and 2, we have
$\mathbb{E}\left[\left\{\frac{e_{10}(X_{1})}{p_{1}-p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{10}(X_{2})}{p_{1}-p_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}\bigg{\rvert}X_{1},X_{2}\right]=\frac{e_{10}(X_{1})e_{10}(X_{2})}{e_{10}^{2}},$
which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by the Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$. This ensures the identification of $\tau_{h}^{10}$. In Theorem 6(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
$\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
×$\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2})-D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{1}-p_{0}}\right]\bigg{\rvert}X_{1},X_{2}\right)=\frac{e_{10}(X_{1})e_{10}(X_{2})}{e_{10}^{2}}.$
Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$. Finally, in Theorem 6(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 6(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$. The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.
Remark 2.
Theorem 6 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al.,, 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al., (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
3.2 Estimators based on nonparametric identification formulas
For point estimation, a common practice for operationalizing the nonparametric identification formulas is to assume parametric models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$. We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$. To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment probability, the principal score, and the outcome mean are correctly specified, respectively. Specifically, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Additionally, we use ‘+’ in the subscript to denote conditions where more than one nuisance function is correctly specified. In addition, the union notation, $\cup$’, denotes the condition where at least one nuisance function is correctly specified.
We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts, and expectations with empirical averages. For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$. Along the lines in Jiang et al., (2022), we define the following quantity for any function $f(Y,D,X)$:
[…]
One example of (10) is having $f(Y_{z},D_{z},X)=D_{z}$, in which case we obtain the following function:
$\psi_{D_{z}}(\mathcal{O})=\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X)\}}{\mathbb{P}(Z=z|X)}+p_{z}(X).$ (11)
Equation (11) is useful in the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome (Tsiatis,, 2006):
$\widehat{p}_{z}\equiv\mathbb{E}_{n}\psi_{D_{z}}(\mathcal{O};\widehat{\alpha},\widehat{\beta})=\mathbb{E}_{n}\left[\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X;\widehat{\beta})\}}{\pi(X;\widehat{\alpha})}+p_{z}(X;\widehat{\beta})\right],$ (12)
which improves upon the plug-in estimator $\mathbb{E}_{n}p_{z}(X;\widehat{\beta})$ in terms of estimation efficiency and robustness, and is consistent for $p_{z}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$.
Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart,, 1998). For example, Theorem 6(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
$\widehat{\tau}_{h,\pi+p}^{10}=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{1}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right].$
Theorem 6(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
$\widehat{\tau}_{h,\pi+\mu}^{10}$ $=\mathbb{U}_{n}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1};\widehat{\alpha})-D_{1}(1-Z_{1})/\{1-\pi(X_{1};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\right.$
×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2};\widehat{\alpha})-D_{2}(1-Z_{2})/\{1-\pi(X_{2};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right),$
and Theorem 6(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
$\widehat{\tau}_{h,p+\mu}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{p_{1}(X_{1};\widehat{\beta})-p_{0}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\left\{\frac{p_{1}(X_{2};\widehat{\beta})-p_{0}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right].$
The constructions of plug-in estimators for the remaining PGCEs are similar and, therefore, omitted. These estimators have stronger requirements on model specification to estimate the PGCEs consistently. Specifically, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment probability and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
Remark 3.
The weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score, which is an extension of the inverse propensity score weighted estimator in Mao, (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable. The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately after applying the Taylor’s expansion on the ratio estimators (Jiang et al.,, 2022); see the Online Supplement for additional details. The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be straightforwardly obtained using empirical process theory for the parametric class (van der Vaart,, 1998, §19.2). The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao, (2017).
4 Multiply robust and efficient $U$-estimators
Although implementing the plug-in estimators motivated by the nonparametric identification formulas is straightforward, their various forms and potential inefficiency call for alternative estimators with improved statistical properties. Here, we develop alternative $U$-estimators that exploit information across all nuisance functions, including the treatment assignment, the intermediate variable, and the final outcome of interest.
4.1 The efficient influence function for the PGCE
We first develop more efficient estimators with root-$n$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al.,, 1993). The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators. Several key properties of the EIF suggest its promise for constructing optimal estimators. First, the EIF often integrates more than one nuisance model and facilitates estimators that are robust against misspecification of some nuisance models (Tsiatis,, 2006; Mao,, 2017). Second, the asymptotic variance lower bound is represented by the variance of the EIF, providing a useful benchmark to define efficient estimators. Third, the EIF provides a device for integrating advanced machine learners for the nuisance parameters, which can lead to feasible estimators that achieve the asymptotic variance lower bound (Chernozhukov et al.,, 2018; Rotnitzky et al.,, 2021). We will explore these properties by first formalizing the EIF for the PGCE. Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(X_{1})e_{10}(X_{2})\allowbreak\mu_{1100}(X_{1},X_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$. This suggests that the EIF for $\tau_{h}^{s}$, $\varphi_{h}^{s}(\mathcal{O})$, is of the form:
$\varphi_{h}^{s}(\mathcal{O})\equiv\frac{1}{\tau_{D}^{s}}\varphi_{h,N}^{s}(\mathcal{O})-\frac{\tau_{h,N}^{s}}{\tau_{D}^{s}{}^{2}}\varphi_{D}^{s}(\mathcal{O})=\frac{\varphi_{h,N}^{s}(\mathcal{O})-\tau_{h}^{s}\varphi_{D}^{s}(\mathcal{O})}{\tau_{D}^{s}},$
where $\varphi_{h,N}^{s}(\mathcal{O})$ and $\varphi_{D}^{s}(\mathcal{O})$ are EIFs for $\tau_{h,N}^{s}$ and $\tau_{D}^{s}$, respectively. Let $F_{zd}(\cdot|x)$ denote the conditional distribution of $Y_{i}(z)$ given $(Z_{i},D_{i},X_{i})=(z,d,x)$, and $F_{X}(\cdot)$ denote the marginal distribution of $X$. We can further define marginalized functions of the observed outcome, $\nu_{1,zd}^{s}(Y)=\int e_{s}(x)h(Y,y)\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$ and $\nu_{0,zd}^{s}(Y)=\int e_{s}(x)h(y,Y)\allowbreak\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$, for $z,d\in\{0,1\}$ and $s\in\{10,00,11\}$. Then, the following results are obtained.
Theorem 7 (EIF).
Under Assumptions 1-3, (a) the EIF for $\tau_{h}^{10}$ is $\varphi_{h}^{10}(\mathcal{O})=\{\varphi_{h,N}^{10}(\mathcal{O})-\tau_{h}^{10}\varphi_{D}^{10}(\mathcal{O})\}/(p_{1}-p_{0})^{2}$, where $\varphi_{h,N}^{10}(\mathcal{O})=\phi_{h}^{10}(\mathcal{O})-2\tau_{h,N}^{10}$ and
$\phi_{h}^{10}(\mathcal{O})=$ $~{}\frac{ZDe_{10}(X)}{\pi(X)p_{1}(X)}\left\{\nu_{1,00}^{10}(Y)-m_{1}^{10}(X)\right\}+\frac{(1-Z)(1-D)e_{10}(X)}{\{1-\pi(X)\}\{1-p_{0}(X)\}}\left\{\nu_{0,11}^{10}(Y)-m_{0}^{10}(X)\right\}$
+$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(X)+m_{0}^{10}(X)\right\},$
with $m_{1}^{10}(X)=\int e_{10}(x)\mu_{1100}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{10}(X)=\int e_{10}(x)\mu_{1100}(x,X)\mathrm{d}F_{X}(x)$, and
$\varphi_{D}^{10}(\mathcal{O})=2\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}(p_{1}-p_{0})-2\tau_{D}^{10}.$
(b) The EIF for $\tau_{h}^{00}$ is $\varphi_{h}^{00}(\mathcal{O})=\{\varphi_{h,N}^{00}(\mathcal{O})-\tau_{h}^{00}\varphi_{D}^{00}(\mathcal{O})\}/(1-p_{1})^{2}$, where $\varphi_{h,N}^{00}(\mathcal{O})=\phi_{h}^{00}(\mathcal{O})-2\tau_{h,N}^{00}$ and
$\phi_{h}^{00}(\mathcal{O})=$ $~{}\frac{Z(1-D)}{\pi(X)}\left\{\nu_{1,00}^{00}(Y)-m_{1}^{00}(X)\right\}+\frac{(1-Z)(1-D)e_{00}(X)}{\{1-\pi(X)\}\{1-p_{0}(X)\}}\left\{\nu_{0,10}^{00}(Y)-m_{0}^{00}(X)\right\}$
$+\psi_{1-D_{1}}(\mathcal{O})\left\{m_{1}^{00}(X)+m_{0}^{00}(X)\right\},$
with $m_{1}^{00}(X)=\int e_{00}(x)\mu_{1000}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{00}(X)=\int e_{00}(x)\mu_{1000}(x,X)\mathrm{d}F_{X}(x)$, and
$\varphi_{D}^{00}(\mathcal{O})=2\psi_{1-D_{1}}(\mathcal{O})(1-p_{1})-2\tau_{D}^{00}.$
(c) The EIF for $\tau_{h}^{11}$ is $\varphi_{h}^{11}(\mathcal{O})=\{\varphi_{h,N}^{11}(\mathcal{O})-\tau_{h}^{11}\varphi_{D}^{11}(\mathcal{O})\}/p_{0}^{2}$, where $\varphi_{h,N}^{11}(\mathcal{O})=\phi_{h}^{11}(\mathcal{O})-2\tau_{h,N}^{11}$ and
$\phi_{h}^{11}(\mathcal{O})=$ $~{}\frac{ZDe_{11}(X)}{\pi(X)p_{1}(X)}\left\{\nu_{1,01}^{11}(Y)-m_{1}^{11}(X)\right\}+\frac{(1-Z)D}{1-\pi(X)}\left\{\nu_{0,11}^{11}(Y)-m_{0}^{11}(X)\right\}$
$+\psi_{D_{0}}(\mathcal{O})\left\{m_{1}^{11}(X)+m_{0}^{11}(X)\right\},$
with $m_{1}^{11}(X)=\int e_{11}(x)\mu_{1101}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{11}(X)=\int e_{11}(x)\mu_{1101}(x,X)\mathrm{d}F_{X}(x)$, and
$\varphi_{D}^{11}(\mathcal{O})=2\psi_{D_{0}}(\mathcal{O})p_{0}-2\tau_{D}^{11}.$
The EIFs in Theorem 7 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey, (2022) and Hines et al., (2022) with respect to the distribution of observed data, $\mathcal{O}$. The derivation for $\varphi_{D}^{s}(\mathcal{O})$ is then follows the chain rule. For $\varphi_{h,N}^{s}(\mathcal{O})$, we re-express the PGCE in terms of the observed data as
$\tau_{h}^{s}=\mathbb{E}\{\mathfrak{h}_{1}^{s}(Y)|Z=1\}+\mathbb{E}\{\mathfrak{h}_{0}^{s}(Y)|Z=0\},$
where $\mathfrak{h}_{1}^{s}(y)=\mathbb{E}\{h(y,Y^{\prime})|Z^{\prime}=0,S=S^{\prime}=s\}$ and $\mathfrak{h}_{0}^{s}(y)=\mathbb{E}\{h(Y^{\prime},y)|Z^{\prime}=1,S=S^{\prime}=s\}$, with the conditional expectations taken with respect to $Y^{\prime}$. Importantly, the EIF for the PGCE remains unchanged after this re-expression, since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged. Based on Theorem 7, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
Remark 4.
Without considering the intermediate variable, the EIF for the generalized causal effect estimand, $\tau_{h}$ in (1), $\varphi_{h}(\mathcal{O})$, can be derived using the same approach as $\varphi_{h}(\mathcal{O})=\phi_{h}(\mathcal{O})-2\tau_{h}$, where
$\phi_{h}(\mathcal{O})=\frac{Z}{\pi(X)}\left\{\nu_{1}(Y)-m_{1}(X)\right\}+\frac{1-Z}{1-\pi(X)}\left\{\nu_{0}(Y)-m_{0}(X)\right\}+m_{1}(X)+m_{0}(X),$
with $\nu_{z}(Y)=\int h(Y,y)\mathrm{d}F_{z}(y|x)\mathrm{d}F_{X}(x)$, $m_{1}(X)=\int\mu_{10}(X,x)\mathrm{d}F_{X}(x)$ and $m_{0}(X)=\int\mu_{10}(x,X)\allowbreak\mathrm{d}F_{X}(x)$. The function $\mu_{10}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=1,Z_{2}=0,X_{1},X_{2}\}$ is the pairwise outcome mean within the observed stratum $(Z_{1}=1,Z_{2}=0)$. This coincides with the EIF developed in Mao, (2017) using the projection approach (Tsiatis,, 2006), and intuitively, $\varphi_{h}(\mathcal{O})$ can be viewed as a special case of $\varphi_{h,N}^{s}(\mathcal{O})$ in the absence of the intermediate variable.
4.2 Multiply robust estimators
The EIFs given in Theorem 7 imply the semiparametric efficiency lower bound, and pave the way for formulating estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator. We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 7, thereby achieving semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that each PGCE can be rewritten as the following:
$\tau_{h}^{10}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{10})}{\{\mathbb{E}(\psi_{D_{1}}-\psi_{D_{0}})\}^{2}},~{}~{}~{}\tau_{h}^{00}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{00})}{\{\mathbb{E}(\psi_{1-D_{1}})\}^{2}},~{}~{}~{}\tau_{h}^{11}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{11})}{\{\mathbb{E}(\psi_{D_{0}})\}^{2}}.$ (13)
Continuing with parametric models for nuisance functions, the denominators in (13) can be estimated using the doubly robust estimator given in (12). Specifically, let $\theta$ denote the collection of parametric nuisance function parameters and $\vartheta$ denote a subset of $\theta$, i.e., $\theta=\{\alpha,\beta,\gamma\}$ and $\vartheta=\{\alpha,\beta\}$, with corresponding maximum likelihood estimators $\widehat{\theta}=\{\widehat{\alpha},\widehat{\beta},\widehat{\gamma}\}$ and $\widehat{\vartheta}=\{\widehat{\alpha},\widehat{\beta}\}$, as well as true values $\theta^{*}=\{\alpha^{*},\beta^{*},\gamma^{*}\}$ and $\vartheta^{*}=\{\alpha^{*},\beta^{*}\}$. Then, for example, the denominator of $\tau_{h}^{10}$ given in (13) can be estimated by $[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}$, where $\mathbb{E}_{n}\psi_{D_{z}}(\widehat{\vartheta})\equiv\mathbb{E}_{n}\psi_{D_{z}}(\mathcal{O};\widehat{\alpha},\widehat{\beta})$ defined in (12).
Estimating the numerators in (13) is more complicated. To target these terms, we construct the following functions of the observed pairs:
$g_{h}^{10}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$
$=2^{-1}\bigg{[}\frac{Z_{i}D_{i}e_{10}(X_{i};\beta)}{\pi(X_{i};\alpha)p_{1}(X_{i};\beta)}\times\frac{(1-Z_{j})(1-D_{j})e_{10}(X_{j};\beta)}{\{1-\pi(X_{j};\alpha)\}\{1-p_{0}(X_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1100}(X_{i},X_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}D_{j}e_{10}(X_{j};\beta)}{\pi(X_{j};\alpha)p_{1}(X_{j};\beta)}\times\frac{(1-Z_{i})(1-D_{i})e_{10}(X_{i};\beta)}{\{1-\pi(X_{i};\alpha)\}\{1-p_{0}(X_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1100}(X_{j},X_{i};\gamma)\right\}$
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(X_{i},X_{j};\gamma)$
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(X_{j},X_{i};\gamma)\bigg{]},$
and
$g_{h}^{00}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$
$=2^{-1}\bigg{[}\frac{Z_{i}(1-D_{i})}{\pi(X_{i};\alpha)}\times\frac{(1-Z_{j})(1-D_{j})e_{00}(X_{j};\beta)}{\{1-\pi(X_{j};\alpha)\}\{1-p_{0}(X_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1000}(X_{i},X_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}(1-D_{j})}{\pi(X_{j};\alpha)}\times\frac{(1-Z_{i})(1-D_{i})e_{00}(X_{i};\beta)}{\{1-\pi(X_{i};\alpha)\}\{1-p_{0}(X_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1000}(X_{j},X_{i};\gamma)\right\}$
$\qquad\qquad+\psi_{1-D_{1}}(\mathcal{O}_{i};\alpha,\beta)\psi_{1-D_{1}}(\mathcal{O}_{j};\alpha,\beta)\mu_{1000}(X_{i},X_{j};\gamma)$
$\qquad\qquad+\psi_{1-D_{1}}(\mathcal{O}_{j};\alpha,\beta)\psi_{1-D_{1}}(\mathcal{O}_{i};\alpha,\beta)\mu_{1000}(X_{j},X_{i};\gamma)\bigg{]},$
and
$g_{h}^{11}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ $=2^{-1}\bigg{[}\frac{Z_{i}D_{i}e_{11}(X_{i};\beta)}{\pi(X_{i};\alpha)p_{1}(X_{i};\beta)}\times\frac{(1-Z_{j})D_{j}}{1-\pi(X_{j};\alpha)}\left\{h(Y_{i},Y_{j})-\mu_{1101}(X_{i},X_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}D_{j}e_{11}(X_{j};\beta)}{\pi(X_{j};\alpha)p_{1}(X_{j};\beta)}\times\frac{(1-Z_{i})D_{i}}{1-\pi(X_{i};\alpha)}\left\{h(Y_{j},Y_{i})-\mu_{1101}(X_{j},X_{i};\gamma)\right\}$
$\qquad\qquad+\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\mu_{1101}(X_{i},X_{j};\gamma)$
$\qquad\qquad+\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\mu_{1101}(X_{j},X_{i};\gamma)\bigg{]},$
for $1\leq i\neq j\leq n$. For $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$, it is straightforward to show that $2\mathbb{E}\{g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})|\mathcal{O}_{i}\}=\phi_{h}^{s}(\mathcal{O}_{i})$ and $\mathbb{E}\{g_{h}^{s}(\mathcal{O}_{i},\allowbreak\mathcal{O}_{j};\theta^{*})\}=2^{-1}\mathbb{E}\{\phi_{h}^{s}(\mathcal{O}_{i})\}=\tau_{h,N}^{s}$. Thus, the numerators can be estimated using an empirical average of all pairs as $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})\equiv\binom{n}{2}^{-1}\sum_{i<j}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta})$ for $s\in\{10,00,11\}$. Therefore, we can obtain the following estimators for the PGCEs expressed in (13):
$\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (14)
which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 8.
Theorem 8 (Multiply robustness and efficiency under parametric working models).
Under Assumptions 13, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$. Then, the estimator $\widehat{\tau}_{h}^{s}$ in (14) is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
Under Theorem 8, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified, and is semiparametrically efficient if all three nuisance models are correctly specified. The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator. The denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$. The consistency of the numerator is more involved and requires calculating the asymptotic bias of $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})-\tau_{h,N}^{s}$. The specific derivations are provided in the Online Supplement Section S4.1 due to the complexity of the expressions. The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao, (2017) and Theorem 3 of Jiang et al., (2022). As discussed in Remark 3, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator. Details of the proof are provided in Section S4.2 of the Online Supplement. Finally, The variances of the triply robust estimators given in (14) can be estimated via nonparametric bootstrap.
4.3 Debiased machine learning estimators
Debiased machine learning is a burgeoning method for efficient estimation of the causal effects in experimental and observation studies, and is often built upon the orthogonal moment functions (Chernozhukov et al.,, 2018, 2022) or the EIFs. Because of the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al.,, 2022), perturbations in the first-step estimates typically have no significant influence on local properties of the desired estimator for the causal estimand. In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates (Rotnitzky et al.,, 2021). Compared to parametric working models, leveraging machine learning working models for the nuisance functions is more advantageous for its robustness against potential misspecification. However, machine learning first-step estimates ($\widehat{\pi}(\cdot)$, $\widehat{p}_{z}(\cdot)$, and $\widehat{\mu}_{z_{1}d_{1}z_{2}d_{2}}(\cdot,\cdot)$) have increased complexity and could no longer satisfy the Donsker class condition (Chernozhukov et al.,, 2018). To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting is adopted, where the entire sample is divided into separate parts for the estimation of the nuisance functions and the estimation of the causal estimand respectively (Chernozhukov et al.,, 2018). Mild regulatory conditions, mostly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy,, 2023).
Because of the unique structure of the PGCEs, we adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur, (2023). Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition =$\mathcal{I}=\{I_{1},\ldots,I_{K}\}$. Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $(i,j)$ with $1\leq i<j\leq N$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$. Chernozhukov et al., (2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$. The following example is used to provide a demonstration of $\mathcal{I}$ and $\mathcal{P}$.
Example 9 (Sample splitting for $U$-statistics).
Consider a scenario where $n=10$ and $K=3$. We can first form the partition =$\mathcal{I}=\{I_{1},I_{2},I_{3}\}$ with $I_{1}=\{1,2,3\}$, $I_{2}=\{4,5,6,7\}$, and $I_{3}=\{8,9,10\}$. Then, the partition $\mathcal{P}$ is formed as
$\{(I_{1},I_{1}),(I_{1},I_{2}),(I_{1},I_{3}),(I_{2},I_{2}),(I_{2},I_{3}),(I_{3},I_{3})\}.$
For $\mathcal{P}$, elements formed by $I_{k_{1}}$ and $I_{k_{2}}$ with $k_{1}<k_{2}$ are more straightforward as the Cartesian product of $I_{k_{1}}$ and $I_{k_{2}}$. For instance,
[…]
Elements formed by $I_{k}$ itself, on the other hand, need to exclude pairs consisting of the same indices, e.g., $(I_{1},I_{1})=\{(1,2),(1,3),(2,3)\}$. According to this sampling-splitting scheme, all $\binom{10}{2}=45$ possible pairs are divided into $L=6$ subsets.
With a slight abuse of notation, let $\widehat{\theta}=\{\widehat{\pi},\widehat{p}_{z},\widehat{\mu}_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\widehat{\vartheta}=\{\widehat{\pi},\widehat{p}_{z}\}$ denote machine learning estimators of nuisance functions $\theta=\{\pi,p_{z},\mu_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\vartheta=\{\pi,p_{z}\}$, respectively, as well as true functions $\theta^{*}=\{\pi^{*},p_{z}^{*},\mu_{z_{1}d_{1}z_{2}d_{2}}^{*}\}$ and $\vartheta^{*}=\{\pi^{*},p_{z}^{*}\}$. For a given partition $\mathcal{P}$ (based on the initial partition of singletons $\mathcal{I}$), we obtain intermediate estimates of the target estimand using observations in each $P_{l}$ (or $I_{k}$) with $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$) estimated using observations in $P_{l}^{c}$ (or $I_{k}^{c}$), those are not present in $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$). By pooling these intermediate estimates together, we can obtain the following cross-fitting estimators for the PGCEs:
$\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n,L}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n,K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n,L}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n,K}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n,L}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n,K}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (15)
where
$\mathbb{U}_{n,L}g_{h}^{s}(\widehat{\theta})=\binom{n}{2}^{-1}\sum_{l=1}^{L}\sum_{(i,j)\in P_{l}}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})$
and
$\mathbb{E}_{n,K}\psi_{D_{z}}(\widehat{\vartheta})=n^{-1}\sum_{k=1}^{K}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}).$
Remark 5.
Estimators given in (15) are the type II estimator proposed in Chernozhukov et al., (2018) for estimands without an intermediate variable. Alternatively, one could first obtain separate estimates
$\mathbb{U}_{|P_{l}|}g_{h}^{s}(\widehat{\theta}_{l})=|P_{l}|^{-1}\sum_{(i,j)\in P_{l}}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})~{}~{}~{}\text{and}~{}~{}~{}\mathbb{E}_{|I_{k}|}\psi_{D_{z}}(\hat{\vartheta}_{k})=|I_{k}|^{-1}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}),$
on $P_{l}$ and $I_{k}$, respectively, where $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ are correspondingly estimated on $P_{l}^{c}$ and $I_{k}^{c}$. Then, define
$\overline{\mathbb{U}}_{L}g_{h}^{s}(\widehat{\theta})=L^{-1}\sum_{l=1}^{L}\mathbb{U}_{|P_{l}|}g_{h}^{s}(\widehat{\theta}_{l})~{}~{}~{}\text{and}~{}~{}~{}\overline{\mathbb{E}}_{K}\psi_{D_{z}}(\widehat{\vartheta})=K^{-1}\sum_{k=1}^{K}\mathbb{E}_{|I_{k}|}\psi_{D_{z}}(\widehat{\vartheta}_{k}),$
and the estimator for the PCEs can be formed as
$\widehat{\tau}_{h}^{10}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{10}(\widehat{\theta})}{[\overline{\mathbb{E}}_{K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{00}(\widehat{\theta})}{\{\overline{\mathbb{E}}_{K}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{11}(\widehat{\theta})}{\{\overline{\mathbb{E}}_{K}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$
which are the type I debiased machine learning estimators. We focus on the type II estimator because it is generally recommended for its numerical stability in pooling intermediate estimates based on split data, especially in smaller samples.
We now discuss the asymptotic properties of the debiased machine learning estimators in (15). Assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability. In addition, we assume true nuisance realizations $\theta^{*},\vartheta^{*}\in\Theta^{*}$. The following result summarizes the efficiency properties of (15).
Theorem 10 (Efficiency under nonparametric working models).
Under Assumptions 13, and
  • (a)
    $\mathbb{E}|g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})|^{2}<\infty$ and $\mathbb{E}|\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})|^{2}<\infty$;
  • (b)
[…]
  • (c)
[…]
  • (d)
    there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 7.
Theorem 10 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al., (2022, Lemma 15) and Escanciano and Terschuur, (2023, Lemma 4). To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$. Regulatory conditions in Theorem 10 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners. Conditions (c) and (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al.,, 2018). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (15) via nonparametric bootstrap.
5 Simulation studies
We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 6, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15). In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
5.1 The probability index estimands with an intermediate variable
[…]
$D_{i}(Z_{i})|Z_{i},X_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\},$
$Y_{i}(Z_{i})|D_{i},Z_{i},X_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1),$
for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$. In this simulation study, we consider a binary intermediate variable ($D$ can be considered as the treatment receipt in the context of non-compliance), and assume the interest lies in estimating all three principal probability index estimands ($\tau_{h}^{10}$, $\tau_{h}^{00}$ and $\tau_{h}^{11}$ with $h(u,v)=\mathbbm{1}(u\geq v)$). In addition, we create transformed covariates, $\widetilde{X}_{i}=(\widetilde{X}_{i1},\widetilde{X}_{i2},\widetilde{X}_{i3},\widetilde{X}_{i4})^{\prime}$, where
$\widetilde{X}_{i1}=\exp(0.5X_{i1}),~{}\widetilde{X}_{i2}=\frac{X_{i2}}{1+X_{i1}},~{}\widetilde{X}_{i3}=\left(\frac{X_{i2}X_{i3}}{25}+0.6\right)^{3},~{}\widetilde{X}_{i4}=\frac{X_{i3}X_{i4}}{\sqrt{2}},$
to simulate scenarios where the nuisance functions are misspecified. Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise. Also, it is worth noting that, in practice, model misspecification pertains to the flexible debiased machine learning estimators, because their estimation accuracy could be affected by the quality of the input covariate feature matrix.
For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 4. For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries. We set the number of observations $n=2,000$ and consider $1,000$ replications; violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure 1 in Section S6 of the Online Supplement.
Figure 1 presents the results of the five estimators under comparison in scenarios where, at most, one of the three nuisance functions is misspecified. Plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified, while the triply robust and debiased machine learning estimators are consistent across all four scenarios (correct specification of all nuisance functions or subsets of nuisance functions). Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure 1 in Section S6 of the Online Supplement. As expected, the plug-in moment estimators are generally biased in those scenarios. In comparison to the plug-in moment estimators, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy. Finally, compared to the triply robust estimator, the debiased machine learning estimator leads to a slight improvement and is generally less biased.
(a) $\tau_{h}^{10}$
(b) $\tau_{h}^{00}$
(c) $\tau_{h}^{11}$
Figure 1: Violin plots of estimators for the principal probability index estimand in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). Horizontal dashed lines represent the true values of the PGCEs.
5.2 The win-ratio estimands with an intermediate variable
In the second simulation study, we consider the case of the win-ratio estimands introduced in Examples 2 and 5, with the final potential outcomes being ordinal. We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,\ldots,Q\}$, we consider the following three-category ordinal logistic regression model:
$\log\frac{\mathbb{P}(Y_{i}(Z_{i})\leq q)}{\mathbb{P}(Y_{i}(Z_{i})>q)}=\eta_{q,0}+2Z_{i}-D_{i}+X_{i1}-X_{i2}+1.2X_{i3}-0.8X_{i4},$
for $q\in\{1,2\}$, where $\eta_{1,0}=-1$ and $\eta_{2,0}=1$. Transformed covariates are created following the procedure described in the previous simulation study. For the plug-in moment estimators and the triply robust estimator with parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are similarly estimated via logistic regression. For the debiased machine learning estimator, we used the multinomial generalized linear models with the elastic net penalty from the glmnet package in R to estimate the pairwise outcome mean, and the treatment propensity and principal scores are estimated using the SuperLearner package. The debiased machine learning estimator adopts a five-fold cross-fitting ($K=5$ and $L=15$) procedure. We set the number of observations $n=2,000$ with $1,000$ replications. Violin plots summarizing the simulation results of each considered estimator are given in Figure 2 and Web Figure 2 in Section S6 of the Online Supplement.
From Figure 2, we observe that, similar to the estimation of the probability index estimand in Section 5.1, the plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified, while the triply robust and debiased machine learning estimators are consistent in all four scenarios where at most one of the three nuisance functions is misspecified. When at least two of the three nuisance functions are misspecified (shown in Web Figure 2 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the rest in terms of estimation accuracy (bias).
(a) $\tau_{h}^{10}$
(b) $\tau_{h}^{00}$
(c) $\tau_{h}^{11}$
Figure 2: Violin plots of estimators for the win-ratio estimands in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). Horizontal dashed lines represent the true values of the PGCEs.
6 Data example
We illustrate the proposed methods with a data example from the U.S. Jobs Corps study (Schochet et al.,, 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24. The dataset (named ‘JC’ with $9,240$ observations, available in the R package causalweight) was from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either, the treatment group, i.e., receiving an offer to participate in Job Corps immediately, or, the control group, i.e., not participating in Job Corps until three years later. Follow-up surveys were administered to both groups 24, 36, and 48 months after the random assignment to collect data on the outcomes of interest. As in many open-label randomized controlled trials, substantial noncompliance to treatment assignment as an intermediate outcome occurred 1 year after randomization (15.4% [$857/5,577$] of participants assigned to treatment did not adhere to the assignment, and 50.6% [$1,854/3,663$] assigned to control switched to the treatment condition).
The proposed methods were implemented to examine the causal effect of Job Corps intervention, where the weekly earnings one year after the assignment were selected as the final outcome. As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects. In each data analysis below, we adjust for baseline covariates including baseline characteristics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
6.1 The survivor probability index for continuous outcomes
We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$. In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al., (2009). Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps. For methods based on parametric working models, a logistic model is considered for the employment process within each assignment group, and a log-normal model is fitted for the final earning outcome to induce the pairwise mean outcome function. For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the Super Learner with the random forest and generalized linear models libraries.
The estimated results for SPI are given in Table 1, where $1,000$ bootstrap replicates were obtained for estimating the standard errors. We observe that, for the always-employed (72.9% of all participants by estimation), the probability index is estimated as $0.539$ using the triply robust and the debiased machine learning methods. This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-survivor under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment. Finally, the point estimates under the plug-in moment methods are similar to the triply robust and debiased machine learning methods, which likely suggests that each parametric working model may be approximately correctly specified.
Table 1: Estimated SPI for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). The corresponding 95% confidence interval is given in the following parenthesis.
Methods Survivor proability index estimates (standard errors)
M1 0.521 (0.508, 0.533)
M2 0.518 (0.505, 0.530)
M3 0.525 (0.513, 0.538)
TR 0.539 (0.524, 0.555)
DML 0.539 (0.527, 0.552)
6.2 Principal win-ratio for ordinal outcomes
We provide additional analysis of the Job Corps data to illustrate the application of the proposed methods to address ordinal outcomes with noncompliance. In this example, we use the actual treatment status as an intermediate variable, similar to Chen and Flores, (2015). Under the monotonicity assumption, three strata potentially exist, including the compliers (stratum $10$), the never-takers (stratum $00$), and the always-takers (stratum $11$). To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu et al., (2018). Under this setting, we estimate the principal win-ratios within each principal strata similar to the definition in (3). For methods based on parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are estimated via logistic regression. For the debiased machine learning approach, we used the multinomial generalized linear models with the elastic net penalty to estimate the pairwise outcome mean, and the treatment propensity and principal scores are estimated using Super Learning; the estimator is fitted under a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
The estimated results are given in Table 2, where $1,000$ bootstrap replicates were obtained for estimating the standard errors. The estimated proportions of always-takers, compliers, and never-takers in all participants are 50.1%, 34.0%, and 15.9%, respectively. The Jobs Corps intervention assignment positively affects participants’ earnings one year after the assignment for compliers and always-takers, and the effect appears to be more pronounced for compliers. For example, using the debiased machine learning estimator, there is 27% increase in the chances of being in a higher earning category for a complier under treatment than a complier under control. The 95% lower confidence limits obtained from all five methods are strictly above the null value of 1. For the never-takers, however, the effect of Jobs Corps appears more neutral, as the principal win-ratio is estimated to be closer to 1 and three out of five methods (M1, triply robust and debiased machine learning) produce 95% confidence intervals that include the null value of 1. Interestingly, in this reanalysis using a win-ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable. In particular, the statistically significant effect among the always-takers is indicative of a positive psychological effect. That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence having a large chance to move into a higher earning category) compared to those who switched from their initial control assignment.
Table 2: Estimated win-ratios within different compliance strata for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). The corresponding 95% confidence interval is given in the following parenthesis.
Strata
Methods $10$ $00$ $11$
M1 1.216 (1.128, 1.304) 1.096 (0.955, 1.238) 1.117 (1.016, 1.218)
M2 1.206 (1.122, 1.289) 1.087 (1.025, 1.149) 1.096 (1.025, 1.168)
M3 1.204 (1.121, 1.286) 1.089 (1.025, 1.152) 1.096 (1.025, 1.167)
TR 1.197 (1.112, 1.282) 1.026 (0.920, 1.133) 1.131 (1.040, 1.223)
DML 1.270 (1.146, 1.393) 1.041 (0.940, 1.142) 1.167 (1.082, 1.252)
7 Discussion
In this article, we developed principal stratification methods for addressing a general class of PGCE estimands defined with a binary intermediate variable. Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability that are based on the contrast function, we contributed a semiparametric efficient approach for estimating the PGCE estimands by identifying their efficient influence functions and obtaining the ratio $U$-estimators that combine several different nuisance working functions. We showed that the $U$-estimators are triply robust in that they are consistent and asymptotically normal when two of the three nuisance functions are correctly specified. The $U$-estimators become locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric models or by machine learning models with cross-fitting. In light of the increasing interest in pairwise comparison estimands in studies with a composite endpoint (Buyse,, 2010; Bebu and Lachin,, 2015) in biomedical research, our methods can also be useful when there is a need to additionally address a binary intercurrent event, such as treatment noncompliance or death (Lipkovich et al.,, 2022; Mao,, 2024)
The estimands and efficient $U$-estimators were developed to address a binary treatment and a binary intermediate outcome specifically but without restrictions on the final outcome. Although this is a standard setup for principal stratification analysis, additional work is needed to expand the identification conditions and results to address extensions of the PGCE estimands with a continuous treatment and/or a continuous intermediate variable (Antonelli et al.,, 2023). With continuous treatments, the definition of pairwise comparison estimands becomes non-trivial as there are infinitely many contrast functions (based on an infinitesimal increment of the treatment value) that one can specify to define $\tau_{h}$ even in the absence of an intermediate variable. On the other hand, with a continuous intermediate variable, there are infinitely many principal strata that can lead to conceptual and implementation challenges (Schwartz et al.,, 2011). Focusing on a linear contrast function, Lu et al., (2023) have developed nonparametric identification and semiparametric efficient estimators for principal causal effects with a continuous intermediate variable, which may inspire future theoretical development that allows for a more general nonlinear contrast function in the estimands definition.
Finally, as in all principal stratification analyses that achieve point identification, our $U$-estimators for estimating PGCEs depend on the monotonicity and the mean principal ignorability assumptions, that are unverifiable from the observed data alone. In Section S7 of the Online Supplement, we have described a sensitivity function approach to obtain the bias-corrected $U$-estimators for estimating the PGCEs, when either the monotonicity assumption or the mean principal ignorability assumption is violated. The bias-corrected $U$-estimators remain triply robust to potential working model misspecifications and provide means to generate a range of point estimates under assumed departures from the key identification assumptions.

### v2

@@
 The following two examples illustrate the PGCE with a specific contrast function.
-Example 1 (Probability index under truncation by death).
+Example 2.1 (Probability index under truncation by death).
 When $D$ represents the survival status (alive if $D=1$) that is ascertained before the measurement of the final outcome $Y$ (such as quality of life, which is only well-defined when $D=1$), there exist four principal strata based on the potential survival status.
@@
 In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg,, 1971).
-Example 2 (Win estimands under noncompliance).
+Example 2.2 (Win estimands under noncompliance).
 In experimental studies with ordinal outcomes, Wang and Pocock, (2016) introduced the win-ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\allowbreak\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$.
@@
 That is, with Assumption 3, each one of the subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only.
-Remark 1.
+Remark 2.1.
 […]
@@
 $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\},$
-for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. We provide some examples of the pairwise outcome model models for specific contrast functions below.
-Example 3 (Pairwise mean function for average treatment effect).
+for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. We provide some examples of the pairwise outcome mean models for specific contrast functions below.
+Example 3.1 (Pairwise mean function for average treatment effect).
 For studying the average treatment effect, the contrast function is specified as $h(u,v)=u-v$.
-If the potential outcomes are continuous, a Gaussian regression model is often assumed for the potential outcomes, i.e., $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$.Here, $f_{z}(D,X)$ is the mean function of $D$ and $X$ given treatment status $Z=z$.
+If the potential outcomes are continuous, a Gaussian regression model is often assumed for the potential outcomes, i.e., $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$.
+Here, $f_{z}(D,X)$ is the mean function of $D$ and $X$ given treatment status $Z=z$.
 We then have the pairwise outcome mean function given by
 $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2}).$ (7)
-Example 4 (Pairwise mean function for probability index).
+Example 3.2 (Pairwise mean function for probability index).
 For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$.
-If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3.
+If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3.1.
 In this case, following Wolfe and Hogg, (1971), one can derive the pairwise outcome mean as
 $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\Phi\left\{\frac{f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2})}{\sqrt{2}\sigma}\right\}.$ (8)
-Example 5 (Pairwise outcome mean for win estimands).
-Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand).
+Example 3.3 (Pairwise outcome mean for win estimands).
+Consider the win estimands developed in Example 2.2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand).
 For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao, (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
@@
 We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$.
-These elements then prepare us to state the set of nonparametric identification results in the following Theorem 6.
-Theorem 6 (Nonparametric identification).
+These elements then prepare us to state the set of nonparametric identification results in the following Theorem 3.4.
+Theorem 3.4 (Nonparametric identification).
 Suppose Assumptions 1-3 hold, $e_{s}>0$ for $s\in\{10,00,11\}$, and $10<\pi(X)<$ 1 for all $X\in\mathcal{X}$, the support of covariates.
@@
     $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{p_{0}(X_{1})}{p_{0}}\right\}\left\{\frac{p_{0}(X_{2})}{p_{0}}\right\}\mu_{1101}(X_{1},X_{2})\right].$
-We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 6, and the identification of the other two PGCEs follows the same strategy.
-In Theorem 6(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$.
+We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 3.4, and the identification of the other two PGCEs follows the same strategy.
+In Theorem 3.4(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$.
 Specifically, under Assumptions 1 and 2, we have
@@
 This ensures the identification of $\tau_{h}^{10}$.
-In Theorem 6(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
+In Theorem 3.4(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
 $\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
@@
 Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$.
-Finally, in Theorem 6(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 6(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$.
+Finally, in Theorem 3.4(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 3.4(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$.
 The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.
-Remark 2.
-Theorem 6 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al.,, 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al., (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
+Remark 3.1.
+Theorem 3.4 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al.,, 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al., (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
 3.2 Estimators based on nonparametric identification formulas
@@
 Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart,, 1998).
-For example, Theorem 6(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
+For example, Theorem 3.4(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
 $\widehat{\tau}_{h,\pi+p}^{10}=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{1}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right].$
-Theorem 6(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
+Theorem 3.4(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
 $\widehat{\tau}_{h,\pi+\mu}^{10}$ $=\mathbb{U}_{n}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1};\widehat{\alpha})-D_{1}(1-Z_{1})/\{1-\pi(X_{1};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\right.$
 ×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2};\widehat{\alpha})-D_{2}(1-Z_{2})/\{1-\pi(X_{2};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right),$
-and Theorem 6(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
+and Theorem 3.4(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
 $\widehat{\tau}_{h,p+\mu}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{p_{1}(X_{1};\widehat{\beta})-p_{0}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\left\{\frac{p_{1}(X_{2};\widehat{\beta})-p_{0}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right].$
@@
 Specifically, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment probability and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
-Remark 3.
+Remark 3.2.
 The weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score, which is an extension of the inverse propensity score weighted estimator in Mao, (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable.
@@
 Then, the following results are obtained.
-Theorem 7 (EIF).
+Theorem 4.1 (EIF).
 Under Assumptions 1-3, (a) the EIF for $\tau_{h}^{10}$ is $\varphi_{h}^{10}(\mathcal{O})=\{\varphi_{h,N}^{10}(\mathcal{O})-\tau_{h}^{10}\varphi_{D}^{10}(\mathcal{O})\}/(p_{1}-p_{0})^{2}$, where $\varphi_{h,N}^{10}(\mathcal{O})=\phi_{h}^{10}(\mathcal{O})-2\tau_{h,N}^{10}$ and
@@
 $\varphi_{D}^{11}(\mathcal{O})=2\psi_{D_{0}}(\mathcal{O})p_{0}-2\tau_{D}^{11}.$
-The EIFs in Theorem 7 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey, (2022) and Hines et al., (2022) with respect to the distribution of observed data, $\mathcal{O}$.
+The EIFs in Theorem 4.1 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey, (2022) and Hines et al., (2022) with respect to the distribution of observed data, $\mathcal{O}$.
 The derivation for $\varphi_{D}^{s}(\mathcal{O})$ is then follows the chain rule.
@@
 Importantly, the EIF for the PGCE remains unchanged after this re-expression, since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged.
-Based on Theorem 7, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
-Remark 4.
+Based on Theorem 4.1, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
+Remark 4.1.
 Without considering the intermediate variable, the EIF for the generalized causal effect estimand, $\tau_{h}$ in (1), $\varphi_{h}(\mathcal{O})$, can be derived using the same approach as $\varphi_{h}(\mathcal{O})=\phi_{h}(\mathcal{O})-2\tau_{h}$, where
@@
 4.2 Multiply robust estimators
-The EIFs given in Theorem 7 imply the semiparametric efficiency lower bound, and pave the way for formulating estimators for the PGCEs.
+The EIFs given in Theorem 4.1 imply the semiparametric efficiency lower bound, and pave the way for formulating estimators for the PGCEs.
 However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator.
-We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 7, thereby achieving semiparametric efficiency with correctly specified nuisance functions.
+We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 4.1, thereby achieving semiparametric efficiency with correctly specified nuisance functions.
 To this end, we first notice that each PGCE can be rewritten as the following:
@@
 $\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (14)
-which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 8.
-Theorem 8 (Multiply robustness and efficiency under parametric working models).
-Under Assumptions 13, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$.
+which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 4.2.
+Theorem 4.2 (Multiply robustness and efficiency under parametric working models).
+Under Assumptions 1-3, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$.
 Then, the estimator $\widehat{\tau}_{h}^{s}$ in (14) is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
-Under Theorem 8, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified, and is semiparametrically efficient if all three nuisance models are correctly specified.
+Under Theorem 4.2, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified, and is semiparametrically efficient if all three nuisance models are correctly specified.
 The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator.
@@
 The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao, (2017) and Theorem 3 of Jiang et al., (2022).
-As discussed in Remark 3, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator.
+As discussed in Remark 3.2, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator.
 Details of the proof are provided in Section S4.2 of the Online Supplement.
@@
 The following example is used to provide a demonstration of $\mathcal{I}$ and $\mathcal{P}$.
-Example 9 (Sample splitting for $U$-statistics).
+Example 4.3 (Sample splitting for $U$-statistics).
 Consider a scenario where $n=10$ and $K=3$.
@@
 $\mathbb{E}_{n,K}\psi_{D_{z}}(\widehat{\vartheta})=n^{-1}\sum_{k=1}^{K}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}).$
-Remark 5.
+Remark 4.2.
 Estimators given in (15) are the type II estimator proposed in Chernozhukov et al., (2018) for estimands without an intermediate variable. Alternatively, one could first obtain separate estimates
@@
 The following result summarizes the efficiency properties of (15).
-Theorem 10 (Efficiency under nonparametric working models).
-Under Assumptions 13, and
+Theorem 4.4 (Efficiency under nonparametric working models).
+Under Assumptions 1-3, and
   • (a)
@@
     there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
-then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 7.
-Theorem 10 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al., (2022, Lemma 15) and Escanciano and Terschuur, (2023, Lemma 4).
+then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 4.1.
+Theorem 4.4 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al., (2022, Lemma 15) and Escanciano and Terschuur, (2023, Lemma 4).
 To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$.
-Regulatory conditions in Theorem 10 are mostly mild.
+Regulatory conditions in Theorem 4.4 are mostly mild.
 Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$.
@@
 5 Simulation studies
-We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 6, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15).
+We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 3.4, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15).
 In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
@@
 Also, it is worth noting that, in practice, model misspecification pertains to the flexible debiased machine learning estimators, because their estimation accuracy could be affected by the quality of the input covariate feature matrix.
-For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 4.
+For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 3.2.
 For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries.
@@
 5.2 The win-ratio estimands with an intermediate variable
-In the second simulation study, we consider the case of the win-ratio estimands introduced in Examples 2 and 5, with the final potential outcomes being ordinal.
+In the second simulation study, we consider the case of the win-ratio estimands introduced in Examples 2.2 and 3.3, with the final potential outcomes being ordinal.
 We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study.
@@
 We observe that, for the always-employed (72.9% of all participants by estimation), the probability index is estimated as $0.539$ using the triply robust and the debiased machine learning methods.
-This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-survivor under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment.
+This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment.
 Finally, the point estimates under the plug-in moment methods are similar to the triply robust and debiased machine learning methods, which likely suggests that each parametric working model may be approximately correctly specified.

### v3

@@
 We thus introduce the principal generalized causal effect estimands, which extend the principal average causal effects to accommodate nonlinear contrast functions.
-Under principal ignorability, we expand the theoretical results in Jiang et al., (2022) to a much wider class of causal estimands in the presence of a binary intermediate variable.
+Under principal ignorability, we expand the theoretical results in Jiang et al.
+(2022) to a much wider class of causal estimands in the presence of a binary intermediate variable.
 We develop identification formulas and derive the efficient influence functions of the generalized estimands for principal stratification analyses.
@@
 The proposed methods are illustrated through simulations and the analysis of a data example.
-Keywords: causal inference, efficient influence function, principal stratification, multiply robust estimation, $U$-statistics, win-ratio.
+Keywords: Causal inference, efficient influence function, principal stratification, multiply robust estimation, $U$-statistics, win ratio
 1 Introduction
-For causal inference under the Neyman-Rubin model (Rubin,, 1974), the average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest.
+For causal inference under the Neyman-Rubin model (Rubin, 1974), the average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest.
 Specifically, let $\{Y(1),Y(0)\}$ denote a pair of potential outcomes for a randomly selected individual from a super-population, and then the average causal effect estimand is defined as $\mathbb{E}\{Y(1)-Y(0)\}$.
@@
 That is, given a contrast function $h(u,v)$ defined on the product space, $\mathcal{Y}\times\mathcal{Y}$ ($\mathcal{Y}$: the potential outcome space), a generalized estimand can be defined as $\mathbb{E}[h\{Y(1),Y(0)\}]$.
-Examples include the average causal effect when $h(u,v)=u-v$, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$ (Lu et al.,, 2020).
-Because of the fundamental problem of causal inference, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$.
-However, a closely related quantity, the generalized causal effect (O’brien,, 1988; Mao,, 2017), can be formulated as:
+Examples include the average causal effect when $h(u,v)=u-v$, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$ (Lu et al., 2020).
+Because $Y(1)$ and $Y(0)$ are never observed simultaneously, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$.
+However, a closely related quantity, the generalized causal effect (O’brien, 1988; Mao, 2017), can be formulated as:
 $\tau_{h}=\int\int h(u,v)\mathrm{d}F_{1}(u)\mathrm{d}F_{0}(v),$ (1)
 where $F_{z}(\cdot)$ is the marginal distribution of the potential outcome $Y(z)\in\mathcal{Y}$ for $z=0,1$.
-Consider $\{Y_{1}(1),Y_{1}(0)\}$ and $\{Y_{2}(1),Y_{2}(0)\}$ as two independent replicates of $\{Y(1),Y(0)\}$, and then $\tau_{h}$ can be re-expressed in a familiar form, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}]$.
-Two notable examples are the probability index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al.,, 2015), and the win-ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al.,, 2011).
+Let $\{Y_{1}(1),Y_{1}(0)\}$ and $\{Y_{2}(1),Y_{2}(0)\}$ denote two independent replicates of $\{Y(1),Y(0)\}$, and $\tau_{h}$ can be re-expressed in a familiar form, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}]$.
+Two notable examples are the probability index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al., 2015), and the win ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al., 2011).
 These estimable quantities are alternative measurements for the population causal effect and can be straightforwardly interpreted.
-For example, the probability index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than $F_{0}$, and is robust against extreme values in the tails of the potential outcome distributions.
-When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al.,, 1996), mortality event (Zhang et al.,, 2009) or dropout (Frangakis and Rubin,, 1999), often have important implications even on the definition of a causal effect and require careful adjustment.
-The intermediate variable is not guaranteed to be balanced via randomization, and direct conditioning on the intermediate variable leads to sample selection bias and non-causal comparisons even in randomized experiments.
-As a solution, the principal stratification framework (Frangakis and Rubin,, 2002) segments the population into subpopulations (principal strata) according to joint potential values of the intermediate variable under alternative conditions.
-The stratum-specific causal effect—the principal causal effect—can be defined and is causally interpretable.
+For example, the probability index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than $F_{0}$ and is robust against extreme values in the tails of the potential outcome distributions.
+When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al., 1996), mortality event (Zhang et al., 2009) or dropout (Frangakis and Rubin, 1999), often have important implications even on the definition of a causal effect and require careful adjustment.
+The intermediate variable is not guaranteed to be balanced via randomization, and direct conditioning on the intermediate variable leads to sample selection bias and non-causal comparisons, even in randomized experiments.
+As a solution, the principal stratification framework (Frangakis and Rubin, 2002) segments the population into principal strata (subpopulations) according to joint potential values of the intermediate variable under alternative conditions.
+The stratum-specific causal effect—the principal causal effect—can thus be defined and is causally interpretable.
 However, a predominant focus in prior works has been on the average causal effect within each principal stratum, and there is a lack of related development when the generalized causal effect is of interest within each subpopulation.
 Therefore, in this article, we propose a new class of estimands, the principal generalized causal effect (PGCE), which accommodates an arbitrary and potentially nonlinear contrast function.
-Under a linear contrast function, the PGCE reduces to the usual principal average causal effect.
-As we demonstrate in due course, the PGCE can represent the stratum-specific probability index and strata-specific win-ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within important subpopulations.
-Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al.,, 1996) and principal ignorability (Jo and Stuart,, 2009; Ding and Lu,, 2016), which are typically invoked for principal stratification analyses.
+The PGCE, under a linear contrast function, reduces to the usual principal average causal effect.
+As we demonstrate in due course, the PGCE can represent the stratum-specific probability index and strata-specific win ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within subpopulations of interest.
+Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al., 1996) and principal ignorability (Jo and Stuart, 2009; Ding and Lu, 2016), which are typically invoked for principal stratification analyses.
 The monotonicity assumption states that the treatment has a nonnegative causal effect on the intermediate outcome of each unit, effectively restricting the number of potential strata.
-This condition is considered plausible when, for example, the intermediate variable is death, and the treatment is known to benefit survival or when the intermediate variable is the actual treatment received and the treatment does not discourage treatment uptake.
+This condition is considered plausible when, for example, the intermediate variable is death, and the treatment is known to benefit survival, or when the intermediate variable is the actual treatment received and the treatment does not discourage treatment uptake.
 The principal ignorability assumption, on the other hand, requires that the observed pre-treatment covariates are sufficient to control for confounding between the intermediate variable and the final outcome.
-For the principal average causal effect, under monotonicity and principal ignorability, Ding and Lu, (2016) formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang et al., (2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and further developed multiply robust and locally efficient estimators under the additional assumption of treatment unconfoundedness.
+For the principal average causal effect, under monotonicity and principal ignorability, Ding and Lu (2016) formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang et al.
+(2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and further developed multiply robust and locally efficient estimators under the additional assumption of treatment unconfoundedness.
 In contrast, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
-The contributions of this article are several-folded.
-First, we propose a class of PGCE estimands in the presence of an intermediate variable, and develop complementary nonparametric identification formulas under treatment unconfoundedness, monotonicity, and principal ignorability.
-Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands, and seek efficiency improvement by deriving their corresponding efficient influence functions (Bickel et al.,, 1993, EIFs).
-Specifically, we show that, when the contrast function $h(u,v)=u-v$, our EIFs reduce to those derived by Jiang et al., (2022) as a special case, but are otherwise more general to accommodate arbitrary contrast functions.
-A major complexity in the context of a nonlinear contrast function, however, is that the estimating equations based on the EIFs requires nuisance functions defined over pairs of observations.
+The contributions of this article are several-fold.
+First, we propose a class of PGCE estimands in the presence of an intermediate variable and develop complementary nonparametric identification formulas under treatment unconfoundedness, monotonicity, and principal ignorability.
+Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands and seek efficiency improvement by deriving their corresponding efficient influence functions (Bickel et al., 1993, EIFs).
+Specifically, we show that when the contrast function $h(u,v)=u-v$, our EIFs reduce to those derived by Jiang et al.
+(2022) as a special case but are otherwise more general to accommodate arbitrary contrast functions.
+A major complexity in the context of a nonlinear contrast function, however, is that the estimating equations based on the EIFs require nuisance functions defined over pairs of observations.
 Thus, to obtain a feasible estimator, we define ratio $U$-statistics whose influence functions are the same as the EIFs.
@@
 When all nuisance functions are correctly specified, the $U$-estimators are locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds.
-The asymptotic properties of the $U$-estimators are developed upon the theory of empirical process (van der Vaart,, 1998, §19), $U$-process (Arcones and Gine,, 1993), as well as the theory of $U$-statistics (van der Vaart,, 1998, §11-12).
+The asymptotic properties of the $U$-estimators are developed upon the theory of empirical process (van der Vaart, 1998, §19), $U$-process (Arcones and Gine, 1993), as well as the theory of $U$-statistics (van der Vaart, 1998, §11-12).
 Finally, the construction of the multiply robust $U$-estimators facilitates the integration of machine learning estimators for the nuisance functions.
-To achieve full asymptotic efficiency, we learn each nuisance function using machine learning methods, and operationalize the debiased machine learning $U$-estimators via cross-fitting (Chernozhukov et al.,, 2018).
-We discuss the modification needed for cross-fitting when a nonlinear contrast function is considered when defining the PGCE.
-We prove the multiple robustness and efficiency property of the debiased machine learning estimators under mild regulation conditions (Chernozhukov et al.,, 2022; Escanciano and Terschuur,, 2023).
+To achieve full asymptotic efficiency, we learn each nuisance function using machine learning methods and operationalize the debiased machine learning $U$-estimators via cross-fitting (Chernozhukov et al., 2018).
+We discuss the modification needed for cross-fitting if a nonlinear contrast function is considered when defining the PGCE and prove the multiple robustness and efficiency property of the debiased machine learning estimators under mild regulation conditions (Chernozhukov et al., 2022; Escanciano and Terschuur, 2023).
 The rest of this article is organized as follows.
@@
 Section 3 discusses the nonparametric identification formulas for the PGCE and associated moment estimators.
-Section 4 derives the EIF for each PGCE estimand, and develops the EIF-motivated feasible $U$-estimators and the debiased machine learning estimators.
+Section 4 derives the EIF for each PGCE estimand and develops the EIF-motivated feasible $U$-estimators and the debiased machine learning estimators.
 Section 5 demonstrates the finite-sample performance of the proposed estimators via simulations.
-Section 6 illustrates the proposed methods with a data example from the United States Jobs Corps study (Schochet et al.,, 2001), and Section 7 concludes.
-Detailed proofs of all Theorems are available in the Online Supplement.
+Section 6 illustrates the proposed methods with a data example from the United States Jobs Corps study (Schochet et al., 2001), and Section 7 concludes.
+Detailed proofs of all theoretical results are available in the Online Supplement.
 2 Notation, estimands and assumptions
 2.1 Setting and notation
-We consider a study (experimental or observational) with $n$ participants and two treatment conditions, and let $Z_{i}$ denote the treatment assignment of participant $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 00 the control.
+We consider a study (experimental or observational) with $n$ participants and two treatment conditions and let $Z_{i}$ denote the treatment assignment of participant $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 00 the control.
 We use $D_{i}(z)$ and $Y_{i}(z)$ to denote the potential outcomes of the intermediate variable and the final outcome variable of interest for participant $i$ under treatment $Z_{i}=z$ ($z\in\{0,1\}$), respectively.
 The intermediate variable $D_{i}\in\{0,1\}$ is assumed to be binary, but there is no restriction on the type of variable for $Y_{i}$.
-Under the stable unit treatment value assumption (Rubin,, 1980, SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$.
+Under the stable unit treatment value assumption (Rubin, 1980, SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$.
 Let $X_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{X_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d).
 2.2 Principal generalized causal effects
-We operate under the principal stratification framework (Frangakis and Rubin,, 2002), and define distinct principal strata according to the joint potential values of the intermediate variable under alternative conditions.
+We operate under the principal stratification framework (Frangakis and Rubin, 2002), and define distinct principal strata according to the joint potential values of the intermediate variable under alternative conditions.
 Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{11,10,01,00\}$.
-A key insight under the principal stratification framework is that, although $D$ is a post-randomization variable, the strata variable $S$ is unaffected by the assignment and can be viewed as a pre-treatment variable.
-Therefore, based on any contrast function $h$, we define the principal generalized causal effect estimand as
+A key insight under the principal stratification framework is that although $D$ is a post-randomization variable, the stratum variable $S$ is unaffected by the assignment and can be viewed as a pre-treatment variable.
+Therefore, based on any contrast function $h$, we define the principal generalized causal effect (PGCE) estimand as
 $\tau_{h}^{s}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=S_{2}=s],~{}\text{for}~{}s\in\{11,10,01,00\},$ (2)
@@
 If the two units are selected from different principal strata, the resulting quantity, $\tau_{h}^{s_{1},s_{2}}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2}]$ with $s_{1}\neq s_{2}$, remains estimable under this framework but generally lacks a meaningful causal interpretation; we thus do not further pursue this direction.
-Evidently, when a linear contrast function is specified as $h(u,v)=u-v$, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature; see, for example, Frangakis and Rubin, (2002), Jiang et al., (2022) and Lipkovich et al., (2022).
+Evidently, when a linear contrast function is specified as $h(u,v)=u-v$, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature; see, for example, Frangakis and Rubin (2002), Jiang et al.
+(2022) and Lipkovich et al.
+(2022).
 When a nonlinear contrast function is considered in defining the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum.
 The following two examples illustrate the PGCE with a specific contrast function.
-Example 2.1 (Probability index under truncation by death).
+Example 1 (Probability index under truncation by death)
 When $D$ represents the survival status (alive if $D=1$) that is ascertained before the measurement of the final outcome $Y$ (such as quality of life, which is only well-defined when $D=1$), there exist four principal strata based on the potential survival status.
 That is, $s=11$ represents the always-survivors who would survive until the measurement of the final outcome regardless of the assignment; $s=10$ and $s=01$ represent the protected and harmed, who would survive until the measurement of the final outcome only under the treatment condition and control condition, respectively; $s=00$ represents the never-survivors who would not survive until the measurement of the final outcome regardless of the assignment.
-Because the final outcome is oftentimes only well-defined without ambiguity when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al.,, 2009; Chen et al.,, 2024).
-In clinical research, Acion et al., (2006) argued that a more intuitive effect size can be measured by the probability index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$.
+Because the final outcome is oftentimes only well-defined without ambiguity when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al., 2009; Chen et al., 2024).
+In clinical research, Acion et al.
+(2006) argued that a more intuitive effect size can be measured by the probability index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$.
 Based on this contrast function, we can define the survivor probability index (SPI) as
 $\mathrm{SPI}=\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=11\}.$ (3)
-Compared to the SACE, the SPI is less sensitive to skewness of the underlying distributions of the potential outcomes and heteroskedasticity between conditions and outliers.
-In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg,, 1971).
-Example 2.2 (Win estimands under noncompliance).
-In experimental studies with ordinal outcomes, Wang and Pocock, (2016) introduced the win-ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\allowbreak\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$.
+Compared to the SACE, the SPI is less sensitive to the skewness of the underlying distributions of the potential outcomes and heteroskedasticity between conditions and outliers.
+In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg, 1971).
+Example 2 (Win estimands under noncompliance)
+In experimental studies with ordinal outcomes, Wang and Pocock (2016) introduced the win ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$.
 This estimand corresponds to a two-dimensional contrast function specified as $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$.
-If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{11,10,01,00\}$ refers to always-takers, compliers, defiers, and never-takers, respectively (Angrist et al.,, 1996).
-Therefore, the win-ratio estimands can be extended to define the compliers win-ratio (CWR) as
+If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{11,10,01,00\}$ refers to always-takers, compliers, defiers, and never-takers, respectively (Angrist et al., 1996).
+Therefore, the win ratio estimands can be extended to define the compliers win ratio (CWR) as
 $\mathrm{CWR}=\frac{\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=S_{2}=10\}}{\mathbb{P}\{Y_{1}(1)<Y_{2}(0)|S_{1}=S_{2}=10\}}.$ (4)
 The CWR estimand represents the odds of a randomly selected unit from all-treated compliers having a more favorable outcome than one from the all-control compliers subpopulation.
-Alternatively, the complier win-difference (CWD) or complier net benefit estimand on the difference scale is given as
+Alternatively, the complier win difference (CWD) or complier net benefit estimand on the difference scale is given as
 $\mathrm{CWD}=\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=S_{2}=10\}-\mathbb{P}\{Y_{1}(1)<Y_{2}(0)|S_{1}=S_{2}=10\}.$ (5)
-The win estimands within other principal strata can be defined similarly. Mao, (2024) considered the point and interval identification of a similar estimand using the instrumental variable approach.
+The win estimands within other principal strata can be defined similarly.
+In the context of noncompliance, Mao (2024) defined a local Mann-Whitney treatment effect estimand given by $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=10\}$, which is also a special case of the PGCE.
+Different from our approach, Mao (2024) used an instrumental variable approach and discussed point and interval identification of this estimand based on monotonicity and exclusion restriction.
 2.3 Identification assumptions
@@
 To identify the PGCE from the observed data, we require the following assumptions.
-Assumption 1 (Treatment ignorability).
+Assumption 1 (Treatment ignorability)
 […]
@@
 For observational studies, Assumption 1 can be satisfied if all confounders affecting the treatment assignment, the intermediate variable, and the final outcome are measured and included in the measured set of pre-treatment covariates, $X$.
-Assumption 2 (Monotonicity).
+Assumption 2 (Monotonicity)
 $D_{i}(1)\geq D_{i}(0)$ for all $i$.
-Assumption 2 states that the treatment assignment only affects the intermediate outcome in a non-negative (or non-positive if assume $D_{i}(1)\leq D_{i}(0)$) direction, and restricts the number of possible strata by excluding the harmed or defiers stratum $S=01$ (or $S=10$).
+Assumption 2 states that the treatment assignment only affects the intermediate outcome in a non-negative (or non-positive if assume $D_{i}(1)\leq D_{i}(0)$) direction and restricts the number of possible strata by excluding the harmed or defiers stratum $S=01$ (or $S=10$).
 For example, when $D$ represents the survival status, monotonicity holds when the treatment is expected not to harm the survival of each unit.
 When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to active treatment.
-Assumption 3 (Mean principal ignorability).
+Assumption 3 (Mean principal ignorability)
 Given the same values of pre-treatment covariates for a pair of units ($X_{1}=x_{1}$ and $X_{2}=x_{2}$), the expected contrast function is exchangeable across certain pairs of principal strata.
@@
 In other words, within the observed data stratum defined by $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the combinations of possible latent principal strata, $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$, are exchangeable with respect to the expected contrast.
-Similarly, the conditional expected contrasts within group $\mathcal{S}_{11}$ and $\mathcal{S}_{00}$ are all equal to estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=2,X_{1},X_{2}]$ and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, respectively.
+Similarly, the conditional expected contrasts within group $\mathcal{S}_{11}$ and $\mathcal{S}_{00}$ are all equal to estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$ and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, respectively.
 That is, with Assumption 3, each one of the subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only.
-Remark 2.1.
+Remark 1
 […]
-Assumptions 2 and 3 are instrumental in achieving nonparametric identification of the PGCE, however, their applicability, in general, cannot be directly tested using observed data alone without prior knowledge.
-Although our primary focus is on identifying and efficiently estimating the PGCE under these two structural assumptions, we provide a confounding function approach for sensitivity analyses under assumed violations of these two structural assumptions in Section S7 of the Online Supplement.
+Assumptions 2 and 3 are instrumental in achieving nonparametric identification of the PGCE; however, their applicability, in general, cannot be directly tested using observed data alone without prior knowledge.
+Although our primary focus is identifying and efficiently estimating the PGCE under these two structural assumptions, we provide a confounding function approach for sensitivity analyses under assumed violations of these two structural assumptions in Section S7 of the Online Supplement.
 3 Nonparametric identification
@@
 To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution, i.e., the treatment probability, the principal score, and the mean of the outcome contrast conditional on covariates (referred to as pairwise outcome mean hereafter for brevity).
-We consider both the observational and experimental studies, and define the treatment probability or the propensity score, by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin,, 1983).
-In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu,, 2016), for $s\in\{10,00,11\}$ under Assumption 2.
+We consider both the observational and experimental studies and define the treatment probability or the propensity score, by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin, 1983).
+In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu, 2016), for $s\in\{10,00,11\}$ under Assumption 2.
 Furthermore, we use averages over the distribution of covariates, $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$, to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively.
@@
 $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\},$
-for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. We provide some examples of the pairwise outcome mean models for specific contrast functions below.
-Example 3.1 (Pairwise mean function for average treatment effect).
+for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Below, we provide some examples of the pairwise outcome mean models for specific contrast functions.
+Example 3 (Pairwise mean function for average treatment effect)
 For studying the average treatment effect, the contrast function is specified as $h(u,v)=u-v$.
@@
 $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2}).$ (7)
-Example 3.2 (Pairwise mean function for probability index).
+Example 4 (Pairwise mean function for probability index)
 For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$.
-If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3.1.
-In this case, following Wolfe and Hogg, (1971), one can derive the pairwise outcome mean as
+If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3.
+In this case, following Wolfe and Hogg (1971), one can derive the pairwise outcome mean as
 $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\Phi\left\{\frac{f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2})}{\sqrt{2}\sigma}\right\}.$ (8)
-Example 3.3 (Pairwise outcome mean for win estimands).
-Consider the win estimands developed in Example 2.2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand).
-For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao, (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
+Example 5 (Pairwise outcome mean for win estimands)
+Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand).
+For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
 $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ (9)
@@
 where $p_{zq}(D,X)$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $X$, $Q$ is the largest category of the ordinal outcome.
-For estimating the principal scores, we first define conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$.
-Then under Assumption 2, the principal scores can be expressed as (Jiang et al.,, 2022)
+For estimating the principal scores, we first define the conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$.
+Then under Assumption 2, the principal scores can be expressed as (Jiang et al., 2022)
 $e_{10}(X)=p_{1}(X)-p_{0}(X),~{}~{}~{}e_{00}(X)=1-p_{1}(X),~{}~{}~{}e_{11}(X)=p_{0}(X).$
@@
 We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$.
-These elements then prepare us to state the set of nonparametric identification results in the following Theorem 3.4.
-Theorem 3.4 (Nonparametric identification).
+These elements then prepare us to state the set of nonparametric identification results in the following Theorem 1.
+Theorem 1 (Nonparametric identification)
 Suppose Assumptions 1-3 hold, $e_{s}>0$ for $s\in\{10,00,11\}$, and $10<\pi(X)<$ 1 for all $X\in\mathcal{X}$, the support of covariates.
@@
     $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{p_{0}(X_{1})}{p_{0}}\right\}\left\{\frac{p_{0}(X_{2})}{p_{0}}\right\}\mu_{1101}(X_{1},X_{2})\right].$
-We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 3.4, and the identification of the other two PGCEs follows the same strategy.
-In Theorem 3.4(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$.
+We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 1, and the identification of the other two PGCEs follows the same strategy.
+In Theorem 1(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$.
 Specifically, under Assumptions 1 and 2, we have
 $\mathbb{E}\left[\left\{\frac{e_{10}(X_{1})}{p_{1}-p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{10}(X_{2})}{p_{1}-p_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}\bigg{\rvert}X_{1},X_{2}\right]=\frac{e_{10}(X_{1})e_{10}(X_{2})}{e_{10}^{2}},$
-which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by the Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$.
+which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$.
 This ensures the identification of $\tau_{h}^{10}$.
-In Theorem 3.4(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
+In Theorem 1(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
 $\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
@@
 Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$.
-Finally, in Theorem 3.4(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 3.4(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$.
+Finally, in Theorem 1(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 1(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$.
 The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.
-Remark 3.1.
-Theorem 3.4 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al.,, 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al., (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
+Remark 2
+Theorem 1 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al., 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al.
+(2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
 3.2 Estimators based on nonparametric identification formulas
@@
 Specifically, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$.
-Additionally, we use ‘+’ in the subscript to denote conditions where more than one nuisance function is correctly specified.
-In addition, the union notation, $\cup$’, denotes the condition where at least one nuisance function is correctly specified.
-We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts, and expectations with empirical averages.
+Additionally, we use ‘+’ in the subscript to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
+We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts and expectations with empirical averages.
 For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$.
-Along the lines in Jiang et al., (2022), we define the following quantity for any function $f(Y,D,X)$:
+Along the lines in Jiang et al.
+(2022), we define the following quantity for any function $f(Y,D,X)$:
 […]
@@
 $\psi_{D_{z}}(\mathcal{O})=\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X)\}}{\mathbb{P}(Z=z|X)}+p_{z}(X).$ (11)
-Equation (11) is useful in the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome (Tsiatis,, 2006):
+Equation (11) is useful in the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome (Tsiatis, 2006):
 $\widehat{p}_{z}\equiv\mathbb{E}_{n}\psi_{D_{z}}(\mathcal{O};\widehat{\alpha},\widehat{\beta})=\mathbb{E}_{n}\left[\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X;\widehat{\beta})\}}{\pi(X;\widehat{\alpha})}+p_{z}(X;\widehat{\beta})\right],$ (12)
 which improves upon the plug-in estimator $\mathbb{E}_{n}p_{z}(X;\widehat{\beta})$ in terms of estimation efficiency and robustness, and is consistent for $p_{z}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$.
-Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart,, 1998).
-For example, Theorem 3.4(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
+Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart, 1998).
+For example, Theorem 1(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
 $\widehat{\tau}_{h,\pi+p}^{10}=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{1}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right].$
-Theorem 3.4(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
+Theorem 1(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
 $\widehat{\tau}_{h,\pi+\mu}^{10}$ $=\mathbb{U}_{n}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1};\widehat{\alpha})-D_{1}(1-Z_{1})/\{1-\pi(X_{1};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\right.$
 ×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2};\widehat{\alpha})-D_{2}(1-Z_{2})/\{1-\pi(X_{2};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right),$
-and Theorem 3.4(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
+and Theorem 1(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
 $\widehat{\tau}_{h,p+\mu}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{p_{1}(X_{1};\widehat{\beta})-p_{0}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\left\{\frac{p_{1}(X_{2};\widehat{\beta})-p_{0}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right].$
@@
 Specifically, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment probability and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
-Remark 3.2.
-The weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score, which is an extension of the inverse propensity score weighted estimator in Mao, (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable.
-The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately after applying the Taylor’s expansion on the ratio estimators (Jiang et al.,, 2022); see the Online Supplement for additional details.
-The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be straightforwardly obtained using empirical process theory for the parametric class (van der Vaart,, 1998, §19.2).
-The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao, (2017).
+Remark 3
+The weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score, which is an extension of the inverse propensity score weighted estimator in Mao (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable.
+The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately after applying the Taylor’s expansion on the ratio estimators (Jiang et al., 2022); see the Online Supplement for additional details.
+The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be straightforwardly obtained using empirical process theory for the parametric class (van der Vaart, 1998, §19.2).
+The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao (2017).
 4 Multiply robust and efficient $U$-estimators
@@
 4.1 The efficient influence function for the PGCE
-We first develop more efficient estimators with root-$n$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al.,, 1993).
+We first develop more efficient estimators with root-$n$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al., 1993).
 The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators.
 Several key properties of the EIF suggest its promise for constructing optimal estimators.
-First, the EIF often integrates more than one nuisance model and facilitates estimators that are robust against misspecification of some nuisance models (Tsiatis,, 2006; Mao,, 2017).
+First, the EIF often integrates more than one nuisance model and facilitates estimators that are robust against misspecification of some nuisance models (Tsiatis, 2006; Mao, 2017).
 Second, the asymptotic variance lower bound is represented by the variance of the EIF, providing a useful benchmark to define efficient estimators.
-Third, the EIF provides a device for integrating advanced machine learners for the nuisance parameters, which can lead to feasible estimators that achieve the asymptotic variance lower bound (Chernozhukov et al.,, 2018; Rotnitzky et al.,, 2021).
+Third, the EIF provides a device for integrating advanced machine learners for the nuisance parameters, which can lead to feasible estimators that achieve the asymptotic variance lower bound (Chernozhukov et al., 2018; Rotnitzky et al., 2021).
 We will explore these properties by first formalizing the EIF for the PGCE.
@@
 Then, the following results are obtained.
-Theorem 4.1 (EIF).
+Theorem 2 (EIF)
 Under Assumptions 1-3, (a) the EIF for $\tau_{h}^{10}$ is $\varphi_{h}^{10}(\mathcal{O})=\{\varphi_{h,N}^{10}(\mathcal{O})-\tau_{h}^{10}\varphi_{D}^{10}(\mathcal{O})\}/(p_{1}-p_{0})^{2}$, where $\varphi_{h,N}^{10}(\mathcal{O})=\phi_{h}^{10}(\mathcal{O})-2\tau_{h,N}^{10}$ and
@@
 $\varphi_{D}^{11}(\mathcal{O})=2\psi_{D_{0}}(\mathcal{O})p_{0}-2\tau_{D}^{11}.$
-The EIFs in Theorem 4.1 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey, (2022) and Hines et al., (2022) with respect to the distribution of observed data, $\mathcal{O}$.
-The derivation for $\varphi_{D}^{s}(\mathcal{O})$ is then follows the chain rule.
+The EIFs in Theorem 2 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey (2022) and Hines et al.
+(2022) with respect to the distribution of observed data, $\mathcal{O}$.
+The derivation for $\varphi_{D}^{s}(\mathcal{O})$ then follows the chain rule.
 For $\varphi_{h,N}^{s}(\mathcal{O})$, we re-express the PGCE in terms of the observed data as
@@
 where $\mathfrak{h}_{1}^{s}(y)=\mathbb{E}\{h(y,Y^{\prime})|Z^{\prime}=0,S=S^{\prime}=s\}$ and $\mathfrak{h}_{0}^{s}(y)=\mathbb{E}\{h(Y^{\prime},y)|Z^{\prime}=1,S=S^{\prime}=s\}$, with the conditional expectations taken with respect to $Y^{\prime}$.
-Importantly, the EIF for the PGCE remains unchanged after this re-expression, since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged.
-Based on Theorem 4.1, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
-Remark 4.1.
+Importantly, the EIF for the PGCE remains unchanged after this re-expression since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged.
+Based on Theorem 2, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
+Remark 4
 Without considering the intermediate variable, the EIF for the generalized causal effect estimand, $\tau_{h}$ in (1), $\varphi_{h}(\mathcal{O})$, can be derived using the same approach as $\varphi_{h}(\mathcal{O})=\phi_{h}(\mathcal{O})-2\tau_{h}$, where
@@
 The function $\mu_{10}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=1,Z_{2}=0,X_{1},X_{2}\}$ is the pairwise outcome mean within the observed stratum $(Z_{1}=1,Z_{2}=0)$.
-This coincides with the EIF developed in Mao, (2017) using the projection approach (Tsiatis,, 2006), and intuitively, $\varphi_{h}(\mathcal{O})$ can be viewed as a special case of $\varphi_{h,N}^{s}(\mathcal{O})$ in the absence of the intermediate variable.
+This coincides with the EIF developed in Mao (2017) using the projection approach (Tsiatis, 2006), and intuitively, $\varphi_{h}(\mathcal{O})$ can be viewed as a special case of $\varphi_{h,N}^{s}(\mathcal{O})$ in the absence of the intermediate variable.
 4.2 Multiply robust estimators
-The EIFs given in Theorem 4.1 imply the semiparametric efficiency lower bound, and pave the way for formulating estimators for the PGCEs.
+The EIFs given in Theorem 2 imply the semiparametric efficiency lower bound and pave the way for formulating estimators for the PGCEs.
 However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator.
-We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 4.1, thereby achieving semiparametric efficiency with correctly specified nuisance functions.
+We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 2, thereby achieving semiparametric efficiency with correctly specified nuisance functions.
 To this end, we first notice that each PGCE can be rewritten as the following:
@@
 $\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (14)
-which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 4.2.
-Theorem 4.2 (Multiply robustness and efficiency under parametric working models).
+which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 3.
+Theorem 3 (Multiply robustness and efficiency under parametric working models)
 Under Assumptions 1-3, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$.
 Then, the estimator $\widehat{\tau}_{h}^{s}$ in (14) is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
-Under Theorem 4.2, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified, and is semiparametrically efficient if all three nuisance models are correctly specified.
+Under Theorem 3, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified.
 The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator.
@@
 The specific derivations are provided in the Online Supplement Section S4.1 due to the complexity of the expressions.
-The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao, (2017) and Theorem 3 of Jiang et al., (2022).
-As discussed in Remark 3.2, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator.
+The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao (2017) and Theorem 3 of Jiang et al.
+(2022).
+As discussed in Remark 3, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator.
 Details of the proof are provided in Section S4.2 of the Online Supplement.
@@
 4.3 Debiased machine learning estimators
-Debiased machine learning is a burgeoning method for efficient estimation of the causal effects in experimental and observation studies, and is often built upon the orthogonal moment functions (Chernozhukov et al.,, 2018, 2022) or the EIFs.
-Because of the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al.,, 2022), perturbations in the first-step estimates typically have no significant influence on local properties of the desired estimator for the causal estimand.
-In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates (Rotnitzky et al.,, 2021).
+Debiased machine learning is a burgeoning method for efficient estimation of the causal effects in experimental and observation studies and is often built upon the orthogonal moment functions (Chernozhukov et al., 2018, 2022) or the EIFs.
+Because the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al., 2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand.
+In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates (Rotnitzky et al., 2021).
 Compared to parametric working models, leveraging machine learning working models for the nuisance functions is more advantageous for its robustness against potential misspecification.
-However, machine learning first-step estimates ($\widehat{\pi}(\cdot)$, $\widehat{p}_{z}(\cdot)$, and $\widehat{\mu}_{z_{1}d_{1}z_{2}d_{2}}(\cdot,\cdot)$) have increased complexity and could no longer satisfy the Donsker class condition (Chernozhukov et al.,, 2018).
-To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting is adopted, where the entire sample is divided into separate parts for the estimation of the nuisance functions and the estimation of the causal estimand respectively (Chernozhukov et al.,, 2018).
-Mild regulatory conditions, mostly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy,, 2023).
-Because of the unique structure of the PGCEs, we adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur, (2023).
+To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting is adopted, where the entire sample is divided into separate parts for the estimation of the nuisance functions and the estimation of the causal estimand respectively (Chernozhukov et al., 2018).
+Mild regulatory conditions, mostly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy, 2023).
+Because of the unique structure of the PGCEs, we adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur (2023).
 Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition =$\mathcal{I}=\{I_{1},\ldots,I_{K}\}$.
-Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $(i,j)$ with $1\leq i<j\leq N$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$.
-Chernozhukov et al., (2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$.
+Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $\{(i,j),1\leq i<j\leq N\}$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$.
+Chernozhukov et al.
+(2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$.
 The following example is used to provide a demonstration of $\mathcal{I}$ and $\mathcal{P}$.
-Example 4.3 (Sample splitting for $U$-statistics).
+Example 6 (Sample splitting for $U$-statistics)
 Consider a scenario where $n=10$ and $K=3$.
@@
 $\mathbb{E}_{n,K}\psi_{D_{z}}(\widehat{\vartheta})=n^{-1}\sum_{k=1}^{K}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}).$
-Remark 4.2.
-Estimators given in (15) are the type II estimator proposed in Chernozhukov et al., (2018) for estimands without an intermediate variable. Alternatively, one could first obtain separate estimates
+Remark 5
+Estimators given in (15) are the type II estimator proposed in Chernozhukov et al. (2018) for estimands without an intermediate variable. Alternatively, one could first obtain separate estimates
 $\mathbb{U}_{|P_{l}|}g_{h}^{s}(\widehat{\theta}_{l})=|P_{l}|^{-1}\sum_{(i,j)\in P_{l}}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})~{}~{}~{}\text{and}~{}~{}~{}\mathbb{E}_{|I_{k}|}\psi_{D_{z}}(\hat{\vartheta}_{k})=|I_{k}|^{-1}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}),$
@@
 The following result summarizes the efficiency properties of (15).
-Theorem 4.4 (Efficiency under nonparametric working models).
+Theorem 4 (Efficiency under nonparametric working models)
 Under Assumptions 1-3, and
@@
   • (d)
-    there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
-then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 4.1.
-Theorem 4.4 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al., (2022, Lemma 15) and Escanciano and Terschuur, (2023, Lemma 4).
+    $g_{h,1}^{s}(\mathcal{O};\theta)\equiv\mathbb{E}g_{h}^{s}(\mathcal{O},o;\theta)$ and $\psi_{D_{z}}(\mathcal{O};\vartheta)$ are $\mathbb{P}$-Donsker with probability approaching 1, and there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
+then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 2.
+Theorem 4 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al.
+(2022, Lemma 15) and Escanciano and Terschuur (2023, Lemma 4).
 To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$.
-Regulatory conditions in Theorem 4.4 are mostly mild.
+Regulatory conditions in Theorem 4 are mostly mild.
 Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$.
 Condition (b) assumes that machine learning first-step estimators are consistent as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners.
-Conditions (c) and (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al.,, 2018).
+Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al., 2018).
+The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator, which has been commonly adopted in debiased machine learning literature (Vansteelandt and Dukes, 2022; Wang et al., 2024).
 Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (15) via nonparametric bootstrap.
 5 Simulation studies
-We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 3.4, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15).
+We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 1, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15).
 In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
@@
 Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise.
-Also, it is worth noting that, in practice, model misspecification pertains to the flexible debiased machine learning estimators, because their estimation accuracy could be affected by the quality of the input covariate feature matrix.
-For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 3.2.
+Also, it is worth noting that, in practice, model misspecification pertains to the flexible debiased machine learning estimators because their estimation accuracy could be affected by the quality of the input covariate feature matrix.
+For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 4.
 For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries.
@@
 Figure 1 presents the results of the five estimators under comparison in scenarios where, at most, one of the three nuisance functions is misspecified.
-Plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified, while the triply robust and debiased machine learning estimators are consistent across all four scenarios (correct specification of all nuisance functions or subsets of nuisance functions).
+Plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified.
+In contrast, the triply robust and debiased machine learning estimators are consistent across all four scenarios (correct specification of all nuisance functions or subsets of nuisance functions).
 Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure 1 in Section S6 of the Online Supplement.
@@
 Horizontal dashed lines represent the true values of the PGCEs.
-5.2 The win-ratio estimands with an intermediate variable
-In the second simulation study, we consider the case of the win-ratio estimands introduced in Examples 2.2 and 3.3, with the final potential outcomes being ordinal.
+5.2 The win ratio estimands with an intermediate variable
+In the second simulation study, we consider the case of the win ratio estimands introduced in Examples 2 and 5, with the final potential outcomes being ordinal.
 We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study.
@@
 (c) $\tau_{h}^{11}$
-Figure 2: Violin plots of estimators for the win-ratio estimands in scenarios where at most one nuisance function is misspecified.
+Figure 2: Violin plots of estimators for the win ratio estimands in scenarios where at most one nuisance function is misspecified.
 Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15).
@@
 We illustrate the proposed methods with a data example from the U.S.
-Jobs Corps study (Schochet et al.,, 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24.
-The dataset (named ‘JC’ with $9,240$ observations, available in the R package causalweight) was from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either, the treatment group, i.e., receiving an offer to participate in Job Corps immediately, or, the control group, i.e., not participating in Job Corps until three years later.
+Jobs Corps study (Schochet et al., 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24.
+The dataset (named ‘JC’ with $9,240$ observations, available in the R package causalweight) was from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, i.e., receiving an offer to participate in Job Corps immediately, or, the control group, i.e., not participating in Job Corps until three years later.
 Follow-up surveys were administered to both groups 24, 36, and 48 months after the random assignment to collect data on the outcomes of interest.
@@
 As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects.
-In each data analysis below, we adjust for baseline covariates including baseline characteristics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
-6.1 The survivor probability index for continuous outcomes
+In each data analysis below, we adjust for baseline covariates, including baseline characteristics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
+6.1 The SPI for continuous outcomes
 We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$.
-In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al., (2009).
+In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al.
+(2009).
 Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps.
@@
 ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15).
-The corresponding 95% confidence interval is given in the following parenthesis.
-Methods Survivor proability index estimates (standard errors)
+The corresponding 95% confidence interval (CI) is given in the following parenthesis.
+Methods SPI estimates
 M1 0.521 (0.508, 0.533)
@@
 DML 0.539 (0.527, 0.552)
-6.2 Principal win-ratio for ordinal outcomes
+6.2 Principal win ratio for ordinal outcomes
 We provide additional analysis of the Job Corps data to illustrate the application of the proposed methods to address ordinal outcomes with noncompliance.
-In this example, we use the actual treatment status as an intermediate variable, similar to Chen and Flores, (2015).
+In this example, we use the actual treatment status as an intermediate variable, similar to Chen and Flores (2015).
 Under the monotonicity assumption, three strata potentially exist, including the compliers (stratum $10$), the never-takers (stratum $00$), and the always-takers (stratum $11$).
-To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu et al., (2018).
-Under this setting, we estimate the principal win-ratios within each principal strata similar to the definition in (4).
+To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu et al.
+(2018).
+Under this setting, we estimate the principal win ratios within each principal strata similar to the definition in (4).
 For methods based on parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are estimated via logistic regression.
@@
 The Jobs Corps intervention assignment positively affects participants’ earnings one year after the assignment for compliers and always-takers, and the effect appears to be more pronounced for compliers.
-For example, using the debiased machine learning estimator, there is 27% increase in the chances of being in a higher earning category for a complier under treatment than a complier under control.
+For example, using the debiased machine learning estimator, there is a 27% increase in the chances of being in a higher earning category for a complier under treatment than a complier under control.
 The 95% lower confidence limits obtained from all five methods are strictly above the null value of 1.
-For the never-takers, however, the effect of Jobs Corps appears more neutral, as the principal win-ratio is estimated to be closer to 1 and three out of five methods (M1, triply robust and debiased machine learning) produce 95% confidence intervals that include the null value of 1.
-Interestingly, in this reanalysis using a win-ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable.
+For the never-takers, however, the effect of Jobs Corps appears more neutral, as the principal win ratio is estimated to be closer to 1, and three out of five methods (M1, triply robust and debiased machine learning) produce 95% confidence intervals that include the null value of 1.
+Interestingly, in this reanalysis using a win ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable.
 In particular, the statistically significant effect among the always-takers is indicative of a positive psychological effect.
 That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence having a large chance to move into a higher earning category) compared to those who switched from their initial control assignment.
-Table 2: Estimated win-ratios within different compliance strata for the Jobs Corps data.
+Table 2: Estimated win ratios within different compliance strata for the Jobs Corps data.
 ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15).
@@
 The $U$-estimators become locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric models or by machine learning models with cross-fitting.
-In light of the increasing interest in pairwise comparison estimands in studies with a composite endpoint (Buyse,, 2010; Bebu and Lachin,, 2015) in biomedical research, our methods can also be useful when there is a need to additionally address a binary intercurrent event, such as treatment noncompliance or death (Lipkovich et al.,, 2022; Mao,, 2024)
+In light of the increasing interest in pairwise comparison estimands in studies with a composite endpoint (Buyse, 2010; Bebu and Lachin, 2015) in biomedical research, our methods can also be useful when there is a need to additionally address a binary intercurrent event, such as treatment noncompliance or death (Bornkamp et al., 2021; Lipkovich et al., 2022; Mao, 2024)
 The estimands and efficient $U$-estimators were developed to address a binary treatment and a binary intermediate outcome specifically but without restrictions on the final outcome.
-Although this is a standard setup for principal stratification analysis, additional work is needed to expand the identification conditions and results to address extensions of the PGCE estimands with a continuous treatment and/or a continuous intermediate variable (Antonelli et al.,, 2023).
+Although this is a standard setup for principal stratification analysis, additional work is needed to expand the identification conditions and results to address extensions of the PGCE estimands with a continuous treatment and/or a continuous intermediate variable (Antonelli et al., 2023).
 With continuous treatments, the definition of pairwise comparison estimands becomes non-trivial as there are infinitely many contrast functions (based on an infinitesimal increment of the treatment value) that one can specify to define $\tau_{h}$ even in the absence of an intermediate variable.
-On the other hand, with a continuous intermediate variable, there are infinitely many principal strata that can lead to conceptual and implementation challenges (Schwartz et al.,, 2011).
-Focusing on a linear contrast function, Lu et al., (2023) have developed nonparametric identification and semiparametric efficient estimators for principal causal effects with a continuous intermediate variable, which may inspire future theoretical development that allows for a more general nonlinear contrast function in the estimands definition.
-Finally, as in all principal stratification analyses that achieve point identification, our $U$-estimators for estimating PGCEs depend on the monotonicity and the mean principal ignorability assumptions, that are unverifiable from the observed data alone.
-In Section S7 of the Online Supplement, we have described a sensitivity function approach to obtain the bias-corrected $U$-estimators for estimating the PGCEs, when either the monotonicity assumption or the mean principal ignorability assumption is violated.
+On the other hand, with a continuous intermediate variable, there are infinitely many principal strata that can lead to conceptual and implementation challenges (Schwartz et al., 2011).
+Focusing on a linear contrast function, Lu et al.
+(2023) have developed nonparametric identification and semiparametric efficient estimators for principal causal effects with a continuous intermediate variable, which may inspire future theoretical development that allows for a more general nonlinear contrast function in the estimands definition.
+Finally, as in all principal stratification analyses that achieve point identification, our $U$-estimators for estimating PGCEs depend on the monotonicity and the mean principal ignorability assumptions that are unverifiable from the observed data alone.
+In Section S7 of the Online Supplement, we have described a sensitivity function approach to obtain the bias-corrected $U$-estimators for estimating the PGCEs when either the monotonicity assumption or the mean principal ignorability assumption is violated.
 The bias-corrected $U$-estimators remain triply robust to potential working model misspecifications and provide means to generate a range of point estimates under assumed departures from the key identification assumptions.
+Acknowledgement
+Research in this article was partially supported by the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (NHLBI, grant number R00-HL141678).
+All statements in this report, including its findings and conclusions, are solely those of the authors and do not necessarily represent the views of the NIH.

### v4

@@
 While the principal average treatment effects have traditionally been the default target of inference, it may not be sufficient when the interest lies in the relative favorability of one potential outcome over the other within the principal stratum.
-We thus introduce the principal generalized causal effect estimands, which extend the principal average causal effects to accommodate nonlinear contrast functions.
+We introduce the principal generalized causal effect estimands to extend the principal average causal effects, accommodating nonlinear contrast functions.
 Under principal ignorability, we expand the theoretical results in Jiang et al.
-(2022) to a much wider class of causal estimands in the presence of a binary intermediate variable.
-We develop identification formulas and derive the efficient influence functions of the generalized estimands for principal stratification analyses.
+(2022) to a much broader class of causal estimands in the presence of a binary intermediate variable.
+We develop identification results and derive the efficient influence functions of the generalized estimands for principal stratification analyses.
 These efficient influence functions motivate a set of multiply robust estimators and lay the ground for obtaining efficient debiased machine learning estimators via cross-fitting based on $U$-statistics.
 The proposed methods are illustrated through simulations and the analysis of a data example.
-Keywords: Causal inference, efficient influence function, principal stratification, multiply robust estimation, $U$-statistics, win ratio
+Keywords: Causal inference, efficient influence function, principal stratification, multiply robust estimation, win ratio, probabilistic index
 1 Introduction
-For causal inference under the Neyman-Rubin model (Rubin, 1974), the average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest.
+The average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest in causal inference.
 Specifically, let $\{Y(1),Y(0)\}$ denote a pair of potential outcomes for a randomly selected individual from a super-population, and then the average causal effect estimand is defined as $\mathbb{E}\{Y(1)-Y(0)\}$.
@@
 That is, given a contrast function $h(u,v)$ defined on the product space, $\mathcal{Y}\times\mathcal{Y}$ ($\mathcal{Y}$: the potential outcome space), a generalized estimand can be defined as $\mathbb{E}[h\{Y(1),Y(0)\}]$.
-Examples include the average causal effect when $h(u,v)=u-v$, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$ (Lu et al., 2020).
-Because $Y(1)$ and $Y(0)$ are never observed simultaneously, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$.
-However, a closely related quantity, the generalized causal effect (O’brien, 1988; Mao, 2017), can be formulated as:
+Examples include the average causal effect when h(u,v)=uvh(u,v)=u-v, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$.
+Because $Y(1)$ and $Y(0)$ are never observed simultaneously, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$ (see Lu et al.
+(2020) for an example for bounding such an estimand with ordinal outcomes).
+However, a closely related quantity, the generalized causal effect, can be formulated:
 $\tau_{h}=\int\int h(u,v)\mathrm{d}F_{1}(u)\mathrm{d}F_{0}(v),$ (1)
@@
 Let $\{Y_{1}(1),Y_{1}(0)\}$ and $\{Y_{2}(1),Y_{2}(0)\}$ denote two independent replicates of $\{Y(1),Y(0)\}$, and $\tau_{h}$ can be re-expressed in a familiar form, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}]$.
-Two notable examples are the probability index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al., 2015), and the win ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al., 2011).
-These estimable quantities are alternative measurements for the population causal effect and can be straightforwardly interpreted.
-For example, the probability index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than $F_{0}$ and is robust against extreme values in the tails of the potential outcome distributions.
-When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al., 1996), mortality event (Zhang et al., 2009) or dropout (Frangakis and Rubin, 1999), often have important implications even on the definition of a causal effect and require careful adjustment.
-The intermediate variable is not guaranteed to be balanced via randomization, and direct conditioning on the intermediate variable leads to sample selection bias and non-causal comparisons, even in randomized experiments.
-As a solution, the principal stratification framework (Frangakis and Rubin, 2002) segments the population into principal strata (subpopulations) according to joint potential values of the intermediate variable under alternative conditions.
-The stratum-specific causal effect—the principal causal effect—can thus be defined and is causally interpretable.
+Two notable examples of the generalized causal effect are the probabilistic index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al., 2015), and the win ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al., 2011).
+These estimable quantities are alternative measurements for the population effect and carry a causal interpretation (Fay and Li, 2024).
+In certain cases, such estimands offer advantages that make them relevant in modern applications.
+First, the probabilistic index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than a unit from $F_{0}$.
+This formulation generally exhibits greater robustness to distributional features.
+That is, unlike mean contrasts, which can be distorted by heavy tails, skewness, or outliers, probabilistic indices require fewer distribution assumptions and are less sensitive to irregularities in the outcome distribution (Hauck et al., 2000; Acion et al., 2006).
+Second, probabilistic indices are also well suited for ordinal outcomes, where the scientific question itself concerns relative ordering rather than pure numerical differences.
+Hence, as De Schryver and De Neve (2019) noted, “the probabilistic index provides an ordinal answer to an ordinal question”, and it relies only on the rank comparison rather than the metric scale or magnitude of differences.
+Third, the win-type estimand is now increasingly used for prioritized composite outcomes, where rank-based win-loss comparisons naturally align with clinical priorities across multiple endpoints.
+Introduced by Pocock et al.
+(2011), the win ratio has gained substantial traction in clinical trials with composite outcomes, as it offers a principled way to combine endpoints of varying clinical importance while providing an interpretable summary of treatment benefit in terms of the probability of winning versus losing.
+When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al., 1996) and death (Zhang et al., 2009) often have important implications on the definition of a causal effect and require careful adjustment.
+The principal stratification framework (Frangakis and Rubin, 2002) offers one solution by segmenting the population into principal strata (subpopulations) based on the joint potential values of the intermediate variable under alternative conditions.
+The stratum-specific causal effect, referred to as the principal causal effect, can thus be defined and is causally interpretable.
 However, a predominant focus in prior works has been on the average causal effect within each principal stratum, and there is a lack of related development when the generalized causal effect is of interest within each subpopulation.
 Therefore, in this article, we propose a new class of estimands, the principal generalized causal effect (PGCE), which accommodates an arbitrary and potentially nonlinear contrast function.
-The PGCE, under a linear contrast function, reduces to the usual principal average causal effect.
-As we demonstrate in due course, the PGCE can represent the stratum-specific probability index and strata-specific win ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within subpopulations of interest.
-Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al., 1996) and principal ignorability (Jo and Stuart, 2009; Ding and Lu, 2016), which are typically invoked for principal stratification analyses.
-The monotonicity assumption states that the treatment has a nonnegative causal effect on the intermediate outcome of each unit, effectively restricting the number of potential strata.
+As we demonstrate in due course, the PGCE can represent the stratum-specific probabilistic index and strata-specific win ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within subpopulations of interest.
+Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al., 1996) and principal ignorability (Ding and Lu, 2016), which are typically invoked for principal stratification analyses.
+The monotonicity assumption states that the treatment has a nonnegative causal effect on the intermediate outcome of each unit, restricting the number of potential strata.
 This condition is considered plausible when, for example, the intermediate variable is death, and the treatment is known to benefit survival, or when the intermediate variable is the actual treatment received and the treatment does not discourage treatment uptake.
-The principal ignorability assumption, on the other hand, requires that the observed pre-treatment covariates are sufficient to control for confounding between the intermediate variable and the final outcome.
+The principal ignorability assumption requires that the observed pre-treatment covariates are sufficient to control for confounding between the intermediate variable and the outcome.
 For the principal average causal effect, under monotonicity and principal ignorability, Ding and Lu (2016) formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang et al.
-(2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and further developed multiply robust and locally efficient estimators under the additional assumption of treatment unconfoundedness.
-In contrast, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
+(2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and developed multiply robust and locally efficient estimators; Cheng et al.
+(2023, 2024) extended the multiply robust and locally efficient estimators to estimate principal survival causal effect with a right-censored time-to-event outcome; Tong et al.
+(2024) expanded multiply robust and locally efficient estimators to estimate fine principal causal effects with multiple treatments.
+Despite these advancements, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
 The contributions of this article are several-fold.
-First, we propose a class of PGCE estimands in the presence of an intermediate variable and develop complementary nonparametric identification formulas under treatment unconfoundedness, monotonicity, and principal ignorability.
+First, we propose a class of PGCE estimands in the presence of a binary intermediate variable and develop complementary nonparametric identification formulas under the assumptions of treatment unconfoundedness, monotonicity, and principal ignorability.
 Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands and seek efficiency improvement by deriving their corresponding efficient influence functions (Bickel et al., 1993, EIFs).
-Specifically, we show that when the contrast function $h(u,v)=u-v$, our EIFs reduce to those derived by Jiang et al.
-(2022) as a special case but are otherwise more general to accommodate arbitrary contrast functions.
-A major complexity in the context of a nonlinear contrast function, however, is that the estimating equations based on the EIFs require nuisance functions defined over pairs of observations.
-Thus, to obtain a feasible estimator, we define ratio $U$-statistics whose influence functions are the same as the EIFs.
-We show that the $U$-estimators are triply robust in the sense that they are consistent and asymptotically normal when two of the three nuisance functions (including the treatment propensity score, principal score, and the potential outcome models) are correctly specified.
-When all nuisance functions are correctly specified, the $U$-estimators are locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds.
-The asymptotic properties of the $U$-estimators are developed upon the theory of empirical process (van der Vaart, 1998, §19), $U$-process (Arcones and Gine, 1993), as well as the theory of $U$-statistics (van der Vaart, 1998, §11-12).
-Finally, the construction of the multiply robust $U$-estimators facilitates the integration of machine learning estimators for the nuisance functions.
-To achieve full asymptotic efficiency, we learn each nuisance function using machine learning methods and operationalize the debiased machine learning $U$-estimators via cross-fitting (Chernozhukov et al., 2018).
-We discuss the modification needed for cross-fitting if a nonlinear contrast function is considered when defining the PGCE and prove the multiple robustness and efficiency property of the debiased machine learning estimators under mild regulation conditions (Chernozhukov et al., 2022; Escanciano and Terschuur, 2023).
+Specifically, we show that when the contrast function is h(u,v)=uvh(u,v)=u-v, our EIFs reduce to those in Jiang et al.
+(2022) as a special case.
+More broadly, our framework accommodates pairwise comparison estimands with arbitrary contrast functions under principal stratification, offering a substantially more general treatment and establishing the present work as a standalone contribution.
+A major complexity for nonlinear contrast functions, however, is that the estimating equations based on the EIFs require nuisance functions defined over pairs of observations, thereby requiring the theory of $U$-statistics to construct feasible estimators.
+We show that the proposed $U$-statistics are triply robust in the sense that they are consistent and asymptotically normal when two of the three nuisance functions are correctly specified.
+When all nuisance functions are correctly specified, our estimators are locally efficient.
+Finally, the construction of the multiply robust $U$-statistics facilitates the integration of nonparametric estimators for the nuisance functions.
+To achieve full asymptotic efficiency, we also estimate each nuisance function using data-adaptive machine learners under the cross-fitting scheme (Chernozhukov et al., 2018).
+We discuss the modifications needed for cross-fitting when a nonlinear contrast function is considered in defining the PGCE, and prove the nonparametric efficiency properties of the debiased machine learning estimators under mild regularity conditions.
 The rest of this article is organized as follows.
@@
 Section 3 discusses the nonparametric identification formulas for the PGCE and associated moment estimators.
-Section 4 derives the EIF for each PGCE estimand and develops the EIF-motivated feasible $U$-estimators and the debiased machine learning estimators.
+Section 4 derives the EIF for each PGCE estimand and develops the EIF-motivated estimators.
 Section 5 demonstrates the finite-sample performance of the proposed estimators via simulations.
 Section 6 illustrates the proposed methods with a data example from the United States Jobs Corps study (Schochet et al., 2001), and Section 7 concludes.
-Detailed proofs of all theoretical results are available in the Online Supplement.
+Detailed proofs of all theoretical results are presented in the Online Supplement.
 2 Notation, estimands and assumptions
 2.1 Setting and notation
-We consider a study (experimental or observational) with $n$ participants and two treatment conditions and let $Z_{i}$ denote the treatment assignment of participant $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 00 the control.
+We consider a study, either experimental or observational, with $n$ individuals and two treatment conditions.
+We let $Z_{i}$ denote the treatment assignment of individual $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 0 the control.
 We use $D_{i}(z)$ and $Y_{i}(z)$ to denote the potential outcomes of the intermediate variable and the final outcome variable of interest for participant $i$ under treatment $Z_{i}=z$ ($z\in\{0,1\}$), respectively.
 The intermediate variable $D_{i}\in\{0,1\}$ is assumed to be binary, but there is no restriction on the type of variable for $Y_{i}$.
-Under the stable unit treatment value assumption (Rubin, 1980, SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$.
-Let $X_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{X_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d).
+Under the Stable Unit Treatment Value Assumption (SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$.
+Let $X_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{X_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d.).
 2.2 Principal generalized causal effects
 We operate under the principal stratification framework (Frangakis and Rubin, 2002), and define distinct principal strata according to the joint potential values of the intermediate variable under alternative conditions.
-Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{11,10,01,00\}$.
+Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{10,01,00,11\}$.
 A key insight under the principal stratification framework is that although $D$ is a post-randomization variable, the stratum variable $S$ is unaffected by the assignment and can be viewed as a pre-treatment variable.
 Therefore, based on any contrast function $h$, we define the principal generalized causal effect (PGCE) estimand as
-$\tau_{h}^{s}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=S_{2}=s],~{}\text{for}~{}s\in\{11,10,01,00\},$ (2)
+$\tau_{h}^{s}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=S_{2}=s],~\text{for}~s\in\{10,01,00,11\},$ (2)
 where $Y_{1}(1)$ and $Y_{2}(0)$ denote the potential outcomes for two randomly selected units from the subpopulation with strata membership $s$ ($S_{1}=s$ and $S_{2}=s$ denote the strata membership for these two selected units).
-Of note, to ensure a proper causal interpretation of $\tau_{h}^{s}$, it is required that both elements of the pair for comparison are from the same principal stratum such that the comparison is made within a common subpopulation.
-If the two units are selected from different principal strata, the resulting quantity, $\tau_{h}^{s_{1},s_{2}}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2}]$ with $s_{1}\neq s_{2}$, remains estimable under this framework but generally lacks a meaningful causal interpretation; we thus do not further pursue this direction.
-Evidently, when a linear contrast function is specified as $h(u,v)=u-v$, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature; see, for example, Frangakis and Rubin (2002), Jiang et al.
-(2022) and Lipkovich et al.
-(2022).
-When a nonlinear contrast function is considered in defining the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum.
-The following two examples illustrate the PGCE with a specific contrast function.
-Example 1 (Probability index under truncation by death)
+When a linear contrast function is specified as h(u,v)=uvh(u,v)=u-v, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature (Frangakis and Rubin, 2002).
+Before proceeding, we provide a caveat on the definition of estimands in Remark 1.
+Remark 1
+One may wonder whether quantities of the form $\tau^{s_{1},s_{2}}_{h}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2}]$ with $s_{1}\neq s_{2}$ can also be considered.
+While such cross-stratum estimands are in principle identifiable under structural assumptions, they lack a meaningful causal interpretation.
+Following the definition of causal effects in Frangakis and Rubin (2002), the comparison of potential outcomes must be made within a well-defined target population.
+If $s_{1}\neq s_{2}$, then the two potential outcomes $Y_{1}(1)$ and $Y_{2}(0)$ are now drawn from distinct subpopulations characterized by different joint potential intermediate outcomes.
+In this case, $\tau^{s_{1},s_{2}}_{h}$ represents a cross-world comparison across heterogeneous subgroups rather than a treatment effect within a common subpopulation, and thus does not conform to the spirit of the original principal stratification framework regarding subpopulation-specific causal effects (Frangakis and Rubin, 2002).
+As an example, when $D$ denotes the survival status in the truncation-by-death setup (Example 1), comparing $Y_{1}(1)$ among the protected (those who survive only under treatment) with $Y_{2}(0)$ among the always-survivors (those who would survive regardless of assignment) conflates two sources of difference: one due to the treatment condition and another due to inherent heterogeneity across subpopulations.
+This contrast, therefore, cannot be solely attributed to treatment and does not carry a causal interpretation.
+For this reason, we restrict attention to $\tau_{h}^{s}$ with $s_{1}=s_{2}=s$ to ensure that pairwise comparisons are drawn from the same principal stratum.
+Given $S_{1}=S_{2}=s$, we emphasize that the contrast function $h$ is evaluated on two independent replicates drawn from the same principal stratum.
+This pairwise construction in the causal estimand also motivates the definition of the estimable pairwise mean function in Section 3.1, which serves as the building block for our identification strategy and $U$-statistic estimators.
+That is, both the estimand and the estimable pairwise mean function require two independent replicates to be properly defined.
+To illustrate the PGCE estimand, we next present two examples, while leaving the technical details on constructing $U$-statistic estimators to Section 3.
+Importantly, when a nonlinear contrast function is used to define the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum.
+Example 1 (Probabilistic index under truncation by death)
 When $D$ represents the survival status (alive if $D=1$) that is ascertained before the measurement of the final outcome $Y$ (such as quality of life, which is only well-defined when $D=1$), there exist four principal strata based on the potential survival status.
 That is, $s=11$ represents the always-survivors who would survive until the measurement of the final outcome regardless of the assignment; $s=10$ and $s=01$ represent the protected and harmed, who would survive until the measurement of the final outcome only under the treatment condition and control condition, respectively; $s=00$ represents the never-survivors who would not survive until the measurement of the final outcome regardless of the assignment.
-Because the final outcome is oftentimes only well-defined without ambiguity when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al., 2009; Chen et al., 2024).
+Because the final outcome is oftentimes only well-defined when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al., 2009; Chen et al., 2024).
 In clinical research, Acion et al.
-(2006) argued that a more intuitive effect size can be measured by the probability index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$.
-Based on this contrast function, we can define the survivor probability index (SPI) as
+(2006) argued that a more intuitive effect size can be measured by the probabilistic index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$.
+Based on this contrast function, we can define the survivor probabilistic index (SPI) as
 $\mathrm{SPI}=\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=11\}.$ (3)
-Compared to the SACE, the SPI is less sensitive to the skewness of the underlying distributions of the potential outcomes and heteroskedasticity between conditions and outliers.
+Compared to the SACE, the SPI is less sensitive to the skewness of the distributions of the potential outcomes, heteroskedasticity, and outliers.
 In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg, 1971).
@@
 This estimand corresponds to a two-dimensional contrast function specified as $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$.
-If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{11,10,01,00\}$ refers to always-takers, compliers, defiers, and never-takers, respectively (Angrist et al., 1996).
+If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{10,01,00,11\}$ refers to always-takers, compliers, defiers, and never-takers, respectively.
 Therefore, the win ratio estimands can be extended to define the compliers win ratio (CWR) as
@@
 The CWR estimand represents the odds of a randomly selected unit from all-treated compliers having a more favorable outcome than one from the all-control compliers subpopulation.
-Alternatively, the complier win difference (CWD) or complier net benefit estimand on the difference scale is given as
+Alternatively, the complier win difference (CWD) or complier net benefit estimand is
 $\mathrm{CWD}=\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=S_{2}=10\}-\mathbb{P}\{Y_{1}(1)<Y_{2}(0)|S_{1}=S_{2}=10\}.$ (5)
-The win estimands within other principal strata can be defined similarly.
 In the context of noncompliance, Mao (2024) defined a local Mann-Whitney treatment effect estimand given by $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=10\}$, which is also a special case of the PGCE.
-Different from our approach, Mao (2024) used an instrumental variable approach and discussed point and interval identification of this estimand based on monotonicity and exclusion restriction.
+Mao (2024) employed an instrumental variable approach and discussed the point and interval identification of this estimand, based on monotonicity and the exclusion restriction.
 2.3 Identification assumptions
-Even in experimental studies, the PGCE estimands cannot be identified without additional assumptions because the pair of potential intermediate variables and that of potential final outcomes are never observed simultaneously.
 To identify the PGCE from the observed data, we require the following assumptions.
@@
 […]
-Assumption 1 rules out confounding between the treatment assignment and intermediate as well as final outcomes conditional on observed pre-treatment covariates.
+Assumption 1 rules out confounding between the treatment assignment and intermediate variable as well as final outcomes conditional on observed pre-treatment covariates.
 This assumption holds by design in randomized experiments.
-For observational studies, Assumption 1 can be satisfied if all confounders affecting the treatment assignment, the intermediate variable, and the final outcome are measured and included in the measured set of pre-treatment covariates, $X$.
+For observational studies, Assumption 1 can be satisfied if all confounders (common causes) affecting the treatment assignment and the intermediate variable, and those affecting the treatment assignment and the final outcome are measured and included in $X$.
 Assumption 2 (Monotonicity)
@@
 When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to active treatment.
-Assumption 3 (Mean principal ignorability)
+Assumption 3 (Pairwise mean principal ignorability)
 Given the same values of pre-treatment covariates for a pair of units ($X_{1}=x_{1}$ and $X_{2}=x_{2}$), the expected contrast function is exchangeable across certain pairs of principal strata.
@@
 $\begin{split}&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2},X_{1}=x_{1},X_{2}=x_{2}]\\ =~&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1}^{\prime},S_{2}=s_{2}^{\prime},X_{1}=x_{1},X_{2}=x_{2}]\end{split}$ (6)
-for (a) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{10}=\{(10,10),(11,10),(10,00),(11,00)\}$; (b) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{11}=\{(11,11),(10,11)\}$; (c) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{00}=\{(00,00),(00,10)\}$.
-Assumption 3 rules out unmeasured confounding between the principal strata membership and the final outcome, ensuring that the PGCE components can be directly estimated from the observed data.
-For example, under Assumptions 1-3, the conditional expectations in group $\mathcal{S}_{10}$ are all equal to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$.
-In other words, within the observed data stratum defined by $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the combinations of possible latent principal strata, $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$, are exchangeable with respect to the expected contrast.
-Similarly, the conditional expected contrasts within group $\mathcal{S}_{11}$ and $\mathcal{S}_{00}$ are all equal to estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$ and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, respectively.
-That is, with Assumption 3, each one of the subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only.
-Remark 1
-[…]
-Assumptions 2 and 3 are instrumental in achieving nonparametric identification of the PGCE; however, their applicability, in general, cannot be directly tested using observed data alone without prior knowledge.
-Although our primary focus is identifying and efficiently estimating the PGCE under these two structural assumptions, we provide a confounding function approach for sensitivity analyses under assumed violations of these two structural assumptions in Section S7 of the Online Supplement.
+if $(s_{1},s_{2})$ and $(s_{1}^{\prime},s_{2}^{\prime})$ belong to the same set defined by one of the following: (a) $\mathcal{S}_{10}=\{(10,10),(10,00),(11,10),(11,00)\}$; (b) $\mathcal{S}_{11}=\{(10,11),(11,11)\}$; (c) $\mathcal{S}_{00}=\{(00,10),(00,\allowbreak 00)\}$; and (d) $\mathcal{S}_{01}=\{(00,11)\}$.
+Intuitively, Assumption 3 rules out unmeasured confounding between principal stratum membership and the final outcome, thereby ensuring that the PGCE estimands can be identified from the observed data.
+Under Assumptions 1 - 3, for instance, the conditional expectations within group $\mathcal{S}_{10}$ all reduce to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$.
+In other words, within the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the various combinations of possible latent strata including $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$ are exchangeable with respect to the expected contrast.
+Similarly, the conditional expected contrasts within $\mathcal{S}_{11}$, $\mathcal{S}_{00}$, and $\mathcal{S}_{01}$ are all equal to the estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$, respectively (see Table 1(a) for a concrete mapping between the observed data strata and the latent principal strata).
+That is, with Assumption 3, each one of the observed subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only (Table 1(b)).
+Assumption 3 can be viewed as a generalization of the unitwise mean principal ignorability assumption in Jiang et al.
+(2022), since it reduces to that form in the linear contrast case h(u,v)=uvh(u,v)=u-v.
+For nonlinear contrasts, however, expectations cannot be interchanged within the contrast function, and thus there is no clear nesting relationship between the pairwise and unitwise versions of mean principal ignorability.
+Consider the win estimands under noncompliance (Example 2) with the contrast function $h(u,v)=\mathbbm{1}(u>v)$.
+Applying (LABEL:eq:pi-equation) to the first two pairs in $\mathcal{S}_{10}$ yields $P\{Y_{1}(1)>Y_{2}(0)|S_{1}=10,S_{2}=10,X_{1}=x_{1},X_{2}=x_{2}\}=P\{Y_{1}(1)>Y_{2}(0)|S_{1}=11,S_{2}=10,X_{1}=x_{1},X_{2}=x_{2}\}$.
+This assumption states that, conditional on the same pairwise covariate profiles, the probability that a treated individual has a more favorable outcome than a control individual is the same whether the pair consists of two compliers or of one complier and one always-taker.
+In other words, there do not exist additional unmeasured factors influencing the conditional probability of winning among pairs in $\mathcal{S}_{10}$.
+This is neither stronger nor weaker than the unitwise mean principal ignorability assumption, which requires, for example, $\mathbb{E}\{Y_{1}(1)|S_{1}=10,X_{1}=x_{1}\}=\mathbb{E}\{Y_{1}(1)|S_{1}=11,X_{1}=x_{1}\}$.
+Table 1: Correspondence between latent principal strata pairs, Assumption 3, and observed strata pairs defined by assignment and observed intermediate variable.
+(a) The mapping from latent principal strata pairs to the conditional pairwise mean defined in observed strata, through Assumption 3. There are nine latent strata pairs under monotonicity.
+Latent strata pairs Conditions in Pairwise mean defined in observed strata
+$S_{1}$ $S_{2}$ Assumption 3 $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1},D_{1},Z_{2},D_{2},X_{1},X_{2}]$
+10 10 (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
+10 $0$ (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
+11 10 (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
+11 $0$ (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
+10 11 (b) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=1\}$
+11 11 (b) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=1\}$
+$0$ 10 (c) $\{Z_{1}=1,D_{1}=0\}$ & $\{Z_{2}=0,D_{2}=0\}$
+$0$ $0$ (c) $\{Z_{1}=1,D_{1}=0\}$ & $\{Z_{2}=0,D_{2}=0\}$
+$0$ 11 (d) $\{Z_{1}=1,D_{1}=0\}$ & $\{Z_{2}=0,D_{2}=1\}$
+(b) The mapping from observed strata pairs (defined by assignment and observed intermediate variable) as a mixture of latent principal strata pairs, through Assumption 3.
+As the estimand is constructed by contrasting $Y_{1}(1)$ with $Y_{2}(0)$, only four observed strata pairs that satisfy $Z_{1}=1$ and $Z_{2}=0$ are informative for identification.
+Observed strata pairs Conditions in Mixture components of
+Unit 1 Unit 2 Assumption 3 latent strata pairs
+$\{Z_{1}=1,D_{1}=1\}$ $\{Z_{2}=0,D_{2}=0\}$ (a) $(S_{1},S_{2})\in\mathcal{S}_{10}$
+$\{Z_{1}=1,D_{1}=1\}$ $\{Z_{2}=0,D_{2}=1\}$ (b) $(S_{1},S_{2})\in\mathcal{S}_{11}$
+$\{Z_{1}=1,D_{1}=0\}$ $\{Z_{2}=0,D_{2}=0\}$ (c) $(S_{1},S_{2})\in\mathcal{S}_{00}$
+$\{Z_{1}=1,D_{1}=0\}$ $\{Z_{2}=0,D_{2}=1\}$ (d) $(S_{1},S_{2})\in\mathcal{S}_{01}$
+Finally, we distinguish Assumptions 1 and 2 from Assumption 3.
+The first two are stated at the individual level, but since the observed data are assumed to be i.i.d., they extend naturally to pairs.
+For example, individual-level treatment ignorability implies that $(Z_{1},Z_{2})$ is jointly independent of the potential outcomes and intermediate variables given $(X_{1},X_{2})$.
+By contrast, Assumption 3 must be formulated at the pairwise level in its minimally sufficient form, because identification with nonlinear contrasts requires linking expectations across observed and latent strata pairs as in Table 1.
+Assumptions 2 and 3 are not testable from the observed data alone.
+We therefore complement our main results with sensitivity analyses under violations of these assumptions, provided in Section 4.4 and Section S7 of the Online Supplement.
 3 Nonparametric identification
 3.1 Identification of the PGCE estimands
-To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution, i.e., the treatment probability, the principal score, and the mean of the outcome contrast conditional on covariates (referred to as pairwise outcome mean hereafter for brevity).
-We consider both the observational and experimental studies and define the treatment probability or the propensity score, by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin, 1983).
+To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution.
+We consider both the observational and experimental studies and define the treatment propensity score by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin, 1983).
 In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu, 2016), for $s\in\{10,00,11\}$ under Assumption 2.
-Furthermore, we use averages over the distribution of covariates, $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$, to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively.
-Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as
-$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\},$
-for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Below, we provide some examples of the pairwise outcome mean models for specific contrast functions.
-Example 3 (Pairwise mean function for average treatment effect)
-For studying the average treatment effect, the contrast function is specified as $h(u,v)=u-v$.
-If the potential outcomes are continuous, a Gaussian regression model is often assumed for the potential outcomes, i.e., $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$.
-Here, $f_{z}(D,X)$ is the mean function of $D$ and $X$ given treatment status $Z=z$.
-We then have the pairwise outcome mean function given by
-$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2}).$ (7)
-Example 4 (Pairwise mean function for probability index)
-For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$.
-If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3.
-In this case, following Wolfe and Hogg (1971), one can derive the pairwise outcome mean as
+Furthermore, we use $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$ to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively.
+Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\}$, for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$.
+Below, we provide two examples of the pairwise outcome mean models for specific contrast functions.
+Example 3 (Pairwise mean function for probabilistic index)
+For studying the probabilistic index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$.
+If the potential outcomes are continuous, a Gaussian regression model can be considered such that $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$.
+In this case, one can derive the pairwise outcome mean as
 $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\Phi\left\{\frac{f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2})}{\sqrt{2}\sigma}\right\}.$ (7)
 Example 4 (Pairwise outcome mean for win estimands)
-Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand).
+Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$.
 For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
 $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ (8)
-=$\quad=\left\{\sum_{q=1}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2}),\sum_{q=1}^{Q}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2})\right\},$
-where $p_{zq}(D,X)$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $X$, $Q$ is the largest category of the ordinal outcome.
+$\quad=\left\{\sum_{q=2}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2}),\sum_{q=1}^{Q-1}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2})\right\},$
+where $p_{zq}(D,X)$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $X$, and $Q(\geq 2)$ is the largest category of the ordinal outcome.
 For estimating the principal scores, we first define the conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$.
-Then under Assumption 2, the principal scores can be expressed as (Jiang et al., 2022)
-$e_{10}(X)=p_{1}(X)-p_{0}(X),~{}~{}~{}e_{00}(X)=1-p_{1}(X),~{}~{}~{}e_{11}(X)=p_{0}(X).$
-This is because, with monotonicity, the observed stratum $(Z=0,D=1)$ only contains stratum $S=11$, and observed stratum $(Z=1,D=0)$ only contains stratum $S=00$.
-The remaining principal score for stratum $S=10$ can thus be obtained by subtracting $e_{11}(X)$ and $e_{00}(X)$ from one.
+Then, under Assumption 2, the principal scores can be expressed as $e_{10}(X)=p_{1}(X)-p_{0}(X)$, $e_{00}(X)=1-p_{1}(X)$, and $e_{11}(X)=p_{0}(X)$.
 The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$.
 We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$.
-These elements then prepare us to state the set of nonparametric identification results in the following Theorem 1.
-Theorem 1 (Nonparametric identification)
-Suppose Assumptions 1-3 hold, $e_{s}>0$ for $s\in\{10,00,11\}$, and $10<\pi(X)<$ 1 for all $X\in\mathcal{X}$, the support of covariates.
-Then, based on different components of the observed data distribution, the following identification formulas for each PGCE can be obtained:
-  • (a)
-    Based on the treatment probability and the principal score,
-    $\tau_{h}^{10}$ $=\mathbb{E}\left[\left\{\frac{e_{10}(X_{1})}{p_{1}-p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{10}(X_{2})}{p_{1}-p_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right],$
-    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-D_{1}}{1-p_{1}}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{00}(X_{2})}{1-p_{1}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right],$
-    $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{e_{11}(X_{1})}{p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{D_{2}}{p_{0}}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right];$
-  • (b)
-    Based on the treatment probability and the pairwise outcome mean,
-    $\tau_{h}^{10}$ $=\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
-    ×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2})-D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{1}-p_{0}}\right]\mu_{1100}(X_{1},X_{2})\right),$
-    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-D_{1}Z_{1}/\pi(X_{1})}{1-p_{1}}\right\}\left\{\frac{1-D_{2}Z_{2}/\pi(X_{2})}{1-p_{1}}\right\}\mu_{1000}(X_{1},X_{2})\right],$
-    $\tau_{h}^{11}$ $=\mathbb{E}\left(\left[\frac{D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{0}}\right]\left[\frac{D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{0}}\right]\mu_{1101}(X_{1},X_{2})\right);$
-  • (c)
-    Based on the principal score and the pairwise outcome mean,
-    $\tau_{h}^{10}$ $=\mathbb{E}\left[\left\{\frac{p_{1}(X_{1})-p_{0}(X_{1})}{p_{1}-p_{0}}\right\}\left\{\frac{p_{1}(X_{2})-p_{0}(X_{2})}{p_{1}-p_{0}}\right\}\mu_{1100}(X_{1},X_{2})\right],$
-    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-p_{1}(X_{1})}{1-p_{1}}\right\}\left\{\frac{1-p_{1}(X_{2})}{1-p_{1}}\right\}\mu_{1000}(X_{1},X_{2})\right],$
-    $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{p_{0}(X_{1})}{p_{0}}\right\}\left\{\frac{p_{0}(X_{2})}{p_{0}}\right\}\mu_{1101}(X_{1},X_{2})\right].$
-We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 1, and the identification of the other two PGCEs follows the same strategy.
-In Theorem 1(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$.
-Specifically, under Assumptions 1 and 2, we have
-$\mathbb{E}\left[\left\{\frac{e_{10}(X_{1})}{p_{1}-p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{10}(X_{2})}{p_{1}-p_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}\bigg{\rvert}X_{1},X_{2}\right]=\frac{e_{10}(X_{1})e_{10}(X_{2})}{e_{10}^{2}},$
-which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$.
-This ensures the identification of $\tau_{h}^{10}$.
-In Theorem 1(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
-$\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
-×$\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2})-D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{1}-p_{0}}\right]\bigg{\rvert}X_{1},X_{2}\right)=\frac{e_{10}(X_{1})e_{10}(X_{2})}{e_{10}^{2}}.$
-Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$.
-Finally, in Theorem 1(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 1(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$.
-The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.
-Remark 2
-Theorem 1 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al., 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al.
-(2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
+Based on these elements, we establish three alternative nonparametric identification formulas for each PGCE estimand, as shown in Theorem S2 of the Online Supplement.
+Each formula requires specifying only two of the three models—the treatment propensity score, the principal score, and the pairwise outcome mean—rather than all three simultaneously.
+Furthermore, Theorem S1 of the Online Supplement presents the associated balancing conditions for any pairwise covariate functions.
+These balancing conditions provide the theoretical foundation for model checking of the treatment propensity and principal score models without peeking at the outcome data.
 3.2 Estimators based on nonparametric identification formulas
-For point estimation, a common practice for operationalizing the nonparametric identification formulas is to assume parametric models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$.
+For point estimation, a common practice for implementing the nonparametric identification formulas is through specifying parametric regression models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$.
 We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$.
-To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment probability, the principal score, and the outcome mean are correctly specified, respectively.
-Specifically, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$.
-Additionally, we use ‘+’ in the subscript to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
+To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment propensity, the principal score, and the outcome mean models are correctly specified, respectively.
+That is, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$.
+Additionally, we use the logic notation ‘+’ to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
 We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts and expectations with empirical averages.
 For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$.
-Along the lines in Jiang et al.
-(2022), we define the following quantity for any function $f(Y,D,X)$:
+Define the following quantity for any function $f(Y,D,X)$:
+$\begin{aligned} \psi_{f(Y_{z},D_{z},X)}(\mathcal{O})&=\frac{\mathbbm{1}(Z=z)[f(Y,D,X)-\mathbb{E}\{f(Y,D,X)|X,Z=z\}]}{\mathbb{P}(Z=z|X)}\\ &\qquad+\mathbb{E}\{f(Y,D,X)|X,Z=z\}\end{aligned}.$ (9)
+One example of (9) is $f(Y_{z},D_{z},X)=D_{z}$, in which case we obtain the following function:
+$\psi_{D_{z}}(\mathcal{O})=\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X)\}}{\mathbb{P}(Z=z|X)}+p_{z}(X).$ (10)
+(10) is useful for the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome:
 […]
-One example of (10) is having $f(Y_{z},D_{z},X)=D_{z}$, in which case we obtain the following function:
-$\psi_{D_{z}}(\mathcal{O})=\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X)\}}{\mathbb{P}(Z=z|X)}+p_{z}(X).$ (11)
-Equation (11) is useful in the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome (Tsiatis, 2006):
-$\widehat{p}_{z}\equiv\mathbb{E}_{n}\psi_{D_{z}}(\mathcal{O};\widehat{\alpha},\widehat{\beta})=\mathbb{E}_{n}\left[\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X;\widehat{\beta})\}}{\pi(X;\widehat{\alpha})}+p_{z}(X;\widehat{\beta})\right],$ (12)
-which improves upon the plug-in estimator $\mathbb{E}_{n}p_{z}(X;\widehat{\beta})$ in terms of estimation efficiency and robustness, and is consistent for $p_{z}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$.
-Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart, 1998).
-For example, Theorem 1(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
-$\widehat{\tau}_{h,\pi+p}^{10}=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{1}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right].$
-Theorem 1(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
+[…]
+For any general constrast function $h$, the plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart, 1998, §12).
+For example, the identification results in Theorem S2 of the Online Supplement motivate the following estimators for $\tau_{h}^{10}$:
+$\widehat{\tau}_{h,\pi+p}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right],$
 $\widehat{\tau}_{h,\pi+\mu}^{10}$ $=\mathbb{U}_{n}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1};\widehat{\alpha})-D_{1}(1-Z_{1})/\{1-\pi(X_{1};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\right$.
-×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2};\widehat{\alpha})-D_{2}(1-Z_{2})/\{1-\pi(X_{2};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right),$
-and Theorem 1(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
-$\widehat{\tau}_{h,p+\mu}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{p_{1}(X_{1};\widehat{\beta})-p_{0}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\left\{\frac{p_{1}(X_{2};\widehat{\beta})-p_{0}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right].$
-The constructions of plug-in estimators for the remaining PGCEs are similar and, therefore, omitted.
-These estimators have stronger requirements on model specification to estimate the PGCEs consistently.
-Specifically, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment probability and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
+[…]
+[…]
+Upon closer investigation, $\widehat{\tau}_{h,\pi+p}^{10}$ is based on the estimated treatment propensity score and the principal score, $\widehat{\tau}_{h,\pi+\mu}^{10}$ is based on the estimated treatment propensity score and the pairwise outcome mean, and $\widehat{\tau}_{h,p+\mu}^{10}$ is based on the estimated principal score and the pairwise outcome mean.
+For brevity, we will use $\tau_{h}^{10}$ as the running example for exposition in the remainder of the article.
+Results for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are analogous and are given in the middle of Section S2 of the Online Supplement.
+These estimators have strong requirements on model specification to consistently estimate the PGCEs.
+To reiterate, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment propensity and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
 Remark 2
-The weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score, which is an extension of the inverse propensity score weighted estimator in Mao (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable.
-The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately after applying the Taylor’s expansion on the ratio estimators (Jiang et al., 2022); see the Online Supplement for additional details.
-The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be straightforwardly obtained using empirical process theory for the parametric class (van der Vaart, 1998, §19.2).
+It is worth noting that our weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score for pairs, and can be viewed as an extension of the pairwise inverse propensity score weighted estimator in Mao (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable.
+The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately; See the Online Supplement for additional details.
+The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be obtained using empirical process theory for the parametric class (van der Vaart, 1998, §19.2).
 The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao (2017).
-4 Multiply robust and efficient $U$-estimators
-Although implementing the plug-in estimators motivated by the nonparametric identification formulas is straightforward, their various forms and potential inefficiency call for alternative estimators with improved statistical properties.
-Here, we develop alternative $U$-estimators that exploit information across all nuisance functions, including the treatment assignment, the intermediate variable, and the final outcome of interest.
-4.1 The efficient influence function for the PGCE
-We first develop more efficient estimators with root-$n$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al., 1993).
+4 Multiply robust and efficient $U$-statistics
+4.1 The efficient influence function for the PGCE estimand
+We first develop more efficient estimators with $\sqrt{n}$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al., 1993).
 The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators.
 Several key properties of the EIF suggest its promise for constructing optimal estimators.
-First, the EIF often integrates more than one nuisance model and facilitates estimators that are robust against misspecification of some nuisance models (Tsiatis, 2006; Mao, 2017).
-Second, the asymptotic variance lower bound is represented by the variance of the EIF, providing a useful benchmark to define efficient estimators.
-Third, the EIF provides a device for integrating advanced machine learners for the nuisance parameters, which can lead to feasible estimators that achieve the asymptotic variance lower bound (Chernozhukov et al., 2018; Rotnitzky et al., 2021).
-We will explore these properties by first formalizing the EIF for the PGCE.
+First, the EIF often integrates multiple nuisance models and facilitates estimators that are robust against partial model misspecification.
+Second, the lower bound of the asymptotic variance is the variance of the EIF, providing a benchmark to define efficient estimators.
+Third, the EIF provides a device for incorporating data-adaptive machine learners for the nuisance parameters, suggesting feasible estimators that achieve the asymptotic variance lower bound with $\sqrt{n}$ convergence rate (Chernozhukov et al., 2018).
+We explore these properties by first formalizing the EIF for each PGCE.
 Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(X_{1})e_{10}(X_{2})\allowbreak\mu_{1100}(X_{1},X_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$.
@@
 We can further define marginalized functions of the observed outcome, $\nu_{1,zd}^{s}(Y)=\int e_{s}(x)h(Y,y)\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$ and $\nu_{0,zd}^{s}(Y)=\int e_{s}(x)h(y,Y)\allowbreak\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$, for $z,d\in\{0,1\}$ and $s\in\{10,00,11\}$.
-Then, the following results are obtained.
+Then, the following Theorem 1 is obtained.
 Theorem 1 (EIF)
-Under Assumptions 1-3, (a) the EIF for $\tau_{h}^{10}$ is $\varphi_{h}^{10}(\mathcal{O})=\{\varphi_{h,N}^{10}(\mathcal{O})-\tau_{h}^{10}\varphi_{D}^{10}(\mathcal{O})\}/(p_{1}-p_{0})^{2}$, where $\varphi_{h,N}^{10}(\mathcal{O})=\phi_{h}^{10}(\mathcal{O})-2\tau_{h,N}^{10}$ and
+Under Assumptions 1 - 3, the EIF for $\tau_{h}^{10}$ is $\varphi_{h}^{10}(\mathcal{O})=\{\varphi_{h,N}^{10}(\mathcal{O})-\tau_{h}^{10}\varphi_{D}^{10}(\mathcal{O})\}/(p_{1}-p_{0})^{2}$, where $\varphi_{h,N}^{10}(\mathcal{O})=\phi_{h}^{10}(\mathcal{O})-2\tau_{h,N}^{10}$ and
 $\phi_{h}^{10}(\mathcal{O})=$ $~\frac{ZDe_{10}(X)}{\pi(X)p_{1}(X)}\left\{\nu_{1,00}^{10}(Y)-m_{1}^{10}(X)\right\}+\frac{(1-Z)(1-D)e_{10}(X)}{\{1-\pi(X)\}\{1-p_{0}(X)\}}\left\{\nu_{0,11}^{10}(Y)-m_{0}^{10}(X)\right\}$
-+$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(X)+m_{0}^{10}(X)\right\},$
+$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(X)+m_{0}^{10}(X)\right\},$
 with $m_{1}^{10}(X)=\int e_{10}(x)\mu_{1100}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{10}(X)=\int e_{10}(x)\mu_{1100}(x,X)\mathrm{d}F_{X}(x)$, and
 $\varphi_{D}^{10}(\mathcal{O})=2\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}(p_{1}-p_{0})-2\tau_{D}^{10}.$
-(b) The EIF for $\tau_{h}^{00}$ is $\varphi_{h}^{00}(\mathcal{O})=\{\varphi_{h,N}^{00}(\mathcal{O})-\tau_{h}^{00}\varphi_{D}^{00}(\mathcal{O})\}/(1-p_{1})^{2}$, where $\varphi_{h,N}^{00}(\mathcal{O})=\phi_{h}^{00}(\mathcal{O})-2\tau_{h,N}^{00}$ and
-$\phi_{h}^{00}(\mathcal{O})=$ $~{}\frac{Z(1-D)}{\pi(X)}\left\{\nu_{1,00}^{00}(Y)-m_{1}^{00}(X)\right\}+\frac{(1-Z)(1-D)e_{00}(X)}{\{1-\pi(X)\}\{1-p_{0}(X)\}}\left\{\nu_{0,10}^{00}(Y)-m_{0}^{00}(X)\right\}$
-$+\psi_{1-D_{1}}(\mathcal{O})\left\{m_{1}^{00}(X)+m_{0}^{00}(X)\right\},$
-with $m_{1}^{00}(X)=\int e_{00}(x)\mu_{1000}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{00}(X)=\int e_{00}(x)\mu_{1000}(x,X)\mathrm{d}F_{X}(x)$, and
-$\varphi_{D}^{00}(\mathcal{O})=2\psi_{1-D_{1}}(\mathcal{O})(1-p_{1})-2\tau_{D}^{00}.$
-(c) The EIF for $\tau_{h}^{11}$ is $\varphi_{h}^{11}(\mathcal{O})=\{\varphi_{h,N}^{11}(\mathcal{O})-\tau_{h}^{11}\varphi_{D}^{11}(\mathcal{O})\}/p_{0}^{2}$, where $\varphi_{h,N}^{11}(\mathcal{O})=\phi_{h}^{11}(\mathcal{O})-2\tau_{h,N}^{11}$ and
-$\phi_{h}^{11}(\mathcal{O})=$ $~{}\frac{ZDe_{11}(X)}{\pi(X)p_{1}(X)}\left\{\nu_{1,01}^{11}(Y)-m_{1}^{11}(X)\right\}+\frac{(1-Z)D}{1-\pi(X)}\left\{\nu_{0,11}^{11}(Y)-m_{0}^{11}(X)\right\}$
-$+\psi_{D_{0}}(\mathcal{O})\left\{m_{1}^{11}(X)+m_{0}^{11}(X)\right\},$
-with $m_{1}^{11}(X)=\int e_{11}(x)\mu_{1101}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{11}(X)=\int e_{11}(x)\mu_{1101}(x,X)\mathrm{d}F_{X}(x)$, and
-$\varphi_{D}^{11}(\mathcal{O})=2\psi_{D_{0}}(\mathcal{O})p_{0}-2\tau_{D}^{11}.$
-The EIFs in Theorem 2 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey (2022) and Hines et al.
+The EIFs for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are given in Theorem S3 of Online Supplement Section S3.1.
+The EIFs for the PGCEs are derived following the chain rule via the method of Gâteaux derivatives as formalized in Ichimura and Newey (2022) and Hines et al.
 (2022) with respect to the distribution of observed data, $\mathcal{O}$.
-The derivation for $\varphi_{D}^{s}(\mathcal{O})$ then follows the chain rule.
-For $\varphi_{h,N}^{s}(\mathcal{O})$, we re-express the PGCE in terms of the observed data as
-$\tau_{h}^{s}=\mathbb{E}\{\mathfrak{h}_{1}^{s}(Y)|Z=1\}+\mathbb{E}\{\mathfrak{h}_{0}^{s}(Y)|Z=0\},$
-where $\mathfrak{h}_{1}^{s}(y)=\mathbb{E}\{h(y,Y^{\prime})|Z^{\prime}=0,S=S^{\prime}=s\}$ and $\mathfrak{h}_{0}^{s}(y)=\mathbb{E}\{h(Y^{\prime},y)|Z^{\prime}=1,S=S^{\prime}=s\}$, with the conditional expectations taken with respect to $Y^{\prime}$.
-Importantly, the EIF for the PGCE remains unchanged after this re-expression since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged.
-Based on Theorem 2, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
-Remark 4
-Without considering the intermediate variable, the EIF for the generalized causal effect estimand, $\tau_{h}$ in (1), $\varphi_{h}(\mathcal{O})$, can be derived using the same approach as $\varphi_{h}(\mathcal{O})=\phi_{h}(\mathcal{O})-2\tau_{h}$, where
-$\phi_{h}(\mathcal{O})=\frac{Z}{\pi(X)}\left\{\nu_{1}(Y)-m_{1}(X)\right\}+\frac{1-Z}{1-\pi(X)}\left\{\nu_{0}(Y)-m_{0}(X)\right\}+m_{1}(X)+m_{0}(X),$
-with $\nu_{z}(Y)=\int h(Y,y)\mathrm{d}F_{z}(y|x)\mathrm{d}F_{X}(x)$, $m_{1}(X)=\int\mu_{10}(X,x)\mathrm{d}F_{X}(x)$ and $m_{0}(X)=\int\mu_{10}(x,X)\allowbreak\mathrm{d}F_{X}(x)$.
-The function $\mu_{10}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=1,Z_{2}=0,X_{1},X_{2}\}$ is the pairwise outcome mean within the observed stratum $(Z_{1}=1,Z_{2}=0)$.
-This coincides with the EIF developed in Mao (2017) using the projection approach (Tsiatis, 2006), and intuitively, $\varphi_{h}(\mathcal{O})$ can be viewed as a special case of $\varphi_{h,N}^{s}(\mathcal{O})$ in the absence of the intermediate variable.
+Based on Theorems 1 and S3, the asymptotic variances of estimators for the PGCEs are lower bounded by the respective semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$ for $s\in\{10,00,11\}$.
 4.2 Multiply robust estimators
-The EIFs given in Theorem 2 imply the semiparametric efficiency lower bound and pave the way for formulating estimators for the PGCEs.
-However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator.
-We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 2, thereby achieving semiparametric efficiency with correctly specified nuisance functions.
-To this end, we first notice that each PGCE can be rewritten as the following:
-$\tau_{h}^{10}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{10})}{\{\mathbb{E}(\psi_{D_{1}}-\psi_{D_{0}})\}^{2}},~{}~{}~{}\tau_{h}^{00}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{00})}{\{\mathbb{E}(\psi_{1-D_{1}})\}^{2}},~{}~{}~{}\tau_{h}^{11}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{11})}{\{\mathbb{E}(\psi_{D_{0}})\}^{2}}.$ (13)
-Continuing with parametric models for nuisance functions, the denominators in (13) can be estimated using the doubly robust estimator given in (12).
-Specifically, let $\theta$ denote the collection of parametric nuisance function parameters and $\vartheta$ denote a subset of $\theta$, i.e., $\theta=\{\alpha,\beta,\gamma\}$ and $\vartheta=\{\alpha,\beta\}$, with corresponding maximum likelihood estimators $\widehat{\theta}=\{\widehat{\alpha},\widehat{\beta},\widehat{\gamma}\}$ and $\widehat{\vartheta}=\{\widehat{\alpha},\widehat{\beta}\}$, as well as true values $\theta^{*}=\{\alpha^{*},\beta^{*},\gamma^{*}\}$ and $\vartheta^{*}=\{\alpha^{*},\beta^{*}\}$.
-Then, for example, the denominator of $\tau_{h}^{10}$ given in (13) can be estimated by $[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}$, where $\mathbb{E}_{n}\psi_{D_{z}}(\widehat{\vartheta})\equiv\mathbb{E}_{n}\psi_{D_{z}}(\mathcal{O};\widehat{\alpha},\widehat{\beta})$ defined in (12).
-Estimating the numerators in (13) is more complicated. To target these terms, we construct the following functions of the observed pairs:
+The EIFs imply the semiparametric efficiency lower bound and pave the way for formulating optimal estimators for the PGCEs.
+However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and require care for constructing the final estimator.
+We use $U$-statistics in constructing the final estimator, which could be shown to have the same influence function given in Theorem 1, thereby achieving local semiparametric efficiency with correctly specified nuisance functions.
+To this end, we first notice that $\tau_{h}^{10}$ can be rewritten as
+$\tau_{h}^{10}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{10})}{\{\mathbb{E}(\psi_{D_{1}}-\psi_{D_{0}})\}^{2}}.$ (12)
+[…]
 $g_{h}^{10}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$
@@
 $\qquad\qquad+\frac{Z_{j}D_{j}e_{10}(X_{j};\beta)}{\pi(X_{j};\alpha)p_{1}(X_{j};\beta)}\times\frac{(1-Z_{i})(1-D_{i})e_{10}(X_{i};\beta)}{\{1-\pi(X_{i};\alpha)\}\{1-p_{0}(X_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1100}(X_{j},X_{i};\gamma)\right\}$
-+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(X_{i},X_{j};\gamma)$
-+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(X_{j},X_{i};\gamma)\bigg{]},$
-and
-$g_{h}^{00}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$
-$=2^{-1}\bigg{[}\frac{Z_{i}(1-D_{i})}{\pi(X_{i};\alpha)}\times\frac{(1-Z_{j})(1-D_{j})e_{00}(X_{j};\beta)}{\{1-\pi(X_{j};\alpha)\}\{1-p_{0}(X_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1000}(X_{i},X_{j};\gamma)\right\}$
-$\qquad\qquad+\frac{Z_{j}(1-D_{j})}{\pi(X_{j};\alpha)}\times\frac{(1-Z_{i})(1-D_{i})e_{00}(X_{i};\beta)}{\{1-\pi(X_{i};\alpha)\}\{1-p_{0}(X_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1000}(X_{j},X_{i};\gamma)\right\}$
-$\qquad\qquad+\psi_{1-D_{1}}(\mathcal{O}_{i};\alpha,\beta)\psi_{1-D_{1}}(\mathcal{O}_{j};\alpha,\beta)\mu_{1000}(X_{i},X_{j};\gamma)$
-$\qquad\qquad+\psi_{1-D_{1}}(\mathcal{O}_{j};\alpha,\beta)\psi_{1-D_{1}}(\mathcal{O}_{i};\alpha,\beta)\mu_{1000}(X_{j},X_{i};\gamma)\bigg{]},$
-and
-$g_{h}^{11}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ $=2^{-1}\bigg{[}\frac{Z_{i}D_{i}e_{11}(X_{i};\beta)}{\pi(X_{i};\alpha)p_{1}(X_{i};\beta)}\times\frac{(1-Z_{j})D_{j}}{1-\pi(X_{j};\alpha)}\left\{h(Y_{i},Y_{j})-\mu_{1101}(X_{i},X_{j};\gamma)\right\}$
-$\qquad\qquad+\frac{Z_{j}D_{j}e_{11}(X_{j};\beta)}{\pi(X_{j};\alpha)p_{1}(X_{j};\beta)}\times\frac{(1-Z_{i})D_{i}}{1-\pi(X_{i};\alpha)}\left\{h(Y_{j},Y_{i})-\mu_{1101}(X_{j},X_{i};\gamma)\right\}$
-$\qquad\qquad+\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\mu_{1101}(X_{i},X_{j};\gamma)$
-$\qquad\qquad+\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\mu_{1101}(X_{j},X_{i};\gamma)\bigg{]},$
-for $1\leq i\neq j\leq n$.
-For $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$, it is straightforward to show that $2\mathbb{E}\{g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})|\mathcal{O}_{i}\}=\phi_{h}^{s}(\mathcal{O}_{i})$ and $\mathbb{E}\{g_{h}^{s}(\mathcal{O}_{i},\allowbreak\mathcal{O}_{j};\theta^{*})\}=2^{-1}\mathbb{E}\{\phi_{h}^{s}(\mathcal{O}_{i})\}=\tau_{h,N}^{s}$.
-Thus, the numerators can be estimated using an empirical average of all pairs as $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})\equiv\binom{n}{2}^{-1}\sum_{i<j}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta})$ for $s\in\{10,00,11\}$.
-Therefore, we can obtain the following estimators for the PGCEs expressed in (13):
-$\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (14)
-which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 3.
-Theorem 3 (Multiply robustness and efficiency under parametric working models)
-Under Assumptions 1-3, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$.
-Then, the estimator $\widehat{\tau}_{h}^{s}$ in (14) is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
-Under Theorem 3, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified.
-The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator.
-The denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$.
-The consistency of the numerator is more involved and requires calculating the asymptotic bias of $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})-\tau_{h,N}^{s}$.
-The specific derivations are provided in the Online Supplement Section S4.1 due to the complexity of the expressions.
-The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao (2017) and Theorem 3 of Jiang et al.
-(2022).
-As discussed in Remark 3, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator.
-Details of the proof are provided in Section S4.2 of the Online Supplement.
-Finally, The variances of the triply robust estimators given in (14) can be estimated via nonparametric bootstrap.
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(X_{i},X_{j};\gamma)$
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(X_{j},X_{i};\gamma)\bigg],$
+[…]
+$\widehat{\tau}_{h,\mathrm{tr}}^{10}=\frac{\mathbb{U}_{n}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},$ (13)
+which is essentially a $U$-statistic but in the form of ratios.
+Estimators $\widehat{\tau}_{h,\mathrm{tr}}^{00}$ and $\widehat{\tau}_{h,\mathrm{tr}}^{11}$ are similarly constructed and defined in Section S4.1 of the Online Supplement.
+The asymptotic properties of these estimators are summarized in Theorem 2.
+Theorem 2 (Multiply robustness and efficiency)
+Under Assumptions 1 - 3, and suppose there exists some $\varepsilon\in(0,1)$ such that $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),\allowbreak p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$.
+Then, each estimator $\widehat{\tau}_{h,\mathrm{tr}}^{s}$, $s\in\{10,00,11\}$, is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
+Under Theorem 2, $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified.
+This triple robustness property of $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is obtained by separately investigating the properties of its numerator and the denominator.
+Specifically, the denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$.
+To account for the uncertainty in estimating the parametric nuisance functions, we estimate the variances of the triply robust estimators via a nonparametric bootstrap.
 4.3 Debiased machine learning estimators
-Debiased machine learning is a burgeoning method for efficient estimation of the causal effects in experimental and observation studies and is often built upon the orthogonal moment functions (Chernozhukov et al., 2018, 2022) or the EIFs.
+We propose further improvement of the parametric multiply robust estimator through debiased machine learning, which is similarly based on orthogonal moment functions (Chernozhukov et al., 2018) or EIFs, but allows efficient estimation of causal effects with data-adaptive nuisance estimators.
 Because the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al., 2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand.
-In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates (Rotnitzky et al., 2021).
-Compared to parametric working models, leveraging machine learning working models for the nuisance functions is more advantageous for its robustness against potential misspecification.
-To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting is adopted, where the entire sample is divided into separate parts for the estimation of the nuisance functions and the estimation of the causal estimand respectively (Chernozhukov et al., 2018).
-Mild regulatory conditions, mostly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy, 2023).
-Because of the unique structure of the PGCEs, we adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur (2023).
-Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition =$\mathcal{I}=\{I_{1},\ldots,I_{K}\}$.
+In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates.
+Compared to parametric working models, leveraging machine learning-based working models for the nuisance functions is more advantageous in terms of robustness against potential misspecification.
+To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting—is adopted.
+In this procedure, the entire sample is divided into separate parts for estimating the nuisance functions and the causal estimand, respectively (Chernozhukov et al., 2018).
+Mild regulatory conditions, mainly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy, 2023).
+We adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur (2022).
+Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition $\mathcal{I}=\{I_{1},\ldots,I_{K}\}$.
 Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $\{(i,j),1\leq i<j\leq n\}$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$.
@@
 (2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$.
-The following example is used to provide a demonstration of $\mathcal{I}$ and $\mathcal{P}$.
+We use the following example to demonstrate the definition of $\mathcal{I}$ and $\mathcal{P}$.
 Example 5 (Sample splitting for $U$-statistics)
 Consider a scenario where $n=10$ and $K=3$.
-We can first form the partition =$\mathcal{I}=\{I_{1},I_{2},I_{3}\}$ with $I_{1}=\{1,2,3\}$, $I_{2}=\{4,5,6,7\}$, and $I_{3}=\{8,9,10\}$.
+We can first form the partition $\mathcal{I}=\{I_{1},I_{2},I_{3}\}$ with $I_{1}=\{1,2,3\}$, $I_{2}=\{4,5,6,7\}$, and $I_{3}=\{8,9,10\}$.
 Then, the partition $\mathcal{P}$ is formed as
 $\{(I_{1},I_{1}),(I_{1},I_{2}),(I_{1},I_{3}),(I_{2},I_{2}),(I_{2},I_{3}),(I_{3},I_{3})\}.$
-For $\mathcal{P}$, elements formed by $I_{k_{1}}$ and $I_{k_{2}}$ with $k_{1}<k_{2}$ are more straightforward as the Cartesian product of $I_{k_{1}}$ and $I_{k_{2}}$. For instance,
-[…]
+For $\mathcal{P}$, elements formed by $I_{k_{1}}$ and $I_{k_{2}}$ with $k_{1}<k_{2}$ are straightforward Cartesian product of $I_{k_{1}}$ and $I_{k_{2}}$. For instance,
+$(I_{1},I_{2})=\{(1,4),(1,5),(1,6),(1,7),(2,4),(2,5),(2,6),(2,7),(3,4),\allowbreak(3,5),(3,6),(3,7)\}.$
 Elements formed by $I_{k}$ itself, on the other hand, need to exclude pairs consisting of the same indices, e.g., $(I_{1},I_{1})=\{(1,2),(1,3),(2,3)\}$.
@@
 For a given partition $\mathcal{P}$ (based on the initial partition of singletons $\mathcal{I}$), we obtain intermediate estimates of the target estimand using observations in each $P_{l}$ (or $I_{k}$) with $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$) estimated using observations in $P_{l}^{c}$ (or $I_{k}^{c}$), those are not present in $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$).
-By pooling these intermediate estimates together, we can obtain the following cross-fitting estimators for the PGCEs:
-$\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n,L}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n,K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n,L}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n,K}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n,L}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n,K}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (15)
-where
-$\mathbb{U}_{n,L}g_{h}^{s}(\widehat{\theta})=\binom{n}{2}^{-1}\sum_{l=1}^{L}\sum_{(i,j)\in P_{l}}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})$
-and
-$\mathbb{E}_{n,K}\psi_{D_{z}}(\widehat{\vartheta})=n^{-1}\sum_{k=1}^{K}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}).$
-Remark 5
-Estimators given in (15) are the type II estimator proposed in Chernozhukov et al. (2018) for estimands without an intermediate variable. Alternatively, one could first obtain separate estimates
-$\mathbb{U}_{|P_{l}|}g_{h}^{s}(\widehat{\theta}_{l})=|P_{l}|^{-1}\sum_{(i,j)\in P_{l}}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})~{}~{}~{}\text{and}~{}~{}~{}\mathbb{E}_{|I_{k}|}\psi_{D_{z}}(\hat{\vartheta}_{k})=|I_{k}|^{-1}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}),$
-on $P_{l}$ and $I_{k}$, respectively, where $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ are correspondingly estimated on $P_{l}^{c}$ and $I_{k}^{c}$. Then, define
-$\overline{\mathbb{U}}_{L}g_{h}^{s}(\widehat{\theta})=L^{-1}\sum_{l=1}^{L}\mathbb{U}_{|P_{l}|}g_{h}^{s}(\widehat{\theta}_{l})~{}~{}~{}\text{and}~{}~{}~{}\overline{\mathbb{E}}_{K}\psi_{D_{z}}(\widehat{\vartheta})=K^{-1}\sum_{k=1}^{K}\mathbb{E}_{|I_{k}|}\psi_{D_{z}}(\widehat{\vartheta}_{k}),$
-and the estimator for the PCEs can be formed as
-$\widehat{\tau}_{h}^{10}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{10}(\widehat{\theta})}{[\overline{\mathbb{E}}_{K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{00}(\widehat{\theta})}{\{\overline{\mathbb{E}}_{K}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{11}(\widehat{\theta})}{\{\overline{\mathbb{E}}_{K}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$
-which are the type I debiased machine learning estimators.
-We focus on the type II estimator because it is generally recommended for its numerical stability in pooling intermediate estimates based on split data, especially in smaller samples.
-We now discuss the asymptotic properties of the debiased machine learning estimators in (15).
-Assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability.
+By pooling these intermediate estimates together, we can obtain the following debiased machine learning estimator for $\tau_{h}^{10}$,
+$\widehat{\tau}_{h,\mathrm{dml}}^{10}=\frac{\mathbb{U}_{n,L}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n,K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},$ (14)
+where $\mathbb{U}_{n,L}g_{h}^{10}(\widehat{\theta})=\binom{n}{2}^{-1}\sum_{l=1}^{L}\sum_{(i,j)\in P_{l}}g_{h}^{10}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})$ and $\mathbb{E}_{n,K}\psi_{D_{z}}(\widehat{\vartheta})=n^{-1}\sum_{k=1}^{K}\sum_{i\in I_{k}}\allowbreak\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k})$.
+The debiased machine learning estimators $\widehat{\tau}_{h,\mathrm{dml}}^{00}$ and $\widehat{\tau}_{h,\mathrm{dml}}^{11}$ can be constructed in a similar fashion and are given in Section S5 of the Online Supplement.
+We now discuss the asymptotic properties of $\widehat{\tau}_{h,\mathrm{dml}}^{s}$.
+We first assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability.
 In addition, we assume true nuisance realizations $\theta^{*},\vartheta^{*}\in\Theta^{*}$.
-The following result summarizes the efficiency properties of (15).
-Theorem 4 (Efficiency under nonparametric working models)
+The following Theorem 3 summarizes the convergence and efficiency properties of the debiased machine learning estimators.
+Theorem 3 (Nonparametric efficiency)
 Under Assumptions 1 - 3, and
@@
   • (b)
-[…]
+    $\widehat{\theta}_{l}$ is consistent for $\theta$ and $\widehat{\vartheta}_{k}$ is consistent for $\vartheta$, with $\mathbb{E}|g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})-g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})|^{2}\xrightarrow{p}0$, and $\mathbb{E}|\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k})-\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})|^{2}\xrightarrow{p}0$, for all $l$ and $k$;
   • (c)
-[…]
+    $\|\widehat{\theta}_{l}-\theta^{*}\|=o_{\mathbb{P}}(n^{-1/4})$ and $\|\widehat{\vartheta}_{k}-\vartheta^{*}\|=o_{\mathbb{P}}(n^{-1/4})$, for all $l$ and $k$, where $\|\cdot\|$ denotes the $\mathcal{L}_{2}$-norm;
   • (d)
     $g_{h,1}^{s}(\mathcal{O};\theta)\equiv\mathbb{E}g_{h}^{s}(\mathcal{O},o;\theta)$ and $\psi_{D_{z}}(\mathcal{O};\vartheta)$ are $\mathbb{P}$-Donsker with probability approaching 1, and there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
-then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 2.
-Theorem 4 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al.
-(2022, Lemma 15) and Escanciano and Terschuur (2023, Lemma 4).
-To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$.
+$\widehat{\tau}_{h,\mathrm{dml}}^{s}$ is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the variance of the EIF.
+Theorem 3 establishes the consistency and asymptotic normality of the debiased machine learning estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al.
+(2022, Lemma 15) and Escanciano and Terschuur (2022, Lemma 4).
 Regulatory conditions in Theorem 3 are mostly mild.
 Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$.
-Condition (b) assumes that machine learning first-step estimators are consistent as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners.
+Condition (b) assumes that machine learning first-step estimators are consistent, as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners.
 Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al., 2018).
-The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator, which has been commonly adopted in debiased machine learning literature (Vansteelandt and Dukes, 2022; Wang et al., 2024).
+The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator; similar conditions have been previously invoked in the debiased machine learning literature (Vansteelandt and Dukes, 2022).
 Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (14) via nonparametric bootstrap.
+4.4 Sensitivity analysis for monotonicity
+To point identify PGCEs, a central assumption is monotonicity (Assumption 2) which rules out the $01$ stratum.
+This stratum may be relevant, for example, when comparing two active interventions with unknown survival benefits against non-mortality outcomes.
+We thus describe a sensitivity analysis strategy to evaluate the robustness of the proposed estimators in the event of departure from monotonicity.
+Following the approach in Ding and Lu (2016), we define the sensitivity function $\eta(X)$ as the ratio between the proportion of strata $01$ and $10$ conditional on the covariates, i.e., $\eta(X)=\mathbb{P}(S=01|X)/\mathbb{P}(S=10|X)$, where $\eta(X)\in[0,\infty)$; the monotonicity holds if $\eta(X)=0$, but $\eta(X)>0$ implies the existence of the $01$ stratum.
+The PCGEs in this scenario are thus $\tau_{h,\eta}^{s}$ for $s\in\{10,01,00,11\}$.
+To proceed, we first notice that, for $\eta(X)\neq 1$, the principal scores can be obtained as:
+$e_{10,\eta}(X)=\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},~~e_{01,\eta}(X)=\frac{\eta(X)\{p_{1}(X)-p_{0}(X)\}}{1-\eta(X)},$ (15)
+$e_{00,\eta}(X)=1-p_{0}(X)-\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},~~e_{11,\eta}(X)=p_{1}(X)-\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},$
+with $0\leq\eta(X)\leq 1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}$, and the proportions of principal strata by $e_{s,\eta}=\mathbb{E}\{e_{s,\eta}(X)\}$ for all $s$.
+When $\eta(X)=1$, strata $10$ and $01$ have equal proportions, which corresponds to the situation where the treatment has zero average causal effect on $S$ and the principal scores are not identifiable (Jiang et al., 2022, Section S4.3).
+Thus, this boundary case was ruled out.
+The identification of the PGCEs without monotonicity also requires an alternative version of the principal ignorability (Assumption S1), which we provide in Section S7.1 of the Online Supplement.
+With a fixed value of the sensitivity function $\eta(X)$, we derive bias-corrected multiply robust and debiased machine learning estimators for the PGCEs, leveraging the EIFs obtained without assuming monotonicity.
+The EIFs and the explicit forms of the bias-corrected estimators are given in Section S7.1 of the Online Supplement.
+Importantly, these multiply robust estimators remain triply robust under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$, as summarized in the following result (proofs provided in Section S7.2 of the Online Supplement).
+Theorem 4 (Multiply robustness and efficiency of bias-corrected estimators)
+Suppose that Assumptions 1 and S1 in Section S7.1 of the Online Supplement hold with a known $\eta(X)\neq 1$, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$.
+Then, for $s\in\{10,01,00,11\}$, the estimator $\widehat{\tau}_{h,\eta,\mathrm{tr}}^{s}$ is (a) triply robust such that it is consistent for $\tau_{h,\eta}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; and (b) asymptotically linear with influence function $\varphi_{h,\eta}^{s}$, thereby achieving the local semiparametric efficiency bound under $\mathcal{M}_{\pi+p+\mu}$.
 5 Simulation studies
-We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 1, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15).
-In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
-5.1 The probability index estimands with an intermediate variable
-[…]
-$D_{i}(Z_{i})|Z_{i},X_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\},$
-$Y_{i}(Z_{i})|D_{i},Z_{i},X_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1),$
-for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$.
-In this simulation study, we consider a binary intermediate variable ($D$ can be considered as the treatment receipt in the context of non-compliance), and assume the interest lies in estimating all three principal probability index estimands ($\tau_{h}^{10}$, $\tau_{h}^{00}$ and $\tau_{h}^{11}$ with $h(u,v)=\mathbbm{1}(u\geq v)$).
+We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for the PGCEs, including the three plug-in moment estimators motivated by the nonparametric identification results in Section S2 of the Online Supplement, the triply robust estimators with parametric working models given in (13), and the debiased machine learning estimators given in (14), under different data-generating processes.
+5.1 Probabilistic index with an intermediate variable
+In the first simulation study, we consider the case of the probabilistic index estimand introduced in Examples 1 and 3, where the potential outcomes are continuous and drawn from a normal distribution.
+We first generate four pre-treatment covariates $X_{i}=(X_{i1},X_{i2},X_{i3},X_{i4})^{\prime}$, where $X_{i1},X_{i2},X_{i3}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$, and $X_{i4}\overset{\text{i.i.d.}}{\sim}\mathcal{B}(0.5)$.
+Then, we draw the treatment assignment from $Z_{i}|X_{i}\sim\mathcal{B}\{\mathrm{expit}(-X_{i1}+0.5X_{i2}-0.25X_{i3}-0.1X_{i4})\}$.
+Finally, the potential intermediate variable and the potential outcome are drawn from $D_{i}(Z_{i})|Z_{i},X_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\}$, and $Y_{i}(Z_{i})|D_{i},Z_{i},X_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1)$ for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$.
+We consider a binary intermediate variable ($D$ can be considered as the treatment receipt in the context of non-compliance) and assume the interest lies in estimating all three principal probabilistic index estimands ($\tau_{h}^{10}$, $\tau_{h}^{00}$ and $\tau_{h}^{11}$ with $h(u,v)=\mathbbm{1}(u\geq v)$).
 In addition, we create transformed covariates, $\widetilde{X}_{i}=(\widetilde{X}_{i1},\widetilde{X}_{i2},\widetilde{X}_{i3},\widetilde{X}_{i4})^{\prime}$, where
@@
 Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise.
-Also, it is worth noting that, in practice, model misspecification pertains to the flexible debiased machine learning estimators because their estimation accuracy could be affected by the quality of the input covariate feature matrix.
-For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 4.
-For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries.
-We set the number of observations $n=2,000$ and consider $1,000$ replications; violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure 1 in Section S6 of the Online Supplement.
-Figure 1 presents the results of the five estimators under comparison in scenarios where, at most, one of the three nuisance functions is misspecified.
-Plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified.
-In contrast, the triply robust and debiased machine learning estimators are consistent across all four scenarios (correct specification of all nuisance functions or subsets of nuisance functions).
-Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure 1 in Section S6 of the Online Supplement.
+It is worth noting that, although in theory, machine learning methods are flexible with richer model spaces as $n$ increases, model misspecification can still be relevant to the debiased machine learning estimators, as their estimation accuracy in finite samples can be affected by the quality of the input covariate feature matrix (Cheng and Li, 2025).
+For the plug-in moment estimators and the triply robust estimator, treatment propensity scores and principal scores are estimated using logistic regression, whereas the pairwise outcome mean is estimated using linear regression, as in Example 3.
+For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear model libraries.
+We set the number of observations to $n=2,000$ and consider $1,000$ replications.
+Violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure S1 in Section S6 of the Online Supplement.
+Figure 1 summarizes the results of the five estimators in scenarios where, at most, one of the three nuisance functions is misspecified.
+Plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified.
+In contrast, the triply robust and debiased machine learning estimators are consistent across all four scenarios, regardless of whether all nuisance functions or subsets of nuisance functions are correctly specified.
+Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure S1 in Section S6 of the Online Supplement.
 As expected, the plug-in moment estimators are generally biased in those scenarios.
-In comparison to the plug-in moment estimators, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy.
-Finally, compared to the triply robust estimator, the debiased machine learning estimator leads to a slight improvement and is generally less biased.
+In contrast, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy.
+Finally, compared to the triply robust estimator, the debiased machine learning estimator yields a slight improvement, as it is generally less biased.
 (a) $\tau_{h}^{10}$
@@
 (c) $\tau_{h}^{11}$
-Figure 1: Violin plots of estimators for the principal probability index estimand in scenarios where at most one nuisance function is misspecified.
-Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15).
+Figure 1: Violin plots of estimators for the principal probabilistic index estimand in scenarios where at most one nuisance function is misspecified.
+Labels: ‘tp’ for treatment propensity or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment propensity (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (13); and ‘DML’ for the debiased machine learning estimators given in (14).
 Horizontal dashed lines represent the true values of the PGCEs.
-5.2 The win ratio estimands with an intermediate variable
-In the second simulation study, we consider the case of the win ratio estimands introduced in Examples 2 and 5, with the final potential outcomes being ordinal.
+5.2 Win ratio with an intermediate variable
+We further consider the case of estimating the win ratio estimands introduced in Examples 2 and 4, with the final potential outcomes being ordinal.
 We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study.
-However, for the potential outcome $Y_{i}(z)\in\{1,\ldots,Q\}$, we consider the following three-category ordinal logistic regression model:
-$\log\frac{\mathbb{P}(Y_{i}(Z_{i})\leq q)}{\mathbb{P}(Y_{i}(Z_{i})>q)}=\eta_{q,0}+2Z_{i}-D_{i}+X_{i1}-X_{i2}+1.2X_{i3}-0.8X_{i4},$
-for $q\in\{1,2\}$, where $\eta_{1,0}=-1$ and $\eta_{2,0}=1$.
-Transformed covariates are created following the procedure described in the previous simulation study.
-For the plug-in moment estimators and the triply robust estimator with parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are similarly estimated via logistic regression.
-For the debiased machine learning estimator, we used the multinomial generalized linear models with the elastic net penalty from the glmnet package in R to estimate the pairwise outcome mean, and the treatment propensity and principal scores are estimated using the SuperLearner package.
-The debiased machine learning estimator adopts a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
+However, for the potential outcome $Y_{i}(z)\in\{1,2,3\}$, we consider the following three-category ordinal logistic regression model:
+$\log\frac{\mathbb{P}\{Y_{i}(Z_{i})\leq q\}}{\mathbb{P}\{Y_{i}(Z_{i})>q\}}=\zeta_{q,0}+2Z_{i}-D_{i}+X_{i1}-X_{i2}+1.2X_{i3}-0.8X_{i4},$ (16)
+for $q\in\{1,2\}$, where $\zeta_{1,0}=-1$ and $\zeta_{2,0}=1$.
+Transformed covariates are created following Section 5.1.
+For the plug-in moment estimators and the triply robust estimator with parametric working models, the pairwise outcome mean is estimated using ordered logit regression, while treatment propensity and principal scores are estimated using logistic regression.
+For the debiased machine learning estimator, we used multinomial generalized linear models with the elastic net penalty from the glmnet package in R to estimate the pairwise outcome mean function.
+The treatment propensity and principal scores were estimated using the SuperLearner.
+The debiased machine learning estimator employs a five-fold cross-fitting procedure with $K=5$ and $L=15$.
 We set the number of observations $n=2,000$ with $1,000$ replications.
-Violin plots summarizing the simulation results of each considered estimator are given in Figure 2 and Web Figure 2 in Section S6 of the Online Supplement.
-From Figure 2, we observe that, similar to the estimation of the probability index estimand in Section 5.1, the plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified, while the triply robust and debiased machine learning estimators are consistent in all four scenarios where at most one of the three nuisance functions is misspecified.
-When at least two of the three nuisance functions are misspecified (shown in Web Figure 2 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the rest in terms of estimation accuracy (bias).
+In parallel to Section 5.1, violin plots summarizing the simulation results of each considered estimator are given in Figure 2 and Web Figure S2 in Section S6 of the Online Supplement.
+From Figure 2, we observe that the plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified.
+In contrast, the triply robust and debiased machine learning estimators are consistent in all four scenarios where at most one of the three nuisance functions is misspecified.
+When at least two of the three nuisance functions are misspecified (as shown in Web Figure S2 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the others in terms of estimation accuracy (bias).
 (a) $\tau_{h}^{10}$
@@
 Figure 2: Violin plots of estimators for the win ratio estimands in scenarios where at most one nuisance function is misspecified.
-Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15).
+Labels: ‘tp’ for treatment propensity or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment propensity (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (13); and ‘DML’ for the debiased machine learning estimators given in (14).
 Horizontal dashed lines represent the true values of the PGCEs.
+5.3 Additional simulations under non-monotonicity
+We consider a third simulation study to demonstrate the potential bias in the uncorrected triply robust and debiased machine learning estimators, as well as to validate the bias-corrected estimators under non-monotonicity.
+To achieve these goals, we modify the data-generating processes in Sections 5.1 and 5.2 by integrating the sensitivity function $\eta(X)$.
+Specifically, continuing with $p_{z}(X)=\mathrm{expit}(-1+2z+X_{1}-0.8X_{2}+0.6X_{3}-X_{4})$ for $z\in\{0,1\}$, we adopt the following configuration for the sensitivity function: $\eta(X)=\eta_{0}[1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}]$, where, we let $\eta_{0}\in[0,1]$ to ensure the true sensitivity function is well-defined within the permissible bound.
+We consider two scenarios: one with moderate violations, represented by $\eta_{0}=0.5$, and another with severe violations, represented by $\eta_{0}=0.99$.
+The strata membership $S$ is drawn from from a multinominal distribution $\mathcal{M}\mathcal{N}\{e_{10,\eta}(X),e_{01,\eta}(X),e_{00,\eta}(X),e_{11,\eta}(X)\}$, where $\{e_{s,\eta}(X),s\in\{10,01,00,11\}\}$ are obtained following (15).
+The treatment assignment $Z\sim\mathcal{B}\{\mathrm{expit}(-X_{1}+0.5X_{2}-0.25X_{3}-0.1X_{4})\}$, and the intermediate variable $D$ is determined once $Z$ and $S$ are known.
+That is, $D=1$ if $Z=1,S\in\{10,11\}$ or $Z=0,S\in\{01,11\}$, and $D=0$ if $Z=1,S\in\{01,00\}$ or $Z=0,S\in\{10,00\}$.
+The potential outcome $Y(z)$ for $z\in\{0,1\}$ is generated following (16), based on the same covariates simulated as in Section 5.1.
+To focus on the main ideas in this sensitivity simulation, we simplify without further considering nuisance function misspecification.
+We compare four estimators: the triply robust and debiased machine learning estimators under monotonicity, and the bias-corrected estimators with a known and correctly specified $\eta(X)$.
+The results are presented in Web Figure S3 in Section S6 of the Online Supplement.
+Overall, the simulation results show that the bias-corrected estimators give consistent estimates of all four PGCEs under both non-monotonicity scenarios.
+As $\eta_{0}$ increases from $0.5$ to $0.99$, the biases of the triply robust and debiased machine learning estimators increase accordingly, due to a higher degree of non-monotonicity.
+Interestingly, such increases in bias are not uniform across all PGCEs.
+For example, the biases of $\widehat{\tau}_{h,\mathrm{tr}}^{10}$ and $\widehat{\tau}_{h,\mathrm{dml}}^{10}$ are relatively small under both $\eta_{0}$ values, suggesting that they may be dramatically affected by violations of the monotonicity assumption in our setup.
+On the other hand, the biases of $\widehat{\tau}_{h,\mathrm{tr}}^{00}$, $\widehat{\tau}_{h,\mathrm{dml}}^{00}$, $\widehat{\tau}_{h,\mathrm{tr}}^{11}$, and $\widehat{\tau}_{h,\mathrm{dml}}^{11}$ increased significantly, indicating their relatively higher sensitivity to non-monotonicity.
+In further exploration, we find that as $\eta_{0}$ increases, a larger proportion of the $01$ stratum comes from the $00$ and $11$ strata, hence contributing to increased bias in the estimators for $\tau_{h}^{00}$ and $\tau_{h}^{11}$.
 6 Data example
@@
 Jobs Corps study (Schochet et al., 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24.
-The dataset (named ‘JC’ with $9,240$ observations, available in the R package causalweight) was from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, i.e., receiving an offer to participate in Job Corps immediately, or, the control group, i.e., not participating in Job Corps until three years later.
-Follow-up surveys were administered to both groups 24, 36, and 48 months after the random assignment to collect data on the outcomes of interest.
-As in many open-label randomized controlled trials, substantial noncompliance to treatment assignment as an intermediate outcome occurred 1 year after randomization (15.4% [$857/5,577$] of participants assigned to treatment did not adhere to the assignment, and 50.6% [$1,854/3,663$] assigned to control switched to the treatment condition).
-The proposed methods were implemented to examine the causal effect of Job Corps intervention, where the weekly earnings one year after the assignment were selected as the final outcome.
+The data were from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, which received an immediate offer to participate in Job Corps, or the control group, which did not participate in Job Corps until three years later.
+As in many open-label randomized controlled trials, substantial noncompliance with treatment assignment occurred as an intermediate outcome 1 year after randomization: 15.4% ($857/5,577$) of participants assigned to treatment did not adhere to the assignment, and 50.6% ($1,854/3,663$) assigned to control switched to the treatment condition.
+The proposed methods were implemented to examine the causal effect of the Job Corps intervention, with weekly earnings one year after the assignment selected as the final outcome.
 As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects.
-In each data analysis below, we adjust for baseline covariates, including baseline characteristics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
-6.1 The SPI for continuous outcomes
+In each data analysis below, we adjust for baseline covariates, including demographics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
+6.1 SPI for continuous outcomes
 We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$.
 In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al.
-(2009).
+(2009) for a similar setup.
 Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps.
@@
 For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear models libraries.
-The estimated results for SPI are given in Table 1, where $1,000$ bootstrap replicates were obtained for estimating the standard errors.
-We observe that, for the always-employed (72.9% of all participants by estimation), the probability index is estimated as $0.539$ using the triply robust and the debiased machine learning methods.
+The estimated results for SPI are presented in Table 2, where $1,000$ bootstrap replicates were generated to estimate the standard errors.
+We observe that, for the always-employed (72.9% of all participants by estimation), the probabilistic index is estimated to be $0.539$ using the triply robust and debiased machine learning methods.
 This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment.
-Finally, the point estimates under the plug-in moment methods are similar to the triply robust and debiased machine learning methods, which likely suggests that each parametric working model may be approximately correctly specified.
+Finally, the point estimates obtained under the plug-in moment methods are slightly closer to the null compared to those from the triply robust and debiased machine learning methods, although their 95% confidence intervals continue to exclude the null.
 Table 2: Estimated SPI for the Jobs Corps data.
 ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (13); and ‘DML’ for the debiased machine learning estimators given in (14).
-The corresponding 95% confidence interval (CI) is given in the following parenthesis.
+The corresponding 95% confidence interval (CI) is given in the following parentheses.
 Methods SPI estimates
@@
 DML 0.539 (0.527, 0.552)
-6.2 Principal win ratio for ordinal outcomes
+6.2 Principal win ratio (PWR) for ordinal outcomes
 We provide additional analysis of the Job Corps data to illustrate the application of the proposed methods to address ordinal outcomes with noncompliance.
-In this example, we use the actual treatment status as an intermediate variable, similar to Chen and Flores (2015).
+We use the actual treatment status as an intermediate variable, similar to Chen and Flores (2015).
 Under the monotonicity assumption, three strata potentially exist, including the compliers (stratum $10$), the never-takers (stratum $00$), and the always-takers (stratum $11$).
@@
 (2018).
-Under this setting, we estimate the principal win ratios within each principal strata similar to the definition in (4).
-For methods based on parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are estimated via logistic regression.
-For the debiased machine learning approach, we used the multinomial generalized linear models with the elastic net penalty to estimate the pairwise outcome mean, and the treatment propensity and principal scores are estimated using Super Learning; the estimator is fitted under a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
-The estimated results are given in Table 2, where $1,000$ bootstrap replicates were obtained for estimating the standard errors.
-The estimated proportions of always-takers, compliers, and never-takers in all participants are 50.1%, 34.0%, and 15.9%, respectively.
-The Jobs Corps intervention assignment positively affects participants’ earnings one year after the assignment for compliers and always-takers, and the effect appears to be more pronounced for compliers.
-For example, using the debiased machine learning estimator, there is a 27% increase in the chances of being in a higher earning category for a complier under treatment than a complier under control.
+Under this setting, we estimate the principal win ratios within each principal stratum, as defined in (4).
+For methods based on parametric working models, the pairwise outcome mean is estimated using ordered logit regression, while treatment propensity and principal scores are estimated using logistic regression.
+For the debiased machine learning approach, we used multinomial generalized linear models with the elastic net penalty to estimate the pairwise outcome mean.
+The treatment propensity and principal scores were estimated using the SuperLearner, fitted under a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
+The estimated results are presented in Table 3, where $1,000$ bootstrap replicates were generated to estimate the standard errors.
+The estimated proportions of always-takers, compliers, and never-takers among all participants are 50.1%, 34.0%, and 15.9%, respectively.
+The Jobs Corps intervention assignment has a positive effect on participants’ earnings one year after the assignment for both compliers and always-takers, and this effect appears to be more pronounced for compliers.
+For example, using the debiased machine learning estimator, there is a 27% increase in the chances of being in a higher earning category for a complier under treatment compared to a complier under control.
 The 95% lower confidence limits obtained from all five methods are strictly above the null value of 1.
@@
 Interestingly, in this reanalysis using a win ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable.
-In particular, the statistically significant effect among the always-takers is indicative of a positive psychological effect.
-That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence having a large chance to move into a higher earning category) compared to those who switched from their initial control assignment.
+The statistically significant effect among the always-takers is indicative of a positive psychological effect.
+That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence more likely to move into a higher earning category) compared to those who switched from their initial control assignment.
 Table 3: Estimated win ratios within different compliance strata for the Jobs Corps data.
 ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (13); and ‘DML’ for the debiased machine learning estimators given in (14).
-The corresponding 95% confidence interval is given in the following parenthesis.
+The corresponding 95% confidence interval is given in the following parentheses.
 Strata
@@
 DML 1.270 (1.146, 1.393) 1.041 (0.940, 1.142) 1.167 (1.082, 1.252)
+6.3 Sensitivity analyses for monotonicity
+We conducted additional sensitivity analyses to examine the impact of potential violations of the monotonicity assumption, focusing on both Sections 6.1 and 6.2.
+Following the approach in Section 5.3, we specified the sensitivity function as $\eta(X)=\eta_{0}[1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}]$, with $\eta_{0}\in\{0.1,0.2,0.3,0.4,0.5\}$.
+This formulation allows the sensitivity function to naturally vary with covariates while being governed by a one-dimensional strength parameter $\eta_{0}$.
+The chosen range of $\eta_{0}$ represents mild to moderate departures from monotonicity, which is plausible in the context of the Jobs Corps study, where encouragement for employment and subsequent treatment receipt may not dramatically deviate from monotonicity.
+We applied both the triply robust and the debiased machine learning estimators, with results reported in Figure 3.
+For the SPI analysis (Figure 3(a)), increasing $\eta_{0}$—corresponding to stronger departures from monotonicity—led to slightly larger estimated PGCEs.
+In other words, as the proportion of harmed strata increases, the estimated SPI shifts further from the null, yielding a stronger treatment effect signal, though statistical significance remains unaffected.
+In contrast, for the PWR analyses (Figures 3(b) - (d)), the estimated PGCEs are largely stable across values of $\eta_{0}$.
+Despite the presence of defiers, the estimated win ratios across all three principal strata remain nearly unchanged.
+The only noticeable pattern is a narrowing of the 95% confidence interval for the complier win ratio as $\eta_{0}$ increases.
+Collectively, these findings suggest that the PWR estimators exhibit strong robustness to non-monotonicity, in contrast to the slightly more pronounced sensitivity observed for the SPI.
+(a) SPI
+(b) $\mathrm{PWR}^{10}$
+(c) $\mathrm{PWR}^{00}$
+(d) $\mathrm{PWR}^{11}$
+Figure 3: Estimated SPI and PWRs and 95% confidence intervals under different values of sensitivity analysis parameter $\eta_{0}$.
+The horizontal axis of each plot shows values of $\eta_{0}$, and the vertical axis shows values of the estimated estimands and confidence intervals.
+SPI: the survivor probabilistic index in Section 6.1; $\mathrm{PWR}^{s}$: the principal win ratio of stratum $s$ in Section 6.2, with $s\in\{10,00,11\}$.
 7 Discussion
-In this article, we developed principal stratification methods for addressing a general class of PGCE estimands defined with a binary intermediate variable.
-Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability that are based on the contrast function, we contributed a semiparametric efficient approach for estimating the PGCE estimands by identifying their efficient influence functions and obtaining the ratio $U$-estimators that combine several different nuisance working functions.
-We showed that the $U$-estimators are triply robust in that they are consistent and asymptotically normal when two of the three nuisance functions are correctly specified.
-The $U$-estimators become locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric models or by machine learning models with cross-fitting.
-In light of the increasing interest in pairwise comparison estimands in studies with a composite endpoint (Buyse, 2010; Bebu and Lachin, 2015) in biomedical research, our methods can also be useful when there is a need to additionally address a binary intercurrent event, such as treatment noncompliance or death (Bornkamp et al., 2021; Lipkovich et al., 2022; Mao, 2024)
-The estimands and efficient $U$-estimators were developed to address a binary treatment and a binary intermediate outcome specifically but without restrictions on the final outcome.
-Although this is a standard setup for principal stratification analysis, additional work is needed to expand the identification conditions and results to address extensions of the PGCE estimands with a continuous treatment and/or a continuous intermediate variable (Antonelli et al., 2023).
-With continuous treatments, the definition of pairwise comparison estimands becomes non-trivial as there are infinitely many contrast functions (based on an infinitesimal increment of the treatment value) that one can specify to define $\tau_{h}$ even in the absence of an intermediate variable.
-On the other hand, with a continuous intermediate variable, there are infinitely many principal strata that can lead to conceptual and implementation challenges (Schwartz et al., 2011).
-Focusing on a linear contrast function, Lu et al.
-(2023) have developed nonparametric identification and semiparametric efficient estimators for principal causal effects with a continuous intermediate variable, which may inspire future theoretical development that allows for a more general nonlinear contrast function in the estimands definition.
-Finally, as in all principal stratification analyses that achieve point identification, our $U$-estimators for estimating PGCEs depend on the monotonicity and the mean principal ignorability assumptions that are unverifiable from the observed data alone.
-In Section S7 of the Online Supplement, we have described a sensitivity function approach to obtain the bias-corrected $U$-estimators for estimating the PGCEs when either the monotonicity assumption or the mean principal ignorability assumption is violated.
-The bias-corrected $U$-estimators remain triply robust to potential working model misspecifications and provide means to generate a range of point estimates under assumed departures from the key identification assumptions.
+In this article, we develop principal stratification methods for addressing a general class of PGCE estimands defined in terms of a binary intermediate variable.
+Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability that are based on the contrast function, we contributed a semiparametric efficient approach for estimating the PGCE estimands by identifying their efficient influence functions and obtaining the ratio $U$-statistic estimators that combine several different nuisance working functions.
+We showed that the proposed $U$-statistic estimators are triply robust and become locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric models or machine learning models with cross-fitting.
+As in all principal stratification analyses that achieve point identification, the proposed $U$-statistic estimators for estimating PGCEs also depend on the mean principal ignorability assumptions that are unverifiable from the observed data alone.
+In Section S7 of the Online Supplement, we have described an additional sensitivity function approach to obtain the bias-corrected $U$-statistic estimators for estimating the PGCEs when the mean principal ignorability assumption is violated.
+The bias-corrected estimators remain triply robust to potential working model misspecifications and provide means to generate a range of point estimates under assumed non-ignorability.
+Finally, our current development assumes that the observed data vectors are independent and identically distributed, providing a natural starting point.
+Extending these ideas to correlated data structures is an important direction for future research.
+For clustered data, recent work in cluster randomized trials has emphasized the need to define estimands that explicitly acknowledge the multilevel structure, particularly in the presence of informative cluster size (Kahan et al., 2023).
+A promising avenue is to expand the semiparametric and nonparametric methods of Wang et al.
+(2024) to accommodate cluster-average and individual-average PGCE estimands when the intermediate variable is measured at the individual level.
+For longitudinal data, both intermediate and final outcomes may be repeatedly measured over time, which requires reformulating the identification assumptions to incorporate sequential ignorability of treatment and multilevel monotonicity (Frangakis et al., 2004).
+It is also worth noting that, while our main exposition has focused on a single final outcome, our framework naturally extends to multiple hierarchically ordered outcomes—a setting that has gained considerable traction in clinical research over the past decade.
+Hierarchical composite endpoints can be viewed as clustered outcome vectors within individuals.
+In Section S8 of the Online Supplement, we extend the notation of Bebu and Lachin (2015) to the potential outcomes framework and show how our results apply to this setting by defining pairwise comparison estimands with hierarchical outcome components while maintaining the same binary intermediate variable.
+Further methodological developments along these directions will continue to broaden the applicability of our framework.
 Acknowledgement
-Research in this article was partially supported by the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (NHLBI, grant number R00-HL141678).
+Research in this article was partially supported by the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (NHLBI, grant number 1R01HL178513).
 All statements in this report, including its findings and conclusions, are solely those of the authors and do not necessarily represent the views of the NIH.

### v5

@@
 Principal stratification is a popular framework for causal inference in the presence of an intermediate outcome.
-While the principal average treatment effects have traditionally been the default target of inference, it may not be sufficient when the interest lies in the relative favorability of one potential outcome over the other within the principal stratum.
-We introduce the principal generalized causal effect estimands to extend the principal average causal effects, accommodating nonlinear contrast functions.
-Under principal ignorability, we expand the theoretical results in Jiang et al.
-(2022) to a much broader class of causal estimands in the presence of a binary intermediate variable.
-We develop identification results and derive the efficient influence functions of the generalized estimands for principal stratification analyses.
-These efficient influence functions motivate a set of multiply robust estimators and lay the ground for obtaining efficient debiased machine learning estimators via cross-fitting based on $U$-statistics.
+While the principal average treatment effects are the standard target of inference, they may be insufficient when interest lies in the relative ordering of potential outcomes within a principal stratum.
+We introduce the principal generalized causal effect estimands to accommodate nonlinear contrast functions, providing robust, probability-scale summaries suitable for ordinal outcomes and win-loss comparisons with composite endpoints.
+Under principal ignorability, we expand the theoretical results in Jiang2022 to a broader class of causal estimands in the presence of a binary intermediate variable.
+We develop nonparametric identification results and derive efficient influence functions for the generalized causal estimands in principal stratification analyses.
+These efficient influence functions motivate multiply robust estimators and lay the ground for obtaining efficient debiased machine learning estimators via cross-fitting based on U-statistics.
 The proposed methods are illustrated through simulations and the analysis of a data example.
@@
 1 Introduction
-The average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest in causal inference.
+The average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the standard estimand of interest in causal inference.
 Specifically, let $\{Y(1),Y(0)\}$ denote a pair of potential outcomes for a randomly selected individual from a super-population, and then the average causal effect estimand is defined as $\mathbb{E}\{Y(1)-Y(0)\}$.
-Conceptually, the average causal effect estimand can be expanded to a wider class, allowing arbitrary comparisons.
+Conceptually, the average causal effect estimand can be extended to a broader class, enabling arbitrary comparisons.
 That is, given a contrast function $h(u,v)$ defined on the product space, $\mathcal{Y}\times\mathcal{Y}$ ($\mathcal{Y}$: the potential outcome space), a generalized estimand can be defined as $\mathbb{E}[h\{Y(1),Y(0)\}]$.
 Examples include the average causal effect when h(u,v)=uvh(u,v)=u-v, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$.
-Because $Y(1)$ and $Y(0)$ are never observed simultaneously, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$ (see Lu et al.
-(2020) for an example for bounding such an estimand with ordinal outcomes).
+Because $Y(1)$ and $Y(0)$ are never observed simultaneously, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$ (see lu2020sharp for an example for bounding such an estimand with ordinal outcomes).
 However, a closely related quantity, the generalized causal effect, can be formulated:
@@
 Let $\{Y_{1}(1),Y_{1}(0)\}$ and $\{Y_{2}(1),Y_{2}(0)\}$ denote two independent replicates of $\{Y(1),Y(0)\}$, and $\tau_{h}$ can be re-expressed in a familiar form, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}]$.
-Two notable examples of the generalized causal effect are the probabilistic index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al., 2015), and the win ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al., 2011).
-These estimable quantities are alternative measurements for the population effect and carry a causal interpretation (Fay and Li, 2024).
-In certain cases, such estimands offer advantages that make them relevant in modern applications.
+Two notable examples of the generalized causal effect are the probabilistic index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (vermeulen2015increasing), and the win ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock2012).
+These estimable quantities are alternative measurements for the population effect and carry a causal interpretation (fay2024causal).
+In some instances, such estimands offer advantages that make them relevant in social science and biomedical applications.
 First, the probabilistic index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than a unit from $F_{0}$.
 This formulation generally exhibits greater robustness to distributional features.
-That is, unlike mean contrasts, which can be distorted by heavy tails, skewness, or outliers, probabilistic indices require fewer distribution assumptions and are less sensitive to irregularities in the outcome distribution (Hauck et al., 2000; Acion et al., 2006).
-Second, probabilistic indices are also well suited for ordinal outcomes, where the scientific question itself concerns relative ordering rather than pure numerical differences.
-Hence, as De Schryver and De Neve (2019) noted, “the probabilistic index provides an ordinal answer to an ordinal question”, and it relies only on the rank comparison rather than the metric scale or magnitude of differences.
-Third, the win-type estimand is now increasingly used for prioritized composite outcomes, where rank-based win-loss comparisons naturally align with clinical priorities across multiple endpoints.
-Introduced by Pocock et al.
-(2011), the win ratio has gained substantial traction in clinical trials with composite outcomes, as it offers a principled way to combine endpoints of varying clinical importance while providing an interpretable summary of treatment benefit in terms of the probability of winning versus losing.
-When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al., 1996) and death (Zhang et al., 2009) often have important implications on the definition of a causal effect and require careful adjustment.
-The principal stratification framework (Frangakis and Rubin, 2002) offers one solution by segmenting the population into principal strata (subpopulations) based on the joint potential values of the intermediate variable under alternative conditions.
+That is, unlike mean contrasts, which can be distorted by heavy tails, skewness, or outliers, probabilistic indices require fewer distribution assumptions and are less sensitive to irregularities in the outcome distribution (hauck2000generalized; acion2006probabilistic).
+Second, probabilistic indices are well suited to ordinal outcomes, in which the scientific question concerns relative ordering rather than pure numerical differences.
+As de2019tutorial noted, “the probabilistic index provides an ordinal answer to an ordinal question”, and it relies only on the rank comparison rather than the metric scale or magnitude of differences.
+Third, the win-type estimand is now increasingly used for prioritized composite outcomes, where win-loss comparisons naturally align with clinical priorities across multiple endpoints.
+Introduced by Pocock2012, the win ratio has gained substantial traction in clinical research with composite outcomes, as it offers a principled way to combine endpoints of varying clinical importance while providing an interpretable summary of treatment benefit in terms of the probability of winning versus losing.
+When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist1996) and death (Zhang2008) often have important implications on the definition of a causal effect and require careful adjustment.
+Principal stratification (Frangakis2002) offers one solution by segmenting the population into principal strata based on the joint potential values of the intermediate variable under alternative conditions.
 The stratum-specific causal effect, referred to as the principal causal effect, can thus be defined and is causally interpretable.
-However, a predominant focus in prior works has been on the average causal effect within each principal stratum, and there is a lack of related development when the generalized causal effect is of interest within each subpopulation.
-Therefore, in this article, we propose a new class of estimands, the principal generalized causal effect (PGCE), which accommodates an arbitrary and potentially nonlinear contrast function.
-As we demonstrate in due course, the PGCE can represent the stratum-specific probabilistic index and strata-specific win ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within subpopulations of interest.
-Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al., 1996) and principal ignorability (Ding and Lu, 2016), which are typically invoked for principal stratification analyses.
-The monotonicity assumption states that the treatment has a nonnegative causal effect on the intermediate outcome of each unit, restricting the number of potential strata.
-This condition is considered plausible when, for example, the intermediate variable is death, and the treatment is known to benefit survival, or when the intermediate variable is the actual treatment received and the treatment does not discourage treatment uptake.
-The principal ignorability assumption requires that the observed pre-treatment covariates are sufficient to control for confounding between the intermediate variable and the outcome.
-For the principal average causal effect, under monotonicity and principal ignorability, Ding and Lu (2016) formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang et al.
-(2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and developed multiply robust and locally efficient estimators; Cheng et al.
-(2023, 2024) extended the multiply robust and locally efficient estimators to estimate principal survival causal effect with a right-censored time-to-event outcome; Tong et al.
-(2024) expanded multiply robust and locally efficient estimators to estimate fine principal causal effects with multiple treatments.
+However, prior work has predominantly focused on the average causal effect within each principal stratum, and there has been little related development when the generalized causal effect is of interest within each subpopulation.
+Therefore, we propose a new class of estimands, the principal generalized causal effect (PGCE), that accommodates an arbitrary, potentially nonlinear contrast function.
+As we demonstrate in due course, the PGCE can represent the stratum-specific probabilistic index and the stratum-specific win ratio with an appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within subpopulations.
+Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist1996) and principal ignorability (Ding2016), which are typically invoked for principal stratification analyses.
+For the principal average causal effect, under monotonicity and principal ignorability, Ding2016 formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang2022 provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and developed multiply robust and locally efficient estimators; cheng2023multiply extended the multiply robust and locally efficient estimators to estimating principal survival causal effect with a right-censored time-to-event outcome; tong2024doubly expanded multiply robust and locally efficient estimators to estimate fine principal causal effects with multiple treatments.
 Despite these advancements, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
 The contributions of this article are several-fold.
-First, we propose a class of PGCE estimands in the presence of a binary intermediate variable and develop complementary nonparametric identification formulas under the assumptions of treatment unconfoundedness, monotonicity, and principal ignorability.
-Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands and seek efficiency improvement by deriving their corresponding efficient influence functions (Bickel et al., 1993, EIFs).
-Specifically, we show that when the contrast function is h(u,v)=uvh(u,v)=u-v, our EIFs reduce to those in Jiang et al.
-(2022) as a special case.
-More broadly, our framework accommodates pairwise comparison estimands with arbitrary contrast functions under principal stratification, offering a substantially more general treatment and establishing the present work as a standalone contribution.
-A major complexity for nonlinear contrast functions, however, is that the estimating equations based on the EIFs require nuisance functions defined over pairs of observations, thereby requiring the theory of $U$-statistics to construct feasible estimators.
+First, we propose a class of PGCE estimands in the presence of a binary intermediate variable and develop nonparametric identification formulas under the assumptions of treatment unconfoundedness, monotonicity, and principal ignorability.
+Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands and derive their corresponding efficient influence functions (EIFs, Bickle1993).
+Specifically, we show that when the contrast function is h(u,v)=uvh(u,v)=u-v, our EIFs reduce to those in Jiang2022 as a special case.
+More broadly, our framework accommodates pairwise comparison estimands with arbitrary contrast functions, offering a substantially more general treatment and establishing the present work as a standalone contribution.
+A significant complexity of nonlinear contrast functions, however, is that the estimating equations based on the EIFs require nuisance functions defined on pairs of observations, thereby necessitating the theory of U-statistics to construct feasible estimators.
 We show that the proposed U-statistics are triply robust in the sense that they are consistent and asymptotically normal when two of the three nuisance functions are correctly specified.
@@
 Finally, the construction of the multiply robust U-statistics facilitates the integration of nonparametric estimators for the nuisance functions.
-To achieve full asymptotic efficiency, we also estimate each nuisance function using data-adaptive machine learners under the cross-fitting scheme (Chernozhukov et al., 2018).
-We discuss the modifications needed for cross-fitting when a nonlinear contrast function is considered in defining the PGCE, and prove the nonparametric efficiency properties of the debiased machine learning estimators under mild regularity conditions.
-The rest of this article is organized as follows.
-Section 2 introduces our causal estimands and identification assumptions for principal stratification.
-Section 3 discusses the nonparametric identification formulas for the PGCE and associated moment estimators.
-Section 4 derives the EIF for each PGCE estimand and develops the EIF-motivated estimators.
-Section 5 demonstrates the finite-sample performance of the proposed estimators via simulations.
-Section 6 illustrates the proposed methods with a data example from the United States Jobs Corps study (Schochet et al., 2001), and Section 7 concludes.
-Detailed proofs of all theoretical results are presented in the Online Supplement.
+To achieve full asymptotic efficiency, we estimate each nuisance function using data-adaptive machine learners under the cross-fitting scheme (Chernozhukov2018).
+We discuss the modifications needed for cross-fitting when a nonlinear contrast function is used to define the PGCE, and prove the nonparametric efficiency properties of the debiased machine learning estimators under mild regularity conditions.
+Example R codes are available at https://github.com/erxc/cusps.
 2 Notation, estimands and assumptions
-2.1 Setting and notation
 We consider a study, either experimental or observational, with $n$ individuals and two treatment conditions.
@@
 Under the Stable Unit Treatment Value Assumption (SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$.
-Let $X_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{X_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d.).
+Let $\bm{X}_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{\bm{X}_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d.).
 2.1 Principal generalized causal effects
-We operate under the principal stratification framework (Frangakis and Rubin, 2002), and define distinct principal strata according to the joint potential values of the intermediate variable under alternative conditions.
+We operate within the principal stratification framework (Frangakis2002) and define distinct principal strata based on the joint potential values of the intermediate variable under alternative conditions.
 Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{10,01,00,11\}$.
@@
 $\tau_{h}^{s}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=S_{2}=s],~\text{for}~s\in\{10,01,00,11\},$ (2)
-where $Y_{1}(1)$ and $Y_{2}(0)$ denote the potential outcomes for two randomly selected units from the subpopulation with strata membership $s$ ($S_{1}=s$ and $S_{2}=s$ denote the strata membership for these two selected units).
-When a linear contrast function is specified as h(u,v)=uvh(u,v)=u-v, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature (Frangakis and Rubin, 2002).
-Before proceeding, we provide a caveat on the definition of estimands in Remark 1.
+where $Y_{1}(1)$ and $Y_{2}(0)$ denote the potential outcomes for two randomly selected units from the subpopulation with strata membership $s$.
+When a linear contrast function is specified as h(u,v)=uvh(u,v)=u-v, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature (Frangakis2002).
 Remark 1
@@
 While such cross-stratum estimands are in principle identifiable under structural assumptions, they lack a meaningful causal interpretation.
-Following the definition of causal effects in Frangakis and Rubin (2002), the comparison of potential outcomes must be made within a well-defined target population.
+Following the definition of causal effects in Frangakis2002, the comparison of potential outcomes must be made within a well-defined target population.
 If $s_{1}\neq s_{2}$, then the two potential outcomes $Y_{1}(1)$ and $Y_{2}(0)$ are now drawn from distinct subpopulations characterized by different joint potential intermediate outcomes.
-In this case, $\tau^{s_{1},s_{2}}_{h}$ represents a cross-world comparison across heterogeneous subgroups rather than a treatment effect within a common subpopulation, and thus does not conform to the spirit of the original principal stratification framework regarding subpopulation-specific causal effects (Frangakis and Rubin, 2002).
+In this case, $\tau^{s_{1},s_{2}}_{h}$ represents a cross-world comparison across heterogeneous subgroups rather than a treatment effect within a common subpopulation, and thus does not conform to the spirit of the original principal stratification framework regarding subpopulation-specific causal effects (Frangakis2002).
 As an example, when $D$ denotes the survival status in the truncation-by-death setup (Example 1), comparing $Y_{1}(1)$ among the protected (those who survive only under treatment) with $Y_{2}(0)$ among the always-survivors (those who would survive regardless of assignment) conflates two sources of difference: one due to the treatment condition and another due to inherent heterogeneity across subpopulations.
-This contrast, therefore, cannot be solely attributed to treatment and does not carry a causal interpretation.
-For this reason, we restrict attention to $\tau_{h}^{s}$ with $s_{1}=s_{2}=s$ to ensure that pairwise comparisons are drawn from the same principal stratum.
+This contrast, therefore, cannot be solely attributed to treatment and does not carry a straightforward causal interpretation.
+For this reason, we restrict attention to $\tau_{h}^{s}$ with $s_{1}=s_{2}=s$.
 Given $S_{1}=S_{2}=s$, we emphasize that the contrast function $h$ is evaluated on two independent replicates drawn from the same principal stratum.
-This pairwise construction in the causal estimand also motivates the definition of the estimable pairwise mean function in Section 3.1, which serves as the building block for our identification strategy and $U$-statistic estimators.
-That is, both the estimand and the estimable pairwise mean function require two independent replicates to be properly defined.
-To illustrate the PGCE estimand, we next present two examples, while leaving the technical details on constructing $U$-statistic estimators to Section 3.
+This pairwise construction in the causal estimand also motivates the definition of the estimable pairwise mean function in Section 3.1, which serves as the building block for our identification strategy and U-statistic-based estimators.
+To illustrate the PGCE estimand, we next present two examples, while leaving the technical details on constructing U-statistic-based estimators to Section 3.
 Importantly, when a nonlinear contrast function is used to define the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum.
@@
 That is, $s=11$ represents the always-survivors who would survive until the measurement of the final outcome regardless of the assignment; $s=10$ and $s=01$ represent the protected and harmed, who would survive until the measurement of the final outcome only under the treatment condition and control condition, respectively; $s=00$ represents the never-survivors who would not survive until the measurement of the final outcome regardless of the assignment.
-Because the final outcome is oftentimes only well-defined when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al., 2009; Chen et al., 2024).
-In clinical research, Acion et al.
-(2006) argued that a more intuitive effect size can be measured by the probabilistic index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$.
+Because the final outcome is typically only well-defined when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang2008; Chen2024).
+In clinical research, acion2006probabilistic argued that a more intuitive effect size can be measured by the probabilistic index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$.
 Based on this contrast function, we can define the survivor probabilistic index (SPI) as
 $\mathrm{SPI}=\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=11\}.$ (3)
-Compared to the SACE, the SPI is less sensitive to the skewness of the distributions of the potential outcomes, heteroskedasticity, and outliers.
-In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg, 1971).
+Compared with the SACE, the SPI is less sensitive to skewness in the distributions of potential outcomes, heteroskedasticity, and outliers.
+In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (wolfe1971constructing).
 Example 2 (Win estimands under noncompliance)
-In experimental studies with ordinal outcomes, Wang and Pocock (2016) introduced the win ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$.
+In experimental studies with ordinal outcomes, wang2016win introduced the win ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$.
 This estimand corresponds to a two-dimensional contrast function specified as $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$.
@@
 $\mathrm{CWR}=\frac{\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=S_{2}=10\}}{\mathbb{P}\{Y_{1}(1)<Y_{2}(0)|S_{1}=S_{2}=10\}}.$ (4)
-The CWR estimand represents the odds of a randomly selected unit from all-treated compliers having a more favorable outcome than one from the all-control compliers subpopulation.
+The CWR represents the odds that a randomly selected complier under the treatment condition has a more favorable outcome than under the control condition.
 Alternatively, the complier win difference (CWD) or complier net benefit estimand is
 $\mathrm{CWD}=\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=S_{2}=10\}-\mathbb{P}\{Y_{1}(1)<Y_{2}(0)|S_{1}=S_{2}=10\}.$ (5)
-In the context of noncompliance, Mao (2024) defined a local Mann-Whitney treatment effect estimand given by $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=10\}$, which is also a special case of the PGCE.
-Mao (2024) employed an instrumental variable approach and discussed the point and interval identification of this estimand, based on monotonicity and the exclusion restriction.
+In the context of noncompliance, Mao2024 defined a local Mann-Whitney treatment effect estimand given by $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=10\}$, which is also a special case of the PGCE.
+Mao2024 employed an instrumental variable approach and discussed the point and interval identification of this estimand under monotonicity and exclusion restriction.
 2.2 Identification assumptions
@@
 This assumption holds by design in randomized experiments.
-For observational studies, Assumption 1 can be satisfied if all confounders (common causes) affecting the treatment assignment and the intermediate variable, and those affecting the treatment assignment and the final outcome are measured and included in $X$.
+For observational studies, Assumption 1 can be satisfied if all confounders (common causes) affecting the treatment assignment and the intermediate variable, and those affecting the treatment assignment and the final outcome are measured and included in $\bm{X}$.
 Assumption 2 (Monotonicity)
@@
 Assumption 2 states that the treatment assignment only affects the intermediate outcome in a non-negative (or non-positive if assume $D_{i}(1)\leq D_{i}(0)$) direction and restricts the number of possible strata by excluding the harmed or defiers stratum $S=01$ (or $S=10$).
-For example, when $D$ represents the survival status, monotonicity holds when the treatment is expected not to harm the survival of each unit.
-When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to active treatment.
+For example, when $D$ represents the survival status, monotonicity holds if the treatment is not expected to harm the survival of each unit.
+When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to the active treatment.
 Assumption 3 (Pairwise mean principal ignorability)
-Given the same values of pre-treatment covariates for a pair of units ($X_{1}=x_{1}$ and $X_{2}=x_{2}$), the expected contrast function is exchangeable across certain pairs of principal strata.
+Given the same values of pre-treatment covariates for a pair of units ($\bm{X}_{1}=\bm{x}_{1}$ and $\bm{X}_{2}=\bm{x}_{2}$), the expected contrast function is exchangeable across certain pairs of principal strata.
 That is,
-$\begin{split}&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2},X_{1}=x_{1},X_{2}=x_{2}]\\ =~&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1}^{\prime},S_{2}=s_{2}^{\prime},X_{1}=x_{1},X_{2}=x_{2}]\end{split}$ (6)
+$\begin{split}&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2},\bm{X}_{1}=\bm{x}_{1},\bm{X}_{2}=\bm{x}_{2}]\\ =~&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1}^{\prime},S_{2}=s_{2}^{\prime},\bm{X}_{1}=\bm{x}_{1},\bm{X}_{2}=\bm{x}_{2}]\end{split}$ (6)
 if $(s_{1},s_{2})$ and $(s_{1}^{\prime},s_{2}^{\prime})$ belong to the same set defined by one of the following: (a) $\mathcal{S}_{10}=\{(10,10),(10,00),(11,10),(11,00)\}$; (b) $\mathcal{S}_{11}=\{(10,11),(11,11)\}$; (c) $\mathcal{S}_{00}=\{(00,10),(00,\allowbreak 00)\}$; and (d) $\mathcal{S}_{01}=\{(00,11)\}$.
-Intuitively, Assumption 3 rules out unmeasured confounding between principal stratum membership and the final outcome, thereby ensuring that the PGCE estimands can be identified from the observed data.
-Under Assumptions 1 - 3, for instance, the conditional expectations within group $\mathcal{S}_{10}$ all reduce to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$.
+Assumption 3 rules out unmeasured confounding between principal stratum membership and the final outcome, thereby ensuring that the PGCE estimands can be identified from the observed data.
+Under Assumptions 1 - 3, for instance, the conditional expectations within group $\mathcal{S}_{10}$ all reduce to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,\bm{X}_{1},\bm{X}_{2}]$.
 In other words, within the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the various combinations of possible latent strata including $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$ are exchangeable with respect to the expected contrast.
-Similarly, the conditional expected contrasts within $\mathcal{S}_{11}$, $\mathcal{S}_{00}$, and $\mathcal{S}_{01}$ are all equal to the estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$, respectively (see Table 1(a) for a concrete mapping between the observed data strata and the latent principal strata).
+Similarly, the conditional expected contrasts within $\mathcal{S}_{11}$, $\mathcal{S}_{00}$, and $\mathcal{S}_{01}$ are all equal to the estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,\bm{X}_{1},\bm{X}_{2}]$, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,\bm{X}_{1},\bm{X}_{2}]$, and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1,\bm{X}_{1},\bm{X}_{2}]$, respectively (see Table 1(a) for a mapping between the observed data strata and the latent principal strata).
 That is, with Assumption 3, each one of the observed subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only (Table 1(b)).
-Assumption 3 can be viewed as a generalization of the unitwise mean principal ignorability assumption in Jiang et al.
-(2022), since it reduces to that form in the linear contrast case h(u,v)=uvh(u,v)=u-v.
+Assumption 3 is a generalization of the unitwise mean principal ignorability assumption in Jiang2022, since it reduces to that form in the linear contrast case h(u,v)=uvh(u,v)=u-v.
 For nonlinear contrasts, however, expectations cannot be interchanged within the contrast function, and thus there is no clear nesting relationship between the pairwise and unitwise versions of mean principal ignorability.
 Consider the win estimands under noncompliance (Example 2) with the contrast function $h(u,v)=\mathbbm{1}(u>v)$.
-Applying (LABEL:eq:pi-equation) to the first two pairs in $\mathcal{S}_{10}$ yields $P\{Y_{1}(1)>Y_{2}(0)|S_{1}=10,S_{2}=10,X_{1}=x_{1},X_{2}=x_{2}\}=P\{Y_{1}(1)>Y_{2}(0)|S_{1}=11,S_{2}=10,X_{1}=x_{1},X_{2}=x_{2}\}$.
+Applying (LABEL:eq:pi-equation) to the first two pairs in $\mathcal{S}_{10}$ yields $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=10,S_{2}=10,\bm{X}_{1}=\bm{x}_{1},\bm{X}_{2}=\bm{x}_{2}\}=\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=11,S_{2}=10,\bm{X}_{1}=\bm{x}_{1},\bm{X}_{2}=\bm{x}_{2}\}$.
 This assumption states that, conditional on the same pairwise covariate profiles, the probability that a treated individual has a more favorable outcome than a control individual is the same whether the pair consists of two compliers or of one complier and one always-taker.
 In other words, there do not exist additional unmeasured factors influencing the conditional probability of winning among pairs in $\mathcal{S}_{10}$.
-This is neither stronger nor weaker than the unitwise mean principal ignorability assumption, which requires, for example, $\mathbb{E}\{Y_{1}(1)|S_{1}=10,X_{1}=x_{1}\}=\mathbb{E}\{Y_{1}(1)|S_{1}=11,X_{1}=x_{1}\}$.
+This is neither stronger nor weaker than the unitwise mean principal ignorability assumption, which requires, for example, $\mathbb{E}\{Y_{1}(1)|S_{1}=10,\bm{X}_{1}=\bm{x}_{1}\}=\mathbb{E}\{Y_{1}(1)|S_{1}=11,\bm{X}_{1}=\bm{x}_{1}\}$.
 Table 1: Correspondence between latent principal strata pairs, Assumption 3, and observed strata pairs defined by assignment and observed intermediate variable.
@@
 Latent strata pairs Conditions in Pairwise mean defined in observed strata
-$S_{1}$ $S_{2}$ Assumption 3 $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1},D_{1},Z_{2},D_{2},X_{1},X_{2}]$
+$S_{1}$ $S_{2}$ Assumption 3 $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1},D_{1},Z_{2},D_{2},\bm{X}_{1},\bm{X}_{2}]$
 10 10 (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
@@
 The first two are stated at the individual level, but since the observed data are assumed to be i.i.d., they extend naturally to pairs.
-For example, individual-level treatment ignorability implies that $(Z_{1},Z_{2})$ is jointly independent of the potential outcomes and intermediate variables given $(X_{1},X_{2})$.
-By contrast, Assumption 3 must be formulated at the pairwise level in its minimally sufficient form, because identification with nonlinear contrasts requires linking expectations across observed and latent strata pairs as in Table 1.
+For example, individual-level treatment ignorability implies that $(Z_{1},Z_{2})$ is jointly independent of the potential outcomes and intermediate variables given $(\bm{X}_{1},\bm{X}_{2})$.
+By contrast, Assumption 3 must be formulated at the pair level in its minimally sufficient form, because identification with nonlinear contrasts requires linking expectations across observed and latent strata pairs as in Table 1.
 Assumptions 2 and 3 are not testable from the observed data alone.
@@
 3 Nonparametric identification
-3.1 Identification of the PGCE estimands
-To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution.
-We consider both the observational and experimental studies and define the treatment propensity score by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin, 1983).
-In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu, 2016), for $s\in\{10,00,11\}$ under Assumption 2.
-Furthermore, we use $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$ to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively.
-Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\}$, for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$.
+3.1 Identification of the PGCEs
+To obtain nonparametric identification formulas for the PGCEs, we start by defining components of the observed data distribution.
+We consider both observational and experimental studies and define the treatment propensity score as $\pi(\bm{X})=\mathbb{P}(Z=1|\bm{X})$, the probability that a participant receives the treatment conditional on pre-treatment covariates (Rosenbaum1983).
+In addition, the principal score, $e_{s}(\bm{X})=\mathbb{P}(S=s|\bm{X})$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding2016), for $s\in\{10,00,11\}$ under Assumption 2.
+Furthermore, we define $\pi=\mathbb{E}\{\pi(\bm{X})\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(\bm{X})\}=\mathbb{P}(S=s)$.
+Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as $\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},\bm{X}_{1},\bm{X}_{2}\}$, for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$.
 Below, we provide two examples of the pairwise outcome mean models for specific contrast functions.
 Example 3 (Pairwise mean function for probabilistic index)
-For studying the probabilistic index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$.
-If the potential outcomes are continuous, a Gaussian regression model can be considered such that $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$.
+To study the probabilistic index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$.
+If the potential outcomes are continuous, a Gaussian regression model can be considered such that $Y(z)|D,\bm{X}\sim\mathcal{N}(f_{z}(D,\bm{X}),\sigma^{2})$.
 In this case, one can derive the pairwise outcome mean as
-$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\Phi\left\{\frac{f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2})}{\sqrt{2}\sigma}\right\}.$ (7)
+$\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2})=\Phi\left\{\frac{f_{z_{1}}(d_{1},\bm{X}_{1})-f_{z_{2}}(d_{2},\bm{X}_{2})}{\sqrt{2}\sigma}\right\}.$ (7)
 Example 4 (Pairwise outcome mean for win estimands)
 Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$.
-For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
-$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ (8)
-$\quad=\left\{\sum_{q=2}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2}),\sum_{q=1}^{Q-1}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2})\right\},$
-where $p_{zq}(D,X)$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $X$, and $Q(\geq 2)$ is the largest category of the ordinal outcome.
-For estimating the principal scores, we first define the conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$.
-Then, under Assumption 2, the principal scores can be expressed as $e_{10}(X)=p_{1}(X)-p_{0}(X)$, $e_{00}(X)=1-p_{1}(X)$, and $e_{11}(X)=p_{0}(X)$.
-The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$.
-We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$.
-Based on these elements, we establish three alternative nonparametric identification formulas for each PGCE estimand, as shown in Theorem S2 of the Online Supplement.
-Each formula requires specifying only two of the three models—the treatment propensity score, the principal score, and the pairwise outcome mean—rather than all three simultaneously.
-Furthermore, Theorem S1 of the Online Supplement presents the associated balancing conditions for any pairwise covariate functions.
-These balancing conditions provide the theoretical foundation for model checking of the treatment propensity and principal score models without peeking at the outcome data.
+For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $\bm{X}$, Mao2018 has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
+$\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2})$ (8)
+$\quad=\left\{\sum_{q=2}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},\bm{X}_{1})p_{z_{2}q^{\prime}}(d_{2},\bm{X}_{2}),\sum_{q=1}^{Q-1}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},\bm{X}_{1})p_{z_{2}q^{\prime}}(d_{2},\bm{X}_{2})\right\},$
+where $p_{zq}(D,\bm{X})$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $\bm{X}$, and $Q(\geq 2)$ is the largest category of the ordinal outcome.
+For estimating the principal scores, we first define the conditional probability, $p_{z}(\bm{X})=\mathbb{P}(D=1|Z=z,\bm{X})$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(\bm{X})\}=\mathbb{P}(D=1|Z=z)$.
+Then, under Assumption 2, the principal scores can be expressed as $e_{10}(\bm{X})=p_{1}(\bm{X})-p_{0}(\bm{X})$, $e_{00}(\bm{X})=1-p_{1}(\bm{X})$, and $e_{11}(\bm{X})=p_{0}(\bm{X})$.
+The population fractions of the three principal strata are $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$.
+We will refer to $\{p_{0}(\bm{X}),p_{1}(\bm{X})\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(\bm{X}),e_{00}(\bm{X}),e_{11}(\bm{X})\}$.
+Based on these elements, we establish three alternative nonparametric identification formulas for each PGCE estimand, given in Theorem S2 of the Online Supplement.
+Each formula requires specifying two of the three models—the treatment propensity score, the principal score, and the pairwise outcome mean—rather than all three simultaneously.
+Furthermore, Theorem S1 of the Online Supplement presents the associated covariate balancing conditions.
+These balancing conditions provide the foundation for model checking of the treatment propensity and principal score models without peeking at the outcome data.
 3.2 Estimators based on nonparametric identification formulas
-For point estimation, a common practice for implementing the nonparametric identification formulas is through specifying parametric regression models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$.
+For point estimation, a common approach is to specify parametric regression models for nuisance functions, i.e., $\pi(\bm{X};\alpha)$, $p_{z}(\bm{X};\beta)$ or $e_{s}(\bm{X};\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$.
 We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$.
-To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment propensity, the principal score, and the outcome mean models are correctly specified, respectively.
-That is, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$.
+To aid further discussions of the statistical properties of estimators for the PGCEs, we introduce notation $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote the conditions under which the treatment propensity, principal score, and outcome mean models are correctly specified, respectively.
+That is, under $\mathcal{M}_{\pi}$, we have $\pi(\bm{X};\alpha^{*})=\pi(\bm{X})$; under $\mathcal{M}_{p}$, we have $p_{z}(\bm{X};\beta^{*})=p_{z}(\bm{X})$ and $e_{s}(\bm{X};\beta^{*})=e_{s}(\bm{X})$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$.
 Additionally, we use the logic notation ‘+’ to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
@@
 For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$.
-Define the following quantity for any function $f(Y,D,X)$:
-$\begin{aligned} \psi_{f(Y_{z},D_{z},X)}(\mathcal{O})&=\frac{\mathbbm{1}(Z=z)[f(Y,D,X)-\mathbb{E}\{f(Y,D,X)|X,Z=z\}]}{\mathbb{P}(Z=z|X)}\\ &\qquad+\mathbb{E}\{f(Y,D,X)|X,Z=z\}\end{aligned}.$ (9)
-One example of (9) is $f(Y_{z},D_{z},X)=D_{z}$, in which case we obtain the following function:
-$\psi_{D_{z}}(\mathcal{O})=\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X)\}}{\mathbb{P}(Z=z|X)}+p_{z}(X).$ (10)
-(10) is useful for the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome:
+Define the following quantity for any function $f(Y,D,\bm{X})$:
+$\begin{aligned} \psi_{f(Y_{z},D_{z},\bm{X})}(\mathcal{O})&=\frac{\mathbbm{1}(Z=z)[f(Y,D,\bm{X})-\mathbb{E}\{f(Y,D,\bm{X})|\bm{X},Z=z\}]}{\mathbb{P}(Z=z|\bm{X})}\\ &\qquad+\mathbb{E}\{f(Y,D,\bm{X})|\bm{X},Z=z\}\end{aligned}.$ (9)
+One example of (9) is $f(Y_{z},D_{z},\bm{X})=D_{z}$, in which case we obtain the following function:
+$\psi_{D_{z}}(\mathcal{O})=\frac{\mathbbm{1}(Z=z)\{D-p_{z}(\bm{X})\}}{\mathbb{P}(Z=z|\bm{X})}+p_{z}(\bm{X}).$ (10)
+Equation (10) is useful for the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome:
 […]
 […]
-For any general constrast function $h$, the plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart, 1998, §12).
+For any general contrast function $h$, the plug-in estimators for the PGCEs take the form of U-statistics (vandervaart1998, §12).
 For example, the identification results in Theorem S2 of the Online Supplement motivate the following estimators for $\tau_{h}^{10}$:
-$\widehat{\tau}_{h,\pi+p}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right],$
-$\widehat{\tau}_{h,\pi+\mu}^{10}$ $=\mathbb{U}_{n}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1};\widehat{\alpha})-D_{1}(1-Z_{1})/\{1-\pi(X_{1};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\right$.
+$\widehat{\tau}_{h,\pi+p}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(\bm{X}_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(\bm{X}_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(\bm{X}_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(\bm{X}_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{0}(\bm{X}_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(\bm{X}_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right],$
+$\widehat{\tau}_{h,\pi+\mu}^{10}$ $=\mathbb{U}_{n}\left(\left[\frac{D_{1}Z_{1}/\pi(\bm{X}_{1};\widehat{\alpha})-D_{1}(1-Z_{1})/\{1-\pi(\bm{X}_{1};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\right$.
 […]
 […]
-Upon closer investigation, $\widehat{\tau}_{h,\pi+p}^{10}$ is based on the estimated treatment propensity score and the principal score, $\widehat{\tau}_{h,\pi+\mu}^{10}$ is based on the estimated treatment propensity score and the pairwise outcome mean, and $\widehat{\tau}_{h,p+\mu}^{10}$ is based on the estimated principal score and the pairwise outcome mean.
-For brevity, we will use $\tau_{h}^{10}$ as the running example for exposition in the remainder of the article.
-Results for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are analogous and are given in the middle of Section S2 of the Online Supplement.
-These estimators have strong requirements on model specification to consistently estimate the PGCEs.
+For brevity, we will use $\tau_{h}^{10}$ as a focused example in the remainder of the article.
+Results for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are analogous and are given in Section S2 of the Online Supplement.
+These estimators have strong requirements on model specification to estimate the PGCEs consistently.
 To reiterate, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment propensity and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
-Remark 2
-It is worth noting that our weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score for pairs, and can be viewed as an extension of the pairwise inverse propensity score weighted estimator in Mao (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable.
-The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately; See the Online Supplement for additional details.
-The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be obtained using empirical process theory for the parametric class (van der Vaart, 1998, §19.2).
-The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao (2017).
 4 Multiply robust and efficient U-statistics
 4.1 The efficient influence function for the PGCE estimand
-We first develop more efficient estimators with $\sqrt{n}$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al., 1993).
+We first develop more efficient estimators with $n^{1/2}$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickle1993).
 The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators.
 Several key properties of the EIF suggest its promise for constructing optimal estimators.
-First, the EIF often integrates multiple nuisance models and facilitates estimators that are robust against partial model misspecification.
-Second, the lower bound of the asymptotic variance is the variance of the EIF, providing a benchmark to define efficient estimators.
-Third, the EIF provides a device for incorporating data-adaptive machine learners for the nuisance parameters, suggesting feasible estimators that achieve the asymptotic variance lower bound with $\sqrt{n}$ convergence rate (Chernozhukov et al., 2018).
+First, the EIF integrates multiple nuisance models and facilitates estimators that are robust against partial model misspecification.
+Second, the lower bound on the asymptotic variance is the variance of the EIF, providing a benchmark for defining efficient estimators.
+Third, the EIF provides a device for incorporating data-adaptive machine learners for nuisance parameters, yielding optimal estimators that achieve the asymptotic variance lower bound with a $n^{1/2}$ convergence rate (Chernozhukov2018).
 We explore these properties by first formalizing the EIF for each PGCE.
-Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(X_{1})e_{10}(X_{2})\allowbreak\mu_{1100}(X_{1},X_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$.
+Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(\bm{X}_{1})e_{10}(\bm{X}_{2})\allowbreak\mu_{1100}(\bm{X}_{1},\bm{X}_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$.
 This suggests that the EIF for $\tau_{h}^{s}$, $\varphi_{h}^{s}(\mathcal{O})$, is of the form:
@@
 where $\varphi_{h,N}^{s}(\mathcal{O})$ and $\varphi_{D}^{s}(\mathcal{O})$ are EIFs for $\tau_{h,N}^{s}$ and $\tau_{D}^{s}$, respectively.
-Let $F_{zd}(\cdot|x)$ denote the conditional distribution of $Y_{i}(z)$ given $(Z_{i},D_{i},X_{i})=(z,d,x)$, and $F_{X}(\cdot)$ denote the marginal distribution of $X$.
-We can further define marginalized functions of the observed outcome, $\nu_{1,zd}^{s}(Y)=\int e_{s}(x)h(Y,y)\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$ and $\nu_{0,zd}^{s}(Y)=\int e_{s}(x)h(y,Y)\allowbreak\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$, for $z,d\in\{0,1\}$ and $s\in\{10,00,11\}$.
+Let $F_{zd}(\cdot|\bm{x})$ denote the conditional distribution of $Y_{i}(z)$ given $(Z_{i},D_{i},\bm{X}_{i})=(z,d,\bm{x})$, and $F_{\bm{X}}(\cdot)$ denote the marginal distribution of $\bm{X}$.
+We define marginalized functions of the observed outcome, $\nu_{1,zd}^{s}(Y)=\int e_{s}(\bm{x})h(Y,y)\mathrm{d}F_{zd}(y|\bm{x})\allowbreak\mathrm{d}F_{\bm{X}}(\bm{x})$ and $\nu_{0,zd}^{s}(Y)=\int e_{s}(\bm{x})h(y,Y)\allowbreak\mathrm{d}F_{zd}(y|\bm{x})\allowbreak\mathrm{d}F_{\bm{X}}(\bm{x})$, for $z,d\in\{0,1\}$ and $s\in\{10,00,11\}$.
 Then, the following Theorem 1 is obtained.
@@
 Under Assumptions 1 - 3, the EIF for $\tau_{h}^{10}$ is $\varphi_{h}^{10}(\mathcal{O})=\{\varphi_{h,N}^{10}(\mathcal{O})-\tau_{h}^{10}\varphi_{D}^{10}(\mathcal{O})\}/(p_{1}-p_{0})^{2}$, where $\varphi_{h,N}^{10}(\mathcal{O})=\phi_{h}^{10}(\mathcal{O})-2\tau_{h,N}^{10}$ and
-$\phi_{h}^{10}(\mathcal{O})=$ $~\frac{ZDe_{10}(X)}{\pi(X)p_{1}(X)}\left\{\nu_{1,00}^{10}(Y)-m_{1}^{10}(X)\right\}+\frac{(1-Z)(1-D)e_{10}(X)}{\{1-\pi(X)\}\{1-p_{0}(X)\}}\left\{\nu_{0,11}^{10}(Y)-m_{0}^{10}(X)\right\}$
-$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(X)+m_{0}^{10}(X)\right\},$
-with $m_{1}^{10}(X)=\int e_{10}(x)\mu_{1100}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{10}(X)=\int e_{10}(x)\mu_{1100}(x,X)\mathrm{d}F_{X}(x)$, and
+$\phi_{h}^{10}(\mathcal{O})=$ $~\frac{ZDe_{10}(\bm{X})}{\pi(\bm{X})p_{1}(\bm{X})}\left\{\nu_{1,00}^{10}(Y)-m_{1}^{10}(\bm{X})\right\}+\frac{(1-Z)(1-D)e_{10}(\bm{X})}{\{1-\pi(\bm{X})\}\{1-p_{0}(\bm{X})\}}\left\{\nu_{0,11}^{10}(Y)-m_{0}^{10}(\bm{X})\right\}$
+$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(\bm{X})+m_{0}^{10}(\bm{X})\right\},$
+with $m_{1}^{10}(\bm{X})=\int e_{10}(\bm{x})\mu_{1100}(\bm{X},\bm{x})\mathrm{d}F_{\bm{X}}(\bm{x})$, $m_{0}^{10}(\bm{X})=\int e_{10}(\bm{x})\mu_{1100}(\bm{x},\bm{X})\mathrm{d}F_{\bm{X}}(\bm{x})$, and
 $\varphi_{D}^{10}(\mathcal{O})=2\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}(p_{1}-p_{0})-2\tau_{D}^{10}.$
-The EIFs for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are given in Theorem S3 of Online Supplement Section S3.1.
-The EIFs for the PGCEs are derived following the chain rule via the method of Gâteaux derivatives as formalized in Ichimura and Newey (2022) and Hines et al.
-(2022) with respect to the distribution of observed data, $\mathcal{O}$.
+For brevity, the EIFs for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are given in Theorem S3 of Online Supplement Section S3.1.
+The EIFs for the PGCEs are derived following the chain rule via the method of Gâteaux derivatives with respect to the distribution of observed data, $\mathcal{O}$ (Ichimura2022; Hines2022).
 Based on Theorems 1 and S3, the asymptotic variances of estimators for the PGCEs are lower bounded by the respective semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$ for $s\in\{10,00,11\}$.
 4.2 Multiply robust estimators
-The EIFs imply the semiparametric efficiency lower bound and pave the way for formulating optimal estimators for the PGCEs.
+The EIFs imply the semiparametric efficiency lower bound and pave the way for the formulation of optimal estimators for the PGCEs.
 However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and require care for constructing the final estimator.
-We use $U$-statistics in constructing the final estimator, which could be shown to have the same influence function given in Theorem 1, thereby achieving local semiparametric efficiency with correctly specified nuisance functions.
+We use U-statistics to construct the final estimator, which has the same influence function as in Theorem 1, thereby achieving local semiparametric efficiency with correctly specified nuisance functions.
 To this end, we first notice that $\tau_{h}^{10}$ can be rewritten as
@@
 $g_{h}^{10}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$
-$=2^{-1}\bigg[\frac{Z_{i}D_{i}e_{10}(X_{i};\beta)}{\pi(X_{i};\alpha)p_{1}(X_{i};\beta)}\times\frac{(1-Z_{j})(1-D_{j})e_{10}(X_{j};\beta)}{\{1-\pi(X_{j};\alpha)\}\{1-p_{0}(X_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1100}(X_{i},X_{j};\gamma)\right\}$
-$\qquad\qquad+\frac{Z_{j}D_{j}e_{10}(X_{j};\beta)}{\pi(X_{j};\alpha)p_{1}(X_{j};\beta)}\times\frac{(1-Z_{i})(1-D_{i})e_{10}(X_{i};\beta)}{\{1-\pi(X_{i};\alpha)\}\{1-p_{0}(X_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1100}(X_{j},X_{i};\gamma)\right\}$
-$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(X_{i},X_{j};\gamma)$
-$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(X_{j},X_{i};\gamma)\bigg],$
+$=2^{-1}\bigg[\frac{Z_{i}D_{i}e_{10}(\bm{X}_{i};\beta)}{\pi(\bm{X}_{i};\alpha)p_{1}(\bm{X}_{i};\beta)}\times\frac{(1-Z_{j})(1-D_{j})e_{10}(\bm{X}_{j};\beta)}{\{1-\pi(\bm{X}_{j};\alpha)\}\{1-p_{0}(\bm{X}_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1100}(\bm{X}_{i},\bm{X}_{j};\gamma)\right\}$
+$\qquad\qquad+\frac{Z_{j}D_{j}e_{10}(\bm{X}_{j};\beta)}{\pi(\bm{X}_{j};\alpha)p_{1}(\bm{X}_{j};\beta)}\times\frac{(1-Z_{i})(1-D_{i})e_{10}(\bm{X}_{i};\beta)}{\{1-\pi(\bm{X}_{i};\alpha)\}\{1-p_{0}(\bm{X}_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1100}(\bm{X}_{j},\bm{X}_{i};\gamma)\right\}$
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(\bm{X}_{i},\bm{X}_{j};\gamma)$
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(\bm{X}_{j},\bm{X}_{i};\gamma)\bigg],$
 […]
 $\widehat{\tau}_{h,\mathrm{tr}}^{10}=\frac{\mathbb{U}_{n}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},$ (13)
-which is essentially a $U$-statistic but in the form of ratios.
+which is essentially a scaled U-statistic.
 Estimators $\widehat{\tau}_{h,\mathrm{tr}}^{00}$ and $\widehat{\tau}_{h,\mathrm{tr}}^{11}$ are similarly constructed and defined in Section S4.1 of the Online Supplement.
@@
 Theorem 2 (Multiply robustness and efficiency)
-Under Assumptions 1 - 3, and suppose there exists some $\varepsilon\in(0,1)$ such that $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),\allowbreak p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$.
+Under Assumptions 1 - 3, and suppose there exists some $\varepsilon\in(0,1)$ such that $\varepsilon<\{\pi(\bm{X};\alpha^{*}),\pi(\bm{X};\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(\bm{X};\beta^{*}),\allowbreak p_{1}(\bm{X};\widehat{\beta}),1-p_{0}(\bm{X};\beta^{*}),1-p_{0}(\bm{X};\widehat{\beta})\}>\varepsilon$ for all $\bm{X}\in\bm{\mathcal{X}}$.
 Then, each estimator $\widehat{\tau}_{h,\mathrm{tr}}^{s}$, $s\in\{10,00,11\}$, is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
-Under Theorem 2, $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified.
+By Theorem 2, $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified.
 This triple robustness property of $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is obtained by separately investigating the properties of its numerator and the denominator.
-Specifically, the denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$.
-To account for the uncertainty in estimating the parametric nuisance functions, we estimate the variances of the triply robust estimators via a nonparametric bootstrap.
+To account for the uncertainty in estimating the parametric nuisance functions, we estimate the variances of the triply robust estimators via the nonparametric bootstrap.
 4.3 Debiased machine learning estimators
-We propose further improvement of the parametric multiply robust estimator through debiased machine learning, which is similarly based on orthogonal moment functions (Chernozhukov et al., 2018) or EIFs, but allows efficient estimation of causal effects with data-adaptive nuisance estimators.
-Because the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al., 2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand.
-In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates.
-Compared to parametric working models, leveraging machine learning-based working models for the nuisance functions is more advantageous in terms of robustness against potential misspecification.
+We study further improvements to the parametric multiply robust estimator via debiased machine learning, which is similarly based on orthogonal moment functions (Chernozhukov2018) or EIFs, but allows efficient estimation of causal effects using data-adaptive nuisance estimators.
+Because the Neyman orthogonality holds for the EIF (Chernozhukov2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand.
+In addition, the mixed bias property of the EIF representation allows one to obtain $n^{1/2}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates.
+Compared with parametric working models, machine learning based working models for the nuisance functions are more robust to potential misspecification.
 To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting—is adopted.
-In this procedure, the entire sample is divided into separate parts for estimating the nuisance functions and the causal estimand, respectively (Chernozhukov et al., 2018).
-Mild regulatory conditions, mainly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy, 2023).
-We adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur (2022).
-Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition $\mathcal{I}=\{I_{1},\ldots,I_{K}\}$.
+In this procedure, the entire sample is divided into separate parts for estimating the nuisance functions and the causal estimand, respectively (Chernozhukov2018).
+Mild regulatory conditions, mainly the consistency of machine learning first-step estimates, are required for the $n^{1/2}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (kennedy2023semiparametric).
+We adopt the sample-splitting scheme for U-statistics provided in escanciano2022machine.
+Specifically, the sample index set $\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition $\mathcal{I}=\{I_{1},\ldots,I_{K}\}$.
 Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $\{(i,j),1\leq i<j\leq n\}$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$.
-Chernozhukov et al.
-(2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$.
+Chernozhukov2018 suggests setting $K$ to moderate values, such as 4 or 5, so that the sizes of elements in $\mathcal{I}$ are $O(n)$.
 We use the following example to demonstrate the definition of $\mathcal{I}$ and $\mathcal{P}$.
@@
 For a given partition $\mathcal{P}$ (based on the initial partition of singletons $\mathcal{I}$), we obtain intermediate estimates of the target estimand using observations in each $P_{l}$ (or $I_{k}$) with $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$) estimated using observations in $P_{l}^{c}$ (or $I_{k}^{c}$), those are not present in $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$).
-By pooling these intermediate estimates together, we can obtain the following debiased machine learning estimator for $\tau_{h}^{10}$,
+By pooling these intermediate estimates together, we obtain the debiased machine learning estimator for $\tau_{h}^{10}$ as
 $\widehat{\tau}_{h,\mathrm{dml}}^{10}=\frac{\mathbb{U}_{n,L}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n,K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},$ (14)
@@
 The debiased machine learning estimators $\widehat{\tau}_{h,\mathrm{dml}}^{00}$ and $\widehat{\tau}_{h,\mathrm{dml}}^{11}$ can be constructed in a similar fashion and are given in Section S5 of the Online Supplement.
-We now discuss the asymptotic properties of $\widehat{\tau}_{h,\mathrm{dml}}^{s}$.
-We first assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability.
+To establish the asymptotic properties of $\widehat{\tau}_{h,\mathrm{dml}}^{s}$, we first assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability.
 In addition, we assume true nuisance realizations $\theta^{*},\vartheta^{*}\in\Theta^{*}$.
@@
   • (d)
-    $g_{h,1}^{s}(\mathcal{O};\theta)\equiv\mathbb{E}g_{h}^{s}(\mathcal{O},o;\theta)$ and $\psi_{D_{z}}(\mathcal{O};\vartheta)$ are $\mathbb{P}$-Donsker with probability approaching 1, and there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
-$\widehat{\tau}_{h,\mathrm{dml}}^{s}$ is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the variance of the EIF.
-Theorem 3 establishes the consistency and asymptotic normality of the debiased machine learning estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al.
-(2022, Lemma 15) and Escanciano and Terschuur (2022, Lemma 4).
+    $g_{h,1}^{s}(\mathcal{O}_{1};\theta)\equiv\mathbb{E}g_{h}^{s}(\mathcal{O}_{1},\mathcal{O}_{2};\theta)$ and $\psi_{D_{z}}(\mathcal{O};\vartheta)$ are $\mathbb{P}$-Donsker with probability approaching 1, and there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
+then $\widehat{\tau}_{h,\mathrm{dml}}^{s}$ is $n^{1/2}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the variance of the EIF.
+Theorem 3 establishes the consistency and asymptotic normality of the debiased machine learning estimators, building on the results developed for orthogonal moment functions given in Chernozhukov2022 and escanciano2022machine.
 Regulatory conditions in Theorem 3 are mostly mild.
 Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$.
-Condition (b) assumes that machine learning first-step estimators are consistent, as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners.
-Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al., 2018).
-The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator; similar conditions have been previously invoked in the debiased machine learning literature (Vansteelandt and Dukes, 2022).
-Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (14) via nonparametric bootstrap.
+Condition (b) assumes that machine learning first-step estimators are consistent, as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for existing machine learners.
+Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov2018).
+The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator; similar conditions have been previously invoked in the debiased machine learning literature (Vansteelandt2022; Jiang2022).
+Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (14) via the nonparametric bootstrap.
 4.4 Sensitivity analysis for monotonicity
-To point identify PGCEs, a central assumption is monotonicity (Assumption 2) which rules out the $01$ stratum.
+To point identify PGCEs, a central assumption is monotonicity (Assumption 2), which rules out the $01$ stratum.
 This stratum may be relevant, for example, when comparing two active interventions with unknown survival benefits against non-mortality outcomes.
 We thus describe a sensitivity analysis strategy to evaluate the robustness of the proposed estimators in the event of departure from monotonicity.
-Following the approach in Ding and Lu (2016), we define the sensitivity function $\eta(X)$ as the ratio between the proportion of strata $01$ and $10$ conditional on the covariates, i.e., $\eta(X)=\mathbb{P}(S=01|X)/\mathbb{P}(S=10|X)$, where $\eta(X)\in[0,\infty)$; the monotonicity holds if $\eta(X)=0$, but $\eta(X)>0$ implies the existence of the $01$ stratum.
+Following the approach in Ding2016, we define the sensitivity function $\eta(\bm{X})$ as the ratio between the proportion of strata $01$ and $10$ conditional on the covariates, i.e., $\eta(\bm{X})=\mathbb{P}(S=01|\bm{X})/\mathbb{P}(S=10|\bm{X})$, where $\eta(\bm{X})\in[0,\infty)$; the monotonicity holds if $\eta(\bm{X})=0$, but $\eta(\bm{X})>0$ implies the existence of the $01$ stratum.
 The PCGEs in this scenario are thus $\tau_{h,\eta}^{s}$ for $s\in\{10,01,00,11\}$.
-To proceed, we first notice that, for $\eta(X)\neq 1$, the principal scores can be obtained as:
-$e_{10,\eta}(X)=\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},~~e_{01,\eta}(X)=\frac{\eta(X)\{p_{1}(X)-p_{0}(X)\}}{1-\eta(X)},$ (15)
-$e_{00,\eta}(X)=1-p_{0}(X)-\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},~~e_{11,\eta}(X)=p_{1}(X)-\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},$
-with $0\leq\eta(X)\leq 1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}$, and the proportions of principal strata by $e_{s,\eta}=\mathbb{E}\{e_{s,\eta}(X)\}$ for all $s$.
-When $\eta(X)=1$, strata $10$ and $01$ have equal proportions, which corresponds to the situation where the treatment has zero average causal effect on $S$ and the principal scores are not identifiable (Jiang et al., 2022, Section S4.3).
+To proceed, we first notice that, for $\eta(\bm{X})\neq 1$, the principal scores can be obtained as:
+$e_{10,\eta}(\bm{X})=\frac{p_{1}(\bm{X})-p_{0}(\bm{X})}{1-\eta(\bm{X})},~~e_{01,\eta}(\bm{X})=\frac{\eta(\bm{X})\{p_{1}(\bm{X})-p_{0}(\bm{X})\}}{1-\eta(\bm{X})},$ (15)
+$e_{00,\eta}(\bm{X})=1-p_{0}(\bm{X})-\frac{p_{1}(\bm{X})-p_{0}(\bm{X})}{1-\eta(\bm{X})},~~e_{11,\eta}(\bm{X})=p_{1}(\bm{X})-\frac{p_{1}(\bm{X})-p_{0}(\bm{X})}{1-\eta(\bm{X})},$
+with $0\leq\eta(\bm{X})\leq 1-\{p_{1}(\bm{X})-p_{0}(\bm{X})\}/\min\{p_{1}(\bm{X}),1-p_{0}(\bm{X})\}$, and the proportions of principal strata by $e_{s,\eta}=\mathbb{E}\{e_{s,\eta}(\bm{X})\}$ for all $s$.
+When $\eta(\bm{X})=1$, strata $10$ and $01$ have equal proportions, which corresponds to the situation where the treatment has zero average causal effect on $S$, and the principal scores are not identifiable (Jiang2022, Section S4.3).
 Thus, this boundary case was ruled out.
 The identification of the PGCEs without monotonicity also requires an alternative version of the principal ignorability (Assumption S1), which we provide in Section S7.1 of the Online Supplement.
-With a fixed value of the sensitivity function $\eta(X)$, we derive bias-corrected multiply robust and debiased machine learning estimators for the PGCEs, leveraging the EIFs obtained without assuming monotonicity.
+With a fixed value of the sensitivity function $\eta(\bm{X})$, we derive bias-corrected multiply robust and debiased machine learning estimators for the PGCEs, leveraging the EIFs obtained without assuming monotonicity.
 The EIFs and the explicit forms of the bias-corrected estimators are given in Section S7.1 of the Online Supplement.
-Importantly, these multiply robust estimators remain triply robust under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$, as summarized in the following result (proofs provided in Section S7.2 of the Online Supplement).
-Theorem 4 (Multiply robustness and efficiency of bias-corrected estimators)
-Suppose that Assumptions 1 and S1 in Section S7.1 of the Online Supplement hold with a known $\eta(X)\neq 1$, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$.
-Then, for $s\in\{10,01,00,11\}$, the estimator $\widehat{\tau}_{h,\eta,\mathrm{tr}}^{s}$ is (a) triply robust such that it is consistent for $\tau_{h,\eta}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; and (b) asymptotically linear with influence function $\varphi_{h,\eta}^{s}$, thereby achieving the local semiparametric efficiency bound under $\mathcal{M}_{\pi+p+\mu}$.
+Importantly, these multiply robust estimators remain triply robust under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$, as summarized in the Theorem S6 in Section S7.1 of the Online Supplement.
 5 Simulation studies
-We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for the PGCEs, including the three plug-in moment estimators motivated by the nonparametric identification results in Section S2 of the Online Supplement, the triply robust estimators with parametric working models given in (13), and the debiased machine learning estimators given in (14), under different data-generating processes.
 5.1 Probabilistic index with an intermediate variable
-In the first simulation study, we consider the case of the probabilistic index estimand introduced in Examples 1 and 3, where the potential outcomes are continuous and drawn from a normal distribution.
-We first generate four pre-treatment covariates $X_{i}=(X_{i1},X_{i2},X_{i3},X_{i4})^{\prime}$, where $X_{i1},X_{i2},X_{i3}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$, and $X_{i4}\overset{\text{i.i.d.}}{\sim}\mathcal{B}(0.5)$.
-Then, we draw the treatment assignment from $Z_{i}|X_{i}\sim\mathcal{B}\{\mathrm{expit}(-X_{i1}+0.5X_{i2}-0.25X_{i3}-0.1X_{i4})\}$.
-Finally, the potential intermediate variable and the potential outcome are drawn from $D_{i}(Z_{i})|Z_{i},X_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\}$, and $Y_{i}(Z_{i})|D_{i},Z_{i},X_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1)$ for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$.
+In the first simulation study, we focus on the probabilistic index estimand introduced in Examples 1 and 3, in which the potential outcomes are continuous and drawn from a normal distribution.
+We first generate four pre-treatment covariates $\bm{X}_{i}=(X_{i1},X_{i2},X_{i3},X_{i4})^{\top}$, where $X_{i1},X_{i2},X_{i3}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$, and $X_{i4}\overset{\text{i.i.d.}}{\sim}\mathcal{B}(0.5)$.
+Then, we draw the treatment assignment from $Z_{i}|\bm{X}_{i}\sim\mathcal{B}\{\mathrm{expit}(-X_{i1}+0.5X_{i2}-0.25X_{i3}-0.1X_{i4})\}$.
+Finally, the potential intermediate variable and the potential outcome are drawn from $D_{i}(Z_{i})|Z_{i},\bm{X}_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\}$, and $Y_{i}(Z_{i})|D_{i},Z_{i},\bm{X}_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1)$ for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$.
 We consider a binary intermediate variable ($D$ can be considered as the treatment receipt in the context of non-compliance) and assume the interest lies in estimating all three principal probabilistic index estimands ($\tau_{h}^{10}$, $\tau_{h}^{00}$ and $\tau_{h}^{11}$ with $h(u,v)=\mathbbm{1}(u\geq v)$).
-In addition, we create transformed covariates, $\widetilde{X}_{i}=(\widetilde{X}_{i1},\widetilde{X}_{i2},\widetilde{X}_{i3},\widetilde{X}_{i4})^{\prime}$, where
+In addition, we create transformed covariates, $\widetilde{\bm{X}}_{i}=(\widetilde{X}_{i1},\widetilde{X}_{i2},\widetilde{X}_{i3},\widetilde{X}_{i4})^{\top}$, where
 $\widetilde{X}_{i1}=\exp(0.5X_{i1}),~\widetilde{X}_{i2}=\frac{X_{i2}}{1+X_{i1}},~\widetilde{X}_{i3}=\left(\frac{X_{i2}X_{i3}}{25}+0.6\right)^{3},~\widetilde{X}_{i4}=\frac{X_{i3}X_{i4}}{\sqrt{2}},$
-to simulate scenarios where the nuisance functions are misspecified.
-Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise.
-It is worth noting that, although in theory, machine learning methods are flexible with richer model spaces as $n$ increases, model misspecification can still be relevant to the debiased machine learning estimators, as their estimation accuracy in finite samples can be affected by the quality of the input covariate feature matrix (Cheng and Li, 2025).
+are used to simulate scenarios where the nuisance functions are misspecified.
+We fit the model using $\bm{X}_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{\bm{X}}_{i}$ otherwise.
+It is worth noting that, although in theory, machine learning methods are flexible with richer model spaces as $n$ increases, model misspecification can still be relevant to the debiased machine learning estimators, as their estimation accuracy in finite samples can be affected by the quality of the input covariate matrix (cheng2025inverting).
 For the plug-in moment estimators and the triply robust estimator, treatment propensity scores and principal scores are estimated using logistic regression, whereas the pairwise outcome mean is estimated using linear regression, as in Example 3.
@@
 In contrast, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy.
-Finally, compared to the triply robust estimator, the debiased machine learning estimator yields a slight improvement, as it is generally less biased.
+Finally, compared with the triply robust estimator, the debiased machine learning estimator yields a slight improvement because it is generally less biased.
 (a) $\tau_{h}^{10}$
@@
 We further consider the case of estimating the win ratio estimands introduced in Examples 2 and 4, with the final potential outcomes being ordinal.
-We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study.
+We generate covariates $\bm{X}_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study.
 However, for the potential outcome $Y_{i}(z)\in\{1,2,3\}$, we consider the following three-category ordinal logistic regression model:
@@
 We set the number of observations $n=2,000$ with $1,000$ replications.
-In parallel to Section 5.1, violin plots summarizing the simulation results of each considered estimator are given in Figure 2 and Web Figure S2 in Section S6 of the Online Supplement.
-From Figure 2, we observe that the plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified.
+In parallel to Section 5.1, violin plots summarizing the simulation results of each considered estimator are given in Web Figures S2 and S3 in Section S6 of the Online Supplement.
+From Web Figure S2, we observe that the plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified.
 In contrast, the triply robust and debiased machine learning estimators are consistent in all four scenarios where at most one of the three nuisance functions is misspecified.
-When at least two of the three nuisance functions are misspecified (as shown in Web Figure S2 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the others in terms of estimation accuracy (bias).
-(a) $\tau_{h}^{10}$
-(b) $\tau_{h}^{00}$
-(c) $\tau_{h}^{11}$
-Figure 2: Violin plots of estimators for the win ratio estimands in scenarios where at most one nuisance function is misspecified.
-Labels: ‘tp’ for treatment propensity or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment propensity (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (13); and ‘DML’ for the debiased machine learning estimators given in (14).
-Horizontal dashed lines represent the true values of the PGCEs.
+When at least two of the three nuisance functions are misspecified (Web Figure S3 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the others in terms of estimation accuracy (bias).
 5.3 Additional simulations under non-monotonicity
-We consider a third simulation study to demonstrate the potential bias in the uncorrected triply robust and debiased machine learning estimators, as well as to validate the bias-corrected estimators under non-monotonicity.
-To achieve these goals, we modify the data-generating processes in Sections 5.1 and 5.2 by integrating the sensitivity function $\eta(X)$.
-Specifically, continuing with $p_{z}(X)=\mathrm{expit}(-1+2z+X_{1}-0.8X_{2}+0.6X_{3}-X_{4})$ for $z\in\{0,1\}$, we adopt the following configuration for the sensitivity function: $\eta(X)=\eta_{0}[1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}]$, where, we let $\eta_{0}\in[0,1]$ to ensure the true sensitivity function is well-defined within the permissible bound.
-We consider two scenarios: one with moderate violations, represented by $\eta_{0}=0.5$, and another with severe violations, represented by $\eta_{0}=0.99$.
-The strata membership $S$ is drawn from from a multinominal distribution $\mathcal{M}\mathcal{N}\{e_{10,\eta}(X),e_{01,\eta}(X),e_{00,\eta}(X),e_{11,\eta}(X)\}$, where $\{e_{s,\eta}(X),s\in\{10,01,00,11\}\}$ are obtained following (15).
-The treatment assignment $Z\sim\mathcal{B}\{\mathrm{expit}(-X_{1}+0.5X_{2}-0.25X_{3}-0.1X_{4})\}$, and the intermediate variable $D$ is determined once $Z$ and $S$ are known.
-That is, $D=1$ if $Z=1,S\in\{10,11\}$ or $Z=0,S\in\{01,11\}$, and $D=0$ if $Z=1,S\in\{01,00\}$ or $Z=0,S\in\{10,00\}$.
-The potential outcome $Y(z)$ for $z\in\{0,1\}$ is generated following (16), based on the same covariates simulated as in Section 5.1.
-To focus on the main ideas in this sensitivity simulation, we simplify without further considering nuisance function misspecification.
-We compare four estimators: the triply robust and debiased machine learning estimators under monotonicity, and the bias-corrected estimators with a known and correctly specified $\eta(X)$.
-The results are presented in Web Figure S3 in Section S6 of the Online Supplement.
-Overall, the simulation results show that the bias-corrected estimators give consistent estimates of all four PGCEs under both non-monotonicity scenarios.
-As $\eta_{0}$ increases from $0.5$ to $0.99$, the biases of the triply robust and debiased machine learning estimators increase accordingly, due to a higher degree of non-monotonicity.
-Interestingly, such increases in bias are not uniform across all PGCEs.
-For example, the biases of $\widehat{\tau}_{h,\mathrm{tr}}^{10}$ and $\widehat{\tau}_{h,\mathrm{dml}}^{10}$ are relatively small under both $\eta_{0}$ values, suggesting that they may be dramatically affected by violations of the monotonicity assumption in our setup.
-On the other hand, the biases of $\widehat{\tau}_{h,\mathrm{tr}}^{00}$, $\widehat{\tau}_{h,\mathrm{dml}}^{00}$, $\widehat{\tau}_{h,\mathrm{tr}}^{11}$, and $\widehat{\tau}_{h,\mathrm{dml}}^{11}$ increased significantly, indicating their relatively higher sensitivity to non-monotonicity.
-In further exploration, we find that as $\eta_{0}$ increases, a larger proportion of the $01$ stratum comes from the $00$ and $11$ strata, hence contributing to increased bias in the estimators for $\tau_{h}^{00}$ and $\tau_{h}^{11}$.
+We conduct a third simulation study to demonstrate bias in the uncorrected triply robust and debiased machine learning estimators, and to validate the bias-corrected estimators under non-monotonicity.
+We modify the data-generating processes in Sections 5.1 and 5.2 by integrating the sensitivity function $\eta(\bm{X})$.
+The detailed data-generating models and numerical results are presented in Section S7.2 of the Online Supplement.
+Overall, the results confirm that the bias-corrected estimators give consistent estimates of all four PGCEs across scenarios, suggesting their usefulness when monotonicity cannot be assumed.
 6 Data example
 We illustrate the proposed methods with a data example from the U.S.
-Jobs Corps study (Schochet et al., 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24.
+Jobs Corps study (Schochet2001), in which individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged 16 to 24.
 The data were from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, which received an immediate offer to participate in Job Corps, or the control group, which did not participate in Job Corps until three years later.
@@
 The proposed methods were implemented to examine the causal effect of the Job Corps intervention, with weekly earnings one year after the assignment selected as the final outcome.
-As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects.
+As earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrast may provide complementary insights into the treatment effects.
 In each data analysis below, we adjust for baseline covariates, including demographics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
@@
 We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$.
-In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al.
-(2009) for a similar setup.
+In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); see also Zhang2008 for a similar setup.
 Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps.
-For methods based on parametric working models, a logistic model is considered for the employment process within each assignment group, and a log-normal model is fitted for the final earning outcome to induce the pairwise mean outcome function.
+For methods based on parametric working models, a logistic model is fitted for the employment process within each assignment group, and a log-normal model is fitted for the final earnings outcome to estimate the pairwise mean outcome function.
 For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear models libraries.
 The estimated results for SPI are presented in Table 2, where $1,000$ bootstrap replicates were generated to estimate the standard errors.
-We observe that, for the always-employed (72.9% of all participants by estimation), the probabilistic index is estimated to be $0.539$ using the triply robust and debiased machine learning methods.
-This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment.
+We observe that, for the always-employed (72.9% of all participants, by estimation), the probabilistic index is estimated at $0.539$ using the triply robust and debiased machine learning methods.
+This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this shows that the Jobs Corps positively affects one’s earnings in the one year after the assignment.
 Finally, the point estimates obtained under the plug-in moment methods are slightly closer to the null compared to those from the triply robust and debiased machine learning methods, although their 95% confidence intervals continue to exclude the null.
@@
 We provide additional analysis of the Job Corps data to illustrate the application of the proposed methods to address ordinal outcomes with noncompliance.
-We use the actual treatment status as an intermediate variable, similar to Chen and Flores (2015).
+We use the actual treatment status as an intermediate variable, similar to Chen2015.
 Under the monotonicity assumption, three strata potentially exist, including the compliers (stratum $10$), the never-takers (stratum $00$), and the always-takers (stratum $11$).
-To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu et al.
-(2018).
+To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu2018.
 Under this setting, we estimate the principal win ratios within each principal stratum, as defined in (4).
 For methods based on parametric working models, the pairwise outcome mean is estimated using ordered logit regression, while treatment propensity and principal scores are estimated using logistic regression.
-For the debiased machine learning approach, we used multinomial generalized linear models with the elastic net penalty to estimate the pairwise outcome mean.
-The treatment propensity and principal scores were estimated using the SuperLearner, fitted under a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
+For the debiased machine learning approach, we used multinomial generalized linear models with an elastic net penalty to estimate the pairwise outcome mean.
+The treatment propensity and principal scores were estimated using the SuperLearner, fitted via a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
 The estimated results are presented in Table 3, where $1,000$ bootstrap replicates were generated to estimate the standard errors.
@@
 The Jobs Corps intervention assignment has a positive effect on participants’ earnings one year after the assignment for both compliers and always-takers, and this effect appears to be more pronounced for compliers.
-For example, using the debiased machine learning estimator, there is a 27% increase in the chances of being in a higher earning category for a complier under treatment compared to a complier under control.
+For example, using the debiased machine learning estimator, there is a 27% increase in the probability of being in a higher-earning category among compliers under treatment relative to those under control.
 The 95% lower confidence limits obtained from all five methods are strictly above the null value of 1.
@@
 Interestingly, in this reanalysis using a win ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable.
-The statistically significant effect among the always-takers is indicative of a positive psychological effect.
-That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence more likely to move into a higher earning category) compared to those who switched from their initial control assignment.
+The statistically significant effect among the always-takers indicates a positive psychological effect due to the assignment alone.
+That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence more likely to move into a higher-earning category) than those who switched from their initial control assignment.
 Table 3: Estimated win ratios within different compliance strata for the Jobs Corps data.
@@
 We conducted additional sensitivity analyses to examine the impact of potential violations of the monotonicity assumption, focusing on both Sections 6.1 and 6.2.
-Following the approach in Section 5.3, we specified the sensitivity function as $\eta(X)=\eta_{0}[1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}]$, with $\eta_{0}\in\{0.1,0.2,0.3,0.4,0.5\}$.
+Following the approach in Section 5.3, we specified the sensitivity function as $\eta(\bm{X})=\eta_{0}[1-\{p_{1}(\bm{X})-p_{0}(\bm{X})\}/\min\{p_{1}(\bm{X}),1-p_{0}(\bm{X})\}]$, with $\eta_{0}\in\{0.1,0.2,0.3,0.4,0.5\}$.
 This formulation allows the sensitivity function to naturally vary with covariates while being governed by a one-dimensional strength parameter $\eta_{0}$.
-The chosen range of $\eta_{0}$ represents mild to moderate departures from monotonicity, which is plausible in the context of the Jobs Corps study, where encouragement for employment and subsequent treatment receipt may not dramatically deviate from monotonicity.
-We applied both the triply robust and the debiased machine learning estimators, with results reported in Figure 3.
-For the SPI analysis (Figure 3(a)), increasing $\eta_{0}$—corresponding to stronger departures from monotonicity—led to slightly larger estimated PGCEs.
-In other words, as the proportion of harmed strata increases, the estimated SPI shifts further from the null, yielding a stronger treatment effect signal, though statistical significance remains unaffected.
-In contrast, for the PWR analyses (Figures 3(b) - (d)), the estimated PGCEs are largely stable across values of $\eta_{0}$.
+The chosen range of $\eta_{0}$ represents mild to moderate departures from monotonicity, which is plausible in the context of the Jobs Corps study, where encouragement for employment and subsequent treatment receipt may not deviate dramatically from monotonicity.
+We applied both the triply robust and the debiased machine learning estimators, with results reported in Web Figure S5 in Section S7.2 of the Online Supplement.
+For the SPI analysis (Web Figure S5(a)), increasing $\eta_{0}$, corresponding to stronger departures from monotonicity, led to slightly larger estimated PGCEs.
+In other words, as the proportion of harmed strata increases, the estimated SPI shifts further from the null, yielding a stronger treatment effect signal, although statistical significance remains unchanged.
+In contrast, for the PWR analyses (Web Figures S5(b) - (d)), the estimated PGCEs are largely stable across values of $\eta_{0}$.
 Despite the presence of defiers, the estimated win ratios across all three principal strata remain nearly unchanged.
 The only noticeable pattern is a narrowing of the 95% confidence interval for the complier win ratio as $\eta_{0}$ increases.
-Collectively, these findings suggest that the PWR estimators exhibit strong robustness to non-monotonicity, in contrast to the slightly more pronounced sensitivity observed for the SPI.
-(a) SPI
-(b) $\mathrm{PWR}^{10}$
-(c) $\mathrm{PWR}^{00}$
-(d) $\mathrm{PWR}^{11}$
-Figure 3: Estimated SPI and PWRs and 95% confidence intervals under different values of sensitivity analysis parameter $\eta_{0}$.
-The horizontal axis of each plot shows values of $\eta_{0}$, and the vertical axis shows values of the estimated estimands and confidence intervals.
-SPI: the survivor probabilistic index in Section 6.1; $\mathrm{PWR}^{s}$: the principal win ratio of stratum $s$ in Section 6.2, with $s\in\{10,00,11\}$.
+These findings suggest that the PWR estimators exhibit strong robustness to non-monotonicity, in contrast to the slightly more pronounced sensitivity observed for the SPI.
 7 Discussion
-In this article, we develop principal stratification methods for addressing a general class of PGCE estimands defined in terms of a binary intermediate variable.
-Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability that are based on the contrast function, we contributed a semiparametric efficient approach for estimating the PGCE estimands by identifying their efficient influence functions and obtaining the ratio $U$-statistic estimators that combine several different nuisance working functions.
-We showed that the proposed $U$-statistic estimators are triply robust and become locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric models or machine learning models with cross-fitting.
-As in all principal stratification analyses that achieve point identification, the proposed $U$-statistic estimators for estimating PGCEs also depend on the mean principal ignorability assumptions that are unverifiable from the observed data alone.
-In Section S7 of the Online Supplement, we have described an additional sensitivity function approach to obtain the bias-corrected $U$-statistic estimators for estimating the PGCEs when the mean principal ignorability assumption is violated.
-The bias-corrected estimators remain triply robust to potential working model misspecifications and provide means to generate a range of point estimates under assumed non-ignorability.
+We develop principal stratification methods for addressing a general class of PGCE estimands defined in terms of a binary intermediate variable.
+Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability based on the contrast function, we proposed a semiparametric approach for estimating the PGCEs by identifying their efficient influence functions and obtaining scaled U-statistic-based estimators.
+We showed that the proposed U-statistic-based estimators are triply robust and locally efficient, with asymptotic variances that attain the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric or machine learning models with cross-fitting.
+As in all principal stratification analyses that achieve point identification, the proposed U-statistic-based estimators of PGCEs also depend on the mean principal ignorability assumptions, which are unverifiable from the observed data alone.
+In Section S7 of the Online Supplement, we describe an additional sensitivity-function approach to obtain bias-corrected U-statistic-based estimators of the PGCEs when the mean principal ignorability assumption is violated.
+The bias-corrected estimators remain triply robust to potential working model misspecifications and provide a means to generate a range of point estimates under the assumption of non-ignorability.
 Finally, our current development assumes that the observed data vectors are independent and identically distributed, providing a natural starting point.
 Extending these ideas to correlated data structures is an important direction for future research.
-For clustered data, recent work in cluster randomized trials has emphasized the need to define estimands that explicitly acknowledge the multilevel structure, particularly in the presence of informative cluster size (Kahan et al., 2023).
-A promising avenue is to expand the semiparametric and nonparametric methods of Wang et al.
-(2024) to accommodate cluster-average and individual-average PGCE estimands when the intermediate variable is measured at the individual level.
-For longitudinal data, both intermediate and final outcomes may be repeatedly measured over time, which requires reformulating the identification assumptions to incorporate sequential ignorability of treatment and multilevel monotonicity (Frangakis et al., 2004).
-It is also worth noting that, while our main exposition has focused on a single final outcome, our framework naturally extends to multiple hierarchically ordered outcomes—a setting that has gained considerable traction in clinical research over the past decade.
-Hierarchical composite endpoints can be viewed as clustered outcome vectors within individuals.
-In Section S8 of the Online Supplement, we extend the notation of Bebu and Lachin (2015) to the potential outcomes framework and show how our results apply to this setting by defining pairwise comparison estimands with hierarchical outcome components while maintaining the same binary intermediate variable.
+For clustered data, recent work on cluster randomized trials has emphasized the need to define estimands that explicitly account for the multilevel structure, particularly in the presence of informative cluster size (kahan2023estimands).
+A promising avenue is to expand the semiparametric and nonparametric methods of wang2024model to accommodate cluster-average and individual-average PGCE estimands when the intermediate variable is measured at the individual level.
+For longitudinal data, both intermediate and final outcomes may be repeatedly measured over time, which requires reformulating the identification assumptions to incorporate sequential ignorability of treatment and multilevel monotonicity (frangakis2004methodology).
+It is also worth noting that, while our main exposition has focused on a single final outcome, our framework naturally extends to multiple hierarchically ordered outcomes, a setting that has gained considerable traction in clinical research over the past decade.
+In Section S8 of the Online Supplement, we extend the notation of Bebu2015 to the potential outcomes framework and show how our results apply to this setting by defining pairwise comparison estimands with hierarchical outcome components while maintaining the same binary intermediate variable.
 Further methodological developments along these directions will continue to broaden the applicability of our framework.